### v2

# appReferences in the Appendix

Abstract
Online linear programming (OLP) has found broad applications in revenue management and resource allocation. State-of-the-art OLP algorithms achieve low regret by repeatedly solving linear programming (LP) subproblems that incorporate updated resource information. However, LP-based methods are computationally expensive and often inefficient for large-scale applications. In contrast, recent first-order OLP algorithms are more computationally efficient but typically suffer from worse regret guarantees. To address these shortcomings, we propose a new algorithm that combines the strengths of LP-based and first-order OLP methods. The algorithm re-solves the LP subproblems periodically at a predefined frequency $f$ and uses the latest dual prices to guide online decision-making. In addition, a first-order method runs in parallel during each interval between LP re-solves, smoothing resource consumption. Our algorithm achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret, delivering a “wait-less” online decision-making process that balances the computational efficiency of first-order methods and the superior regret guarantee of LP-based methods.
1 Introduction
Sequential decision-making has garnered increasing attention for its utility in guiding optimal strategies in dynamic environments. It addresses the problems of identifying the optimal decisions and policies in environments where knowledge of the system continuously accumulates and evolves over time. In this paper, we study Online Linear Programming (OLP) [1], a powerful paradigm that encapsulates the core principles of sequential decision-making. OLP methods have been extensively applied to diverse areas, including resource allocation [2], online advertising [19], and inventory management [21].
In the OLP problem, customers arrive sequentially, each presenting a resource request and an associated bidding price. The objective is to determine which resource requests to fulfill to maximize total revenue while ensuring that no resource is ever over-allocated. The key challenge is that decisions must be made immediately and irrevocably, relying solely on historical data without knowledge of future arrivals. The goal is to minimize the cumulative regret with respect to the optimal hindsight linear programming (LP) solution.
Most OLP algorithms use online learning methods to gradually refine their estimates of the optimal dual prices, using these values to guide near-optimal decision-making. Commonly, these algorithms rely either on LP-based or first-order methods. LP-based methods repeatedly solve linear programs at each timestep with all available information, thereby updating the dual prices. However, the substantial computational demands of LP-based methods often restrict their application in time-sensitive settings. In contrast, first-order methods leverage gradient information for quick, incremental updates to the dual prices. However, they generally cannot achieve the same $\mathcal{O}(\log{T})$ regret guarantee as the LP-based methods [16]. These trade-offs motivate an open question at the intersection of online learning and decision making:
Can we simultaneously achieve low regret and efficiency?
We answer this question affirmatively in the well-studied stochastic input model–where the resource request and bidding price for each customer are sampled identically and independently from an unknown distribution—by proposing a decision-making framework that combines LP-based and first-order OLP algorithms. Our algorithm iteratively optimizes the dual prices, which progressively converge to the optimal dual solution.
1.1 Main Contributions
We propose a parallel multi-stage framework for OLP that integrates online learning and decision-making. Building on this framework, we develop a mathematical analysis that reconciles existing OLP approaches, ultimately allowing us to present an algorithm that achieves logarithmic worst-case regret. We formalize the trade-off between decision quality and computational efficiency, enabling end-users to quantify how allocating additional computational resources will reduce regret. In more detail, our contributions can be summarized as follows:
  • Parallel Multi-stage Framework: We separate online learning and decision-making into two separate yet interlinked processes, connected through a feedback loop. We re-solve the LP subproblems periodically at a fixed frequency and pass the result to the decision-making path. These results guide decision-making until the next time that the LP is re-solved. Additionally, to enhance the algorithm’s efficiency, we apply a first-order method during the initial and final stages, restarting with the most recent learning outcomes. By integrating the LP-based and first-order techniques, this approach leverages their respective strengths.
  • Regret Analysis: We develop a mathematical analysis for our algorithm where we derive a new performance metric to account for the inter-dependency between the LP-based and first-order methods. We unify the analysis of these two methods and provide a “spectrum theorem” that bounds regret for any feasible choice of the re-solving frequency in $[1,T]$:
    Theorem 1.1 (Informal version of Theorem 3.2).
    If we re-solve the updated linear programs every $f$ timesteps over the total time horizon of length $T$ and utilize gradient information in the initial and final $f$ steps, our algorithm is guaranteed to achieve a worst-case regret of
    $\mathcal{O}\big{(}\log\big{(}\tfrac{T}{f}\big{)}+\sqrt{f}\big{)}.$ (1)
    When $f=1$, the algorithm is a pure LP-based method and reduces to [16], and when $f=T$, the algorithm is a pure first-order method with regret $\mathcal{O}(\sqrt{T})$. In our analysis, one can choose the “optimal” $f$ to balance the LP-solving times and regret. Compared to [23], our method enables a “wait-less” decision-making system for customers during the whole process.
1.2 Related Literature
Online resource allocation and OLP have been widely studied under two predominant models: the stochastic input model [9, 6] and the stochastic permutation model [1, 11]. In our paper, we study the stochastic input model, where each customer’s resource request and bidding price—represented as columns of the constraint matrix—are drawn i.i.d. from an unknown distribution $\mathcal{P}\in\Xi$.
Table 1: Performances in current OLP literature
Paper Setting Algorithm Regret Decision-Making
[16] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T\log\log T)$ Delay
[4] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
[17] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
[12] Bounded, finite support LP-based $\mathcal{O}(\log^{2}T)$ Delay
[5] Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
[14] Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
[23] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T/f)$ Delay
This paper Bounded, continuous support, non-degeneracy LP-based & First-order ${\color[rgb]{1,0,0}\mathcal{O}(\log(T/f)+\sqrt{f})}$ No Delay
[15] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[2] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[7] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[3] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[20] Bounded, finite support, non-degeneracy First-order $\mathcal{O}(T^{3/8})$ No Delay
[8] Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(T^{1/3})$ No Delay
[17] Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(\log^{2}T)$ No Delay
LP-based Methods
This class of algorithms focuses on the OLP dual problem, repeatedly solving it to update the decision variables. While these methods achieve low regret, they are computationally expensive. The dual LPs are formulated based on the information revealed by each arriving customer. Early versions of these algorithms maintained a fixed average resource consumption rate throughout the time horizon [1], whereas more recent variants adaptively adjust the consumption rate based on the remaining resources, updating the LP accordingly [16]. Under standard non-degeneracy and continuous support assumptions, these LP-based methods can achieve a regret bound of $\mathcal{O}(\log T)$. Several related works explore variants of this approach: [4] studies a multi-secretary problem and also attains $\mathcal{O}(\log T)$ regret; [17] incorporates a regularization term on the resource and also achieves $\mathcal{O}(\log T)$ regret; [12] assumes that the resource distribution has a finite support and obtains $\mathcal{O}(\log^{2}T)$ regret; and [5] makes a stronger assumption that both the resource and price distributions have finite support, leading to a constant regret.
First-order Methods
These algorithms leverage gradient information to update the decision variables directly without explicitly solving LPs. Although less accurate than LP-based methods, they enable more efficient computation and are therefore well-suited for large-scale or time-sensitive applications. First-order methods can attain a regret of $\mathcal{O}(\sqrt{T})$ with mirror descent and subgradient information under distributions with continuous support [15, 2], and $\mathcal{O}(T^{3/8})$ under non-degeneracy assumptions for finite support distributions [20]. Furthermore, [7] attains a $\mathcal{O}(\sqrt{T})$ regret under proximal point updates, while [3] obtains the same regret via a momentum variant of mirror descent. [8] utilizes a restart strategy to decouple the learning and decision-making processes and improves the regret to $\mathcal{O}(T^{1/3})$. [17] integrates the resource consumption rate adjustment from LP-based methods into first-order methods and obtains $\mathcal{O}(\log^{2}T)$ regret.
Delay in Decision-making
This line of research examines the impact of decision delays in online resource allocation. [10] analyzes an online setting with a mix of impatient customers, whose requests must be decided upon immediately and irrevocably, and partially patient customers, for whom decisions can be postponed for several time periods. [22] demonstrates that deferring decisions—either by introducing explicit delays or batching customers—can substantially reduce regret. In concurrent research, [23] partially solves this delay issue by solving LPs in batches. While their approach is similar in spirit to ours, they make a stronger assumption that each resource request has a positive lower bound in expectation and their algorithm still requires customers to wait for decisions in the initial and final batches. Compared with the aforementioned works, our framework imposes minimal assumptions on the problem parameters, attains a regret bound of $\mathcal{O}(\log(T/f)+\sqrt{f})$, and provides a “wait-less” decision-making process throughout the entire time horizon.
The rest of the paper is organized as follows. Section 2 introduces the problem formulation, preliminary dual convergence results, and the main assumptions. Section 3 proposes our novel framework to integrate the LP-based and first-order methods and provides our “spectrum theorem,” which bounds the regret of our algorithm by $\mathcal{O}(\log(T/f)+\sqrt{f})$ for any choice of $f$. Section 4 presents experiments that illustrate the strong performance of our algorithm and validate our theoretical conclusions.
2 Problem Setup
[…]
2.1 Online LP Formulation
Consider an online resource allocation problem over the time horizon $T$. Initially, we have an inventory vector $\bm{b}\in\mathbb{R}^{m}$, representing $m$ types of resources. At each time step $t$, a customer arrives with a request characterized by $(r_{t},\bm{a}_{t})\sim\mathcal{P}$, where $r_{t}\in\mathbb{R}$ is the offered payment (bid), $\bm{a}_{t}\in\mathbb{R}^{m}$ is the customer’s demand for each resource, and $\mathcal{P}$ is a fixed, unknown distribution. We must decide whether to accept or reject this request, represented by the decision variable $x_{t}\in\{0,1\}$. The goal is to maximize the cumulative reward. This problem can be formulated as the following OLP, which we refer to as the primal linear program (PLP):
$\max_{\bm{x}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{t=1}^{T}r_{t}x_{t}$ (PLP)
subject to $\leavevmode\nobreak\ \quad\sum_{t=1}^{T}a_{it}x_{t}\leq b_{i},i=1,...,m$
[…]
The dual problem of (PLP) is given by
$\min_{\bm{p},\bm{y}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=1}^{m}b_{i}p_{i}+\sum_{t=1}^{T}y_{t}$ (DLP)
subject to $\leavevmode\nobreak\ \quad\sum_{i=1}^{m}a_{it}p_{i}+y_{t}\geq r_{t},t=1,...,T$
$\quad\ \ p_{i},y_{t}\geq 0,\forall i,t,$
where $\bm{p}$ is the vector of dual prices. Let $\bm{d}=\bm{b}/T\in\mathbb{R}^{m}$ denote the initial average resource capacity. As demonstrated by [15], (DLP) can be written as
$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f_{T}(\bm{p})\coloneqq\frac{1}{T}\sum_{t=1}^{T}\left[\sum_{i=1}^{m}d_{i}p_{i}+\left(r_{t}-\sum_{i=1}^{m}a_{it}p_{i}\right)^{+}\right]$ (2)
subject to $\leavevmode\nobreak\ \quad p_{i}\geq 0,i=1,...,m$
This can be regarded as a $T$-sample approximation of the following stochastic program:
$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f(\bm{p})\coloneqq\mathbb{E}[f_{T}(\bm{p})]=\bm{d}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (3)
subject to $\leavevmode\nobreak\ \quad\bm{p}\geq\bm{0}$
where the expectation is taken with respect to the coefficient pair $(r,\bm{a})$ drawn from the unknown distribution $\mathcal{P}$.
We define the optimal solutions to the $T$-sample approximation problem (2) and the stochastic program (3) respectively as
$\bm{p}_{T}^{*}=\arg\min_{\bm{p}\geq\textbf{0}}\leavevmode\nobreak\ f_{T}(\bm{p})\quad\text{and}\quad\bm{p}^{*}=\arg\min_{\bm{p}\geq\textbf{0}}\leavevmode\nobreak\ f(\bm{p}).$
The decision variables $x_{t}^{*}$ can be established from the complementary slackness condition as
$x_{t}^{*}=\left\{\begin{array}[]{ll}\ 0,&r_{t}<\bm{a}_{t}^{\top}\bm{p}_{T}^{*},\vspace{3 pt}\\ \ 1,&r_{t}>\bm{a}_{t}^{\top}\bm{p}_{T}^{*}.\end{array}\right.$ (6)
and $x_{t}^{*}\in[0,1]$ if $r_{t}=\bm{a}_{t}^{\top}\bm{p}_{T}^{*}$.
This connection between the primal-dual optimal solutions motivates us to propose dual-based online LP algorithms and then provide a convergence analysis of $\bm{p}_{T}^{*}$ to $\bm{p}^{*}$.
2.2 Online Dual Algorithms
We introduce two basic dual-based algorithms for the OLP problem. We calculate and update the dual prices $\{\bm{p}_{t}\}_{t=1}^{T}$ and compute decisions $\{x_{t}\}_{t=1}^{T}$ based on rule (6).
2.2.1 LP-based Algorithm
In this method, we calculate the dual price by re-solving the online LP problem at each time step. The model is dynamically updated to incorporate real-time capacity constraints derived from the Action-history-dependent learning (AHDL) algorithm [16]. Define $d_{it}=\frac{b_{it}}{T-t}$ to be the average remaining resource capacity for resource $i$ at time $t$. The resulting optimization problem can be viewed as a $t$-sample approximation to the stochastic program specified by $\bm{d}_{t}=(d_{it},...,d_{mt})^{\top}$ as
$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f_{\bm{d}_{t}}(\bm{p})\coloneqq\bm{d}_{t}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (7)
subject to $\leavevmode\nobreak\ \quad\bm{p}\geq\bm{0}.$
We denote the optimal solution to the stochastic program (7) as $\bm{p}_{t}^{*}$. Thus, instead of focusing on a static problem with fixed initial $\bm{d}$, we formulate a dynamic stochastic program that evolves according to the changing constraints. We outline this method in Algorithm 1.
Algorithm 1 LP-based Online AHDL algorithm
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0}.$
for $t$ = $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6) if constraints are not violated. Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Solve the updated dual problem and obtain dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
end for
This method incorporates the past decisions into the optimization of $\bm{p}_{t}$. If resources were over-utilized in earlier periods, the remaining supply decreases, prompting the algorithm to raise the dual price and become more selective with future orders. Conversely, if ample resources remain, the future dual price will be lowered, allowing more consumer requests to be accepted. This adaptive mechanism accounts for past actions by adjusting the available resource capacity.
2.2.2 Subgradient-based Algorithm
In this method, we calculate the dual price by iteratively moving in the direction of the steepest cost gradient. We maintain a static resource constraint, as the historical information is integrated through the gradient updates. Due to its lower computational overhead, the subgradient-based is well-suited for high-dimensional and real-time applications. Algorithm 2 summarizes this method, which is also known as the “first-order” method in [15].
Algorithm 2 First-Order Online algorithm
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0}.$
for $t$ = $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6). Update learning rate $\alpha_{t}=\frac{1}{t+1}$. Compute subgradient and obtain dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ (\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}+\frac{1}{2\alpha_{t}}\lVert\bm{p}-\bm{p}_{t}\rVert^{2}$
end for
We can derive a closed-form solution for the updated dual price from the above rule as
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$
where the subgradient term evaluated at $\bm{p}_{t}$ is calculated as
$\bm{d}-\bm{a}_{t}x_{t}\in\partial_{\bm{p}=\bm{p}_{t}}\big{(}\bm{d}^{\top}\bm{p}+(r_{t}-\bm{a}_{t}^{\top}\bm{p})^{+}\big{)}.$
Thus, this process can be interpreted as a projected stochastic subgradient descent method for solving the optimization problem (2). The method reduces computational costs by requiring only a single pass through the data and eliminates the need for solving LPs.
2.3 Performance Metrics
We evaluate the algorithms using a bi-objective performance measure that involves the optimality gap and constraint violation. The optimality gap measures the difference between the objective value of the algorithm’s output and that of the true optimal solution, while the constraint violation measures the degree to which the algorithm’s output fails to meet the given constraints.
We denote the order-request matrix by $\bm{A}=[\bm{a}_{1},...,\bm{a}_{T}]^{\top}$, the offline optimal solution to problem (PLP) by $\bm{x}^{*}=(x_{1}^{*},...,x_{T}^{*})^{\top}$, and the online optimal solution by $\bm{x}=(x_{1},...,x_{T})^{\top}$. Then, we define the regret $r(\bm{x})$ and resource violation $v(\bm{x})$ as:
$r(\bm{x})$ $\coloneqq\max_{\mathbf{A}\mathbf{x}\leq\bm{b},\mathbf{0}\leq\mathbf{x}\leq\mathbf{1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{t=1}^{T}r_{t}x_{t}^{*}-r_{t}x_{t},$ (8)
$v(\bm{x})$ $\coloneqq\lVert(\bm{Ax}-\bm{b})^{+}\rVert.$ (9)
Therefore, we define the following bi-objectives for evaluating the total regret and worst-case total regret of the algorithms we propose:
$\Delta_{T}=\sup_{\mathcal{P}\in\Xi}\mathbb{E}_{\mathcal{P}}[r(\bm{x})+v(\bm{x})]$ (10)
where $\mathcal{P}$ is the unknown distribution of $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ and $\Xi$ denotes the family of distribution satisfying some regularity assumptions specified later.
This metric is commonly used for first-order OLP algorithms in [7, 15] and is also aligned with the literature on the online convex optimization with constraints in [18, 24]. For LP-based methods, [16] use stopping time to bound the regret. Equation (10) integrates the analysis of subgradient methods and LP-based methods into a single performance metric, enabling more consistent analysis across online learning scenarios. This integrated method promotes more balanced resource consumption over the decision horizon. This flexibility makes our framework well-suited for scenarios without prior knowledge of the constraint support, thereby expanding its applicability to a broad range of online learning and resource allocation problems.
2.4 Assumptions and Auxiliary Results
We propose the following assumptions regarding the stochastic inputs used throughout the paper. These minimal assumptions are standard in the online learning literature [16, 13, 23]. In particular, we require the input data to be bounded, follow a linear growth, and be non-degenerate.
Assumption 1 (Boundedness).
We assume
  1. (a)
    The order inputs $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ are generated i.i.d from an unknown distribution $\mathcal{P}$.
  2. (b)
    There exist constants $\bar{r},\bar{a}>0$ such that $|r_{t}|\leq\bar{r}$ and $\lVert\bm{a}_{t}\rVert_{\infty}\leq\bar{a}$ almost surely for $t=1,...,T$.
  3. (c)
    The average resource capacity $\bm{d}=\bm{b}/T$ satisfies $d_{i}\in[\underline{d},\bar{d}]$ for some $\bar{d}>\underline{d}>0$ for any $i=1,...,m$.
In this assumption, (a) states that $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ are independent of each other, but we allow dependencies between their components. Part (b) introduces the bounds $\bar{r},\bar{a}$ solely for analytical purposes; they are not used within the algorithm itself. This is a minimal requirement on $(r_{t},\bm{a}_{t})$ compared to previous work [23]. Part (c) requires the average resource capacity to grow linearly with $T$, ensuring that a constant fraction of the $x_{t}$ values can be set to 1 in the optimal solutions. Consequently, the number of fulfillable orders is proportional to $T$, facilitating a stable service level over time.
Assumption 2 (Uniform Non-degeneracy).
We assume
  1. (a)
    The second moment matrix $\mathbb{E}[\bm{a}\bm{a}^{\top}]$ is positive definite with minimum eigenvalue $\lambda$.
  2. (b)
    There exists constants $\mu,\nu$ such that for any $(r,\bm{a})\sim\mathcal{P}$,
    $\nu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|\leq|\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}\mid\bm{a})-\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}^{*}\mid\bm{a})|\leq\mu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|$
    holds for all $\bm{p}\in\mathcal{V}_{\bm{p}}\coloneqq\{\bm{p}\in\mathbb{R}^{m}:\bm{p}\geq\bm{0},\lVert\bm{p}\rVert\leq\frac{\bar{r}}{\underline{d}}\}$ and $\bm{d}\in\mathcal{V}_{\bm{d}}\coloneqq[\underline{d},\bar{d}]^{m}$.
  3. (c)
    The optimal solution $\bm{p}^{*}$ satisfies $p_{i}^{*}=0$ if and only if $d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]>0$ for all $\bm{d}\in\mathcal{V}_{\bm{d}}$ and $i=1,...,m$.
In this assumption, part (b) ensures that the cumulative distribution of the reward given the resource consumption request $r|\bm{a}$ is continuous and exhibits a stable growth rate. Part (c) requires strict complementarity for the optimal solutions of the stochastic program in (3), which is a non-degeneracy condition for both the primal and dual LP.
According to stochastic program (3), we define the binding and non-binding index sets as:
$I_{B}$ =$=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})=0]\},$
$I_{N}$ =$=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})>0]\}.$ (11)
Here, $I_{B}$ denotes the set of binding resources—those that are limited and require careful allocation—while $I_{N}$ denotes the set of non-binding resources, which are unlimited and can fulfill any order’s request. By Assumption 2(c), these sets are complements, as $I_{B}\cap I_{N}=\emptyset$ and $I_{B}\cup I_{N}=\{1,...,m\}$.
Let $\Xi$ denote the family of distributions satisfying Assumptions 1 and 2. We propose the following lemma.
Lemma 2.1 (Dimension Stability).
Under Assumptions 1 and 2, there exists a constant $\delta>0$ such that
[…]
This lemma is a consequence of Lemmas 12 and 13 in [16]. The existence of $\delta$ comes from the continuity of $f_{\bm{d}_{t}}(\bm{p})$. Thus, $\delta$ is only associated with stochastic program (7), and it is independent of $T$. Note that $\mathcal{D}\subset\mathcal{V}_{\bm{d}}$. We will analyze our feasible dual solutions derived from the online algorithm with $\bm{d}_{t}\in\mathcal{D}$.
Lemma 2.2 (Quadratic Regularity, Proposition 2 in [16]).
Under Assumptions 1 and 2, for any $\bm{p}\in\mathcal{V}_{\bm{p}}$,
$f(\bm{p})$ $\leq f(\bm{p}^{*})+\nabla f(\bm{p}^{*})^{\top}(\bm{p}-\bm{p}^{*})+\frac{\mu\bar{a}^{2}}{2}\lVert\bm{p}-\bm{p}^{*}\rVert^{2},$
$f(\bm{p})$ $\geq f(\bm{p}^{*})+\nabla f(\bm{p}^{*})^{\top}(\bm{p}-\bm{p}^{*})+\frac{\nu\lambda}{2}\lVert\bm{p}-\bm{p}^{*}\rVert^{2}.$
Moreover, $\bm{p}^{*}$ is the unique optimal solution to (3).
This lemma establishes a local form of semi-strong convexity and smoothness at $\bm{p}^{*}$, which is guaranteed by our assumptions on the distribution $\mathcal{P}\in\Xi$. By focusing on these local properties rather than insisting on strong convexity and global smoothness, we relax the classical requirements typically imposed in such settings. In later sections, we will leverage this result to derive our regret bound.
3 Parallel Multi-Phase Online Learning and Decision-Making
In this section, we present our novel framework for parallel multi-phase online learning and decision-making. Our approach is motivated by the difficulties of existing online resource allocation methods, which involve either high regret or heavy computational costs. Specifically, the LP-based AHDL Algorithm 1 has a strong worst-case regret bound of $\mathcal{O}(\log T)$, but its high computational costs delay decision-making. Meanwhile, the first-order online Algorithm 2 updates decisions efficiently but suffers a high regret of $\mathcal{O}(\sqrt{T})$. By combining the strengths of these two methods, our new framework balances decision-making quality and computational efficiency, thereby overcoming the limitations of existing approaches.
3.1 Algorithm Design
We establish our framework with the following design:
  1. 1.
    We maintain two parallel processes—one dedicated to online learning and the other to online decision-making. The online learning results are periodically sent to the decision-making path to guide subsequent updates.
  2. 2.
    In the online learning path, we employ an enhanced LP-based AHDL algorithm, which only re-solves the updated OLP problem according to a predefined frequency, significantly reducing the computational overhead.
  3. 3.
    In the online decision-making path, we apply the first-order method exclusively during the initial and final re-solving batches, where the learning rate is optimally tuned for these two intervals.
Figure 1: Parallel Multi-Phase Online Learning and Decision-Making Algorithm
Figure 1 illustrates our framework, which generates two parallel sequences of dual solutions: $\{\bm{p}_{t}^{D}\}_{t=1}^{T}$ from a first-order method and $\{\bm{p}_{t}^{L}\}_{t=1}^{T}$ from an LP-based approach. The algorithm proceeds in batches of length $f$. In the first batch, it uses a first-order method to iteratively update dual prices, thereby guiding the decision-making process. At the end of this batch, it applies the LP-based AHDL algorithm to obtain a refined set of dual prices, passing these updated values to the decision-making path. This LP-based re-solving occurs only once per batch, making the approach computationally efficient. Formally, we re-solve the OLP at every time $t$ satisfying:
$t\leq kf\quad\text{and}\quad t\bmod f=0,$ (12)
where $k=\left\lfloor\frac{T}{f}\right\rfloor$ represents the number of batches.
After the first batch, the algorithm does not employ the first-order method until the final batch. Instead, during intermediate batches, the decision-making path uses the most recent dual prices computed from the LP at the end of the previous batch. Only in the final batch does the algorithm restart the first-order updates to guide the final set of decisions. Algorithm 3 summarizes this approach.
Algorithm 3 Parallel Multi-Phase Online Learning and Decision-Making
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of re-solving batches $k=\lfloor T/f\rfloor$.
for $t=$ $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6). Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$. if $t\leq f$ or $t\geq kf$ then
             Compute the subgradient and update the dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$ (13)
       end if
      if $t\bmod f=0$ and constraints are not violated then
             Solve the online dual problem and update dual price $\bm{p}_{t+1}$ as
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$ (14)
       end if
      else
             Update dual price $\bm{p}_{t+1}$ to be the most recent solution $\bm{p}_{t+1}=\bm{p}_{t}$.
       end if
end for
We achieve a strong synergy between an effective online learning method and an efficient decision-making procedure. This structure enables the algorithm to remain responsive to dynamic environments and adapt to the latest information. Our strategy of periodic re-solving greatly reduces the computational costs. Therefore, we establish a “wait-less” online decision-making framework in which each customer’s order is processed immediately without waiting for the completion of earlier requests.
3.2 Algorithm Analysis
We next present several lemmas that characterize the convergence of the online dual solutions $\{\bm{p}_{t}\}_{t=1}^{T}$ generated by Algorithm 3 toward the optimal solution $\bm{p}^{*}$ of the stochastic program (3). In particular, let $\bm{p}_{t}^{D}$ be the dual prices updated via the first-order rule (13), and let $\bm{p}_{t}^{L}$ be those obtained from the LP-based method (14) at time $t$. These lemmas will then be used to establish upper bounds on the worst-case regret of our algorithm.
Lemma 3.1 (Dual Price Boundedness).
Under Assumption 1, the online and offline optimal dual prices are bounded respectively by:
$\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}},\ \ \ \lVert\bm{p}_{t}\rVert\leq\bar{p}\coloneqq\max\left(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right).$ (15)
This lemma establishes that the optimal dual prices remain bounded. Our algorithm maintains these bounds because if $\bm{p}_{t}$ grows large, the algorithm responds by accepting more orders, which in turn reduces $\bm{p}_{t+1}$. This self-correcting mechanism keeps the dual prices within these limits. Moreover, drawing on results of [15], we have that
$\lVert\bm{p}_{t}^{D}\rVert\leq\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})$
and from [16], we have that
$\lVert\bm{p}_{t}^{L}\rVert\leq\frac{\bar{r}}{\underline{d}-\delta}.$
Consequently, we can prove the following convergence lemmas.
Lemma 3.2 (Dual Convergence of LP-based algorithm).
Under Assumptions 1 and 2, there exists a constant $C_{lp}>0$ depending on $\bar{r},\bar{a},\underline{d},m,\nu$, and $\lambda$ such that
$\mathbb{E}\left[\lVert\bm{p}_{t}^{L}-\bm{p}_{t}^{*}\rVert^{2}\right]\leq\frac{C_{lp}}{t}.$ (16)
In addition, the difference between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ satisfies
$\lVert\bm{p}_{t}^{*}-\bm{p}^{*}\rVert^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\lVert\bm{d}_{t}-\bm{d}\rVert^{2}.$ (17)
This lemma establishes that the LP-based online AHDL algorithm [13] produces dual solutions $\bm{p}_{t}$ that converge to $\bm{p}_{t}^{*}$. This convergence highlights the stability of the online dual variables during the intermediate stages of decision-making and ensures a warm start for the first-order method in the last re-solving batch. Moreover, we bound the distance between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ by relating their respective average resource capacities, $\bm{d}_{t}$ and $\bm{d}$, in the associated stochastic programs [16]. These results provide the foundation for analyzing $\lVert\bm{p}_{t}-\bm{p}^{*}\rVert$ in our final result.
Lemma 3.3 (Dual convergence of First-order algorithm).
Under Assumptions 1 and 2, if $\alpha_{t}<\nu\lambda$, the subgradient updates satisfy the following recursion rule:
$\mathbb{E}\left[\lVert\bm{p}_{t+1}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq(1-\alpha_{t}\nu\lambda)\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}.$ (18)
Case 1. if $\alpha_{t}\equiv\alpha<\frac{1}{\nu\lambda}$, then there exists a constant $C_{fo}=\frac{\bar{p}^{2}+m(\bar{a}+\bar{d})^{2}}{\nu\lambda}$ such that
$\mathbb{E}\left[\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq C_{fo}\left(\frac{1}{\alpha t}+\alpha\right).$ (19)
Case 2. if $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, then there exists a constant $C_{fo}=\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}}$ such that
$\mathbb{E}\left[\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq\frac{C_{fo}}{t}.$ (20)
This lemma shows that the first-order method [8] guarantees the convergence of $\bm{p}_{t}$ to $\bm{p}^{*}$. It also highlights a key trade-off in choosing the learning rate $\alpha_{t}$, an aspect that will be central to our later optimality analysis. With this groundwork in place, we now move on to bounding the total regret of our algorithm.
Theorem 3.1 (Total Regret).
Under Assumptions 1 and 2, the total regret $\Delta_{T}$ of Algorithm 3 satisfies:
$\Delta_{T}\ \leq\ \mu\bar{a}^{2}\sum_{t=1}^{T}\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]+\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\left[\lVert(\bm{b}-\bm{Ax})^{B+}\rVert\right]+\mathbb{E}\left[\lVert(\bm{Ax}-\bm{b})^{+}\rVert\right].$ (21)
where $(\cdot)^{B+}$ indicates that we only consider the binding terms and project them onto the positive orthant.
We express the total regret as the sum of two key components: dual convergence and resource consumption. This bound demonstrates that the algorithm encourages (1) smooth and balanced resource utilization, especially for the binding terms, and (2) full resource consumption by the end of the time horizon. Building on Theorem 3.1, we next introduce a “spectrum theorem” for the algorithm’s total regret.
Theorem 3.2 (Spectrum Theorem).
Under Assumptions 1 and 2, the worst-case regret $\Delta_{T}$ of Algorithm 3 is bounded by
$\Delta_{T}\in\mathcal{O}\left(\sqrt{f}+\log\left(\frac{T}{f}\right)+f^{1/3}\right)=\mathcal{O}\left(\log\left(\frac{T}{f}\right)+\sqrt{f}\right),$ (22)
where $T/f$ represents the number of re-solving batches and $f$ is the length of each batch.
Remark 3.1.
The Spectrum Theorem elucidates the trade-offs in total regret induced by our use of LP-based and first-order methods. As stated in Section 2, utilizing pure LP-based method at every step yields a regret of $\mathcal{O}(\log T)$, while applying pure first-order method continuously attains a regret of $\mathcal{O}(\sqrt{T})$. Our algorithm sets a re-solving frequency $f\in[1,T]$ and achieves the regret of $\mathcal{O}(\log(T/f)+\sqrt{f})$, which aligns with previous results with the extreme cases of $f=1$ and $f=T$. By selecting an appropriate re-solving frequency $f$, our algorithm leverages the strengths of both approaches. Specifically, first-order method gives us the regret of $\sqrt{f}$ for the first batch and $f^{1/3}$ for the last batch, and LP-based method produces the regret of $\log(T/f)$.
Remark 3.2.
We can obtain a tighter bound of
$\Delta_{T}\in\mathcal{O}(\log(T)+f^{1/3})$ (23)
if we have a warm start of the initial dual price for the first batch satisfying $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ or we use the LP-based method to learn at each time. Both these approaches can give us the improvement from $\sqrt{f}$ to $f^{1/3}$ for the first-batch regret.
Technical Intuitions.
We prove this theorem using several novel techniques. As stated in Theorem 3.1, we develop a comprehensive framework that accommodates both the first-order and LP-based methods, allowing us to decompose the total regret into distinct components associated with each approach. Specifically:
  • For the first-order method, we relate the distance $\bm{Ax}-\bm{b}$ to the distance between dual prices.
  • For the LP-based method, as the algorithm progresses toward the end of the time horizon, if the current average resource consumption rate $\bm{d}_{t}$ exceeds the allowed limit, we manually set the dual prices to zero to accept all future orders. This approach enables us to adapt the concept of a “stopping time” within our new framework and derive a rigorous bound of regret.
Through this new framework, we integrate the LP-based and subgradient-based methods to achieve “wait-less” decision-making, where each customer’s order is processed immediately without delays, while also attaining the best-known logarithmic regret bound.
3.3 Algorithm Extension
Building on Algorithm 3, we propose an enhanced multi-stage learning algorithm that employs the first-order method for decision-making during each re-solving interval. This multi-restart mechanism further improves the accuracy and efficiency of our framework. We illustrate this algorithm in the following figure.
Figure 2: Enhanced Parallel Multi-Phase Online Learning and Decision-Making Algorithm
Figure 2 illustrates this variant of Algorithm 3 with multi-stage exploration and exploitation. Rather than make decisions solely based on the most recently solved dual prices, the algorithm treats the updated dual prices at each re-solving interval as a new starting point for the first-order learning process. For example, in Algorithm 3, the dual price $\bm{p}_{f}^{L}$ is used to guide decision-making during the time interval $t\in[f,2f]$. In contrast, the enhanced algorithm restarts the first-order method by setting $\bm{p}_{f}^{L}=\bm{p}_{f}^{D}$ and then continues to fine-tune the dual prices $\{\bm{p}_{t}^{D}\}_{t=f}^{2f}$ to make up-to-date decisions. Algorithm 4 summarizes this approach.
Algorithm 4 Enhanced Parallel Multi-Phase Online Learning and Decision-Making
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of re-solving batches $k=\lfloor T/f\rfloor$.
for $t=$ $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6). Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$. if $t\bmod f\neq 0$ then
             Compute the subgradient and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$
       end if
      if $t\bmod f=0$ and constraints are not violated then
             Solve the online dual problem and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
       end if
end for
This algorithm establishes an online framework of multi-stage learning and decision-making. The incorporation of an intermediate first-order method improves the algorithm’s stability and ensures smooth resource consumption. These improvements result in better performance during the final batch, thereby reducing the algorithm’s total regret.
3.4 Algorithm Application
Our motivation in designing Algorithm 3 is to develop a hybrid decision-making system that effectively balances computational efficiency and decision accuracy. By integrating the first-order method, we aim to reduce computational costs, while the LP-based method enhances the precision of our decisions. Building on the Spectrum Theorem 3.2, this section aims to translate our theoretical results into practical applications. Specifically, we formulate a new optimization problem to determine the optimal re-solving frequency, taking into account computational resource constraints.
We use $c_{1}(\cdot),c_{2}(\cdot)$ to denote the computational cost functions for the LP-based method and first-order method, respectively, and use $R$ to denote the total computational resource capacity. Then we define the following optimization problem to determine the optimal re-solving frequency $f$:
$\min_{f\in\mathbb{Z}^{+}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \log\Big{(}\frac{T}{f}\Big{)}+\sqrt{f}$ (24)
subject to $\leavevmode\nobreak\ \quad c_{1}(k)+2c_{2}(f)\leq R$
$\leavevmode\nobreak\ \quad 1\leq f\leq T$
We can further analyze the computational requirements of our algorithm by considering the specific LP solvers employed. Recall that $m$ denotes the dimension of the dual price vector. At any given time step $t$, utilizing either the interior-point method or the simplex method as the LP-based solver incurs a computational cost of $m^{2}(m+t)$. In contrast, the first-order method incurs a computational cost of $2mf$. We encapsulate these trade-offs in the following proposition.
Proposition 3.1 (Optimal Re-solving Frequency).
Given a fixed computational resource capacity $R$, if we use the interior-point method or simplex method as the LP solver in Algorithm 3, we can formulate the optimization problem (24) as
$\min_{f\in\mathbb{Z}^{+}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \log\Big{(}\frac{T}{f}\Big{)}+\sqrt{f}$ (25)
subject to $\leavevmode\nobreak\ \quad\sum_{b=1}^{k}\left(m^{2}(m+bf)\right)+2mf\leq R,$
$\leavevmode\nobreak\ \quad 1\leq f\leq T$ (26)
By solving optimization problem (25), we can select an optimal re-solving frequency $f$ that achieves the lowest regret within the computational resource limit $R$.
Therefore, we extend our original optimization problem for resource allocation to include computational resource allocation. This proposition enables users to determine the optimal re-solving frequency that balances regret with computational cost based on their available computational resources.
4 Numerical Experiments
We conduct extensive experiments under various settings to illustrate the performance of our algorithm and validate the theoretical results of our framework. This section is divided into two parts. In the first part (Section 4.1), we evaluate Algorithms 3 and 4 across different choices of the re-solving frequency. In the second part (Section 4.2), we compare our algorithm with the LP-based method in terms of running time and with the first-order method in terms of regret. Both sets of experiments are conducted under different distributions of $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ that satisfy the assumptions in Section 2.
4.1 Performance Evaluation under Varying Re-solving Frequencies
We generate the sequence $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ randomly from distributions that satisfy Assumptions 1 and 2. Specifically, we use the uniform and normal distribution, with details in Table 2. We select the types of resource $m\in\{1,5\}$, time horizon $T$ evenly spanned over $[10^{2},10^{6}]$, and initial average resource capacity $d_{i}$ sampled i.i.d. from $\text{Uniform}[1/3,2/3]$. We run our algorithm for $100$ trials and use the average as the final result for each experiment.
Table 2: Stochastic Input in Experiments
Distribution Model $\bm{a}_{t}$ $r_{t}$
Input I $a_{it}\sim\text{Uniform}[0,2]$ $r_{t}\bot\bm{a}_{t},r_{t}\sim\text{Uniform}[0,10]$
Input II $a_{it}\sim\text{Normal}(0.5,1)$ $r_{t}=\sum_{i=1}^{m}a_{it},r_{t}\sim\text{Normal}(0.5m,m)$
We evaluate our algorithm with $m=1$ under the stochastic input models I and II from Table 2. The experiments are conducted across various choices of $f\in\{T^{1/3},T^{1/2},T^{2/3}\}$ representing high, medium, and low re-solving frequencies. We analyze the absolute value of the regret and its growth rate over time. We illustrate our results in Table 3 and Figure 3 on a logarithmic scale.
Table 3: Regret of Algorithms with various re-solving frequency
Offline Low freq Mid freq High freq
$T=10^{3}$ 12.13 7.76 5.77 4.86
$T=10^{4}$ 38.50 10.96 7.55 5.67
$T=10^{5}$ 122.44 23.90 9.92 8.36
$T=10^{6}$ 404.59 56.70 21.90 8.99
(a) Algorithm 3, Input I
Offline Low freq Mid freq High freq
$T=10^{3}$ 11.44 6.28 4.86 3.95
$T=10^{4}$ 36.50 10.21 7.34 3.81
$T=10^{5}$ 115.57 14.61 11.78 4.66
$T=10^{6}$ 365.99 35.20 15.68 6.26
(b) Algorithm 3, Input II
Offline Low freq Mid freq High freq
$T=10^{3}$ 12.13 6.78 5.20 4.50
$T=10^{4}$ 38.50 10.37 8.03 5.99
$T=10^{5}$ 122.444 22.33 11.57 6.36
$T=10^{6}$ 404.59 48.21 22.44 7.09
(c) Algorithm 4, Input I
Offline Low freq Mid freq High freq
$T=10^{3}$ 11.44 3.20 2.56 1.75
$T=10^{4}$ 36.50 5.48 4.30 2.52
$T=10^{5}$ 115.57 12.35 4.48 3.86
$T=10^{6}$ 365.99 30.48 13.20 4.77
(d) Algorithm 4, Input II
(a) Algorithm 3 with input I
(b) Algorithm 3 with input II
(c) Algorithm 4 with input I
(d) Algorithm 4 with input II
Figure 3: Main Algorithms with Various Frequency
These experiments assess the performance of our two main algorithms under varying re-solving frequencies and across multiple time horizons. As presented in Table 3 and Figure 3, we observe that regret decreases as the re-solving frequency increases from $f=T^{2/3}$ to $f=T^{1/3}$. This trend holds consistently across both algorithms and input types, with higher re-solving frequencies achieving the lowest regret in all cases. In addition, while regret accumulates over longer time horizons, the rate of increase remains stable for algorithms employing higher re-solving frequencies. These findings illustrate the robustness of our algorithms, as more frequent updates enable better adaptation to dynamic environment.
4.2 Comparative Analysis with Baseline Methods
After examining the algorithm under various re-solving frequencies, we compare its performance with the classic first-order method and LP-based method in terms of regret and computation time. We generate $\{(r_{t},\bm{a}_{t}\}_{t=1}^{T}$ from the uniform distribution under the Input I model (Table 2). We set the number of resource types to $m=5$, the time horizon $T$ to range evenly over $[10^{2},10^{6}]$, and the average resource capacity $d_{i}$ to be sampled i.i.d from $\text{Uniform}[1/3,2/3]$. The re-solving frequency is fixed to be $f=T^{1/3}$. Each result is averaged over 100 trial runs, and the findings are summarized in Table 5 and Figure 5.
$T$ Regret Algorithm Compute Time(s)
$10^{3}$ $66.88$ Offline $0.001$
$2.80$ LP-based (Algorithm 1) $4.185$
$25.11$ First-Order (Algorithm 2) $0.002$
7.20 Algorithm 3 0.3
4.50 Algorithm 4 0.3
$10^{4}$ 115.32 Offline $0.008$
3.50 LP-based (Algorithm 1) $123.497$
60.39 First-Order (Algorithm 2) $0.013$
9.12 Algorithm 3 1.8
5.09 Algorithm 4 1.8
$10^{5}$ 203.20 Offline $0.118$
3.88 LP-based (Algorithm 1) $>3600$
73.07 First-Order (Algorithm 2) $0.108$
9.41 Algorithm 3 56.9
6.36 Algorithm 4 56.8
$10^{6}$ 351.91 Offline 1.211
5.50 LP-based (Algorithm 1) $>100000$
115.39 First-Order (Algorithm 2) 1.577
11.65 Algorithm 3 2155.9
7.09 Algorithm 4 2242.1
Table 4: Algorithms Comparison.
Figure 5: Regret for Various Algorithms
Table 5 and Figure 5 provide a comprehensive comparison of all the algorithms discussed in this paper. Focusing on our primary algorithms, Algorithm 3 and 4, we observe:
  1. 1.
    Our algorithms exhibit strong performance in terms of decision optimality. They achieve over a 20-fold improvement in regret compared to the offline solution and more than a $10$-fold improvement compared to the first-order method. These numerical results also corroborate the theoretical bounds in Theorem 3.2.
  2. 2.
    Our algorithms are computationally efficient. Their computation times exhibit a 100-fold improvement compared to the LP-based method while maintaining regret levels that are remarkably close to those of the LP-based method. This significant improvement in efficiency makes our algorithms highly suitable for real-time and large-scale applications.
Therefore, we achieve a balance between effective decision-making and efficient computation. Our algorithms demonstrate better regret than the first-order method and obtain substantial computational speed-ups compared to the LP-based method. Our algorithms exhibit robust scalability and adaptability under various re-solving frequencies and stochastic input models, consistently maintaining superior performance as the problem size grows. This adaptability ensures their effectiveness in diverse scenarios.
5 Conclusion
This paper addresses an online linear programming problem for decision-making in resource allocation. Identifying existing methods’ limitations, we develop a novel two-path framework that decouples online learning and decision-making into independent processes. Our cohesive framework integrates the LP-based method and first-order method, thereby balancing decision-making optimality and computational efficiency. We prove that our algorithm attains $\mathcal{O}(\log(T/f)+\sqrt{f})$ worst-case regret under continuous distributions, and our experimental results demonstrate strong regret and runtime improvements over competitive baselines.
Appendix
[…]
Structure of the Appendix
We organize the appendix as follows. In Section A, we introduce the foundation analysis for LP-based method and first-order method for the discussion of our framework; Section B presents the cohesive mathematical framework which unify all the OLP methods, and the unique properties derived from our parallel multi-phase algorithm; Section C combines all the previous results and proves the main results in our paper; Section D provides some auxiliary results and technical support in the previous three parts.
Appendix A Primary Results and Properties
In this section, we provide primary results of LP-based method and first-order method. We also include some basic analysis of dual algorithms that will help present our main results. We start from the convergence of regular LP-based method. All the analysis are under Assumption (1) and (2).
A.1 LP-based Analysis
We provide the analysis for dual solution obtained from LP-based method, so $\bm{p}_{t}=\bm{p}_{t}^{L}$ in this part.
Lemma A.1 (Boundedness of LP result).
The online dual price $\bm{p}_{t}$ from Algorithm 1 and the optimal dual price $\bm{p}^{*}$ of stochastic program (3) are bounded as
$\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}},\ \ \ \lVert\bm{p}_{t}\rVert\leq\frac{\bar{r}}{\underline{d}-\delta}.$
Proof. By the optimality of $\bm{p}^{*}$ and boundedness of $\bm{d}$, we have
$\underline{d}\lVert\bm{p}^{*}\rVert_{1}\leq\bm{d}^{\top}\bm{p}^{*}\leq\mathbb{E}[r]\leq\bar{r}.$
This holds because if otherwise, $\bm{p}^{*}$ can not be the optimal solution due to $f(\bm{p}^{*})>f(\bm{0})$. Given the non-negativeness of $\bm{p}^{*}$ and $\bm{p}^{*}\in\mathcal{V}_{\bm{d}}$, we obtain $\lVert\bm{p}^{*}\rVert\leq\lVert\bm{p}^{*}\rVert_{1}$, and thus $\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}}$.
Similarly, by the optimality of $\bm{p}_{t}^{*}$ and its associated $\bm{d}_{t}\in\mathcal{D}$ in Lemma 2.1, we know
$(\underline{d}-\delta)\lVert\bm{p}_{t}\rVert_{1}\leq{\bm{d}}_{t}^{\top}\bm{p}_{t}\leq\mathbb{E}[r]\leq\bar{r},$
so we get the bound as $\lVert\bm{p}_{t}\rVert\leq\frac{\bar{r}}{\underline{d}-\delta}$.
Proof of Lemma 3.2.
According to the latest results in [13], the online dual price $\bm{p}_{t}$ achieves a sublinear convergence $\mathcal{O}(\frac{1}{\sqrt{t}})$ to $\bm{p}_{t}^{*}$. Since
$\bm{p}^{*}\in\arg\min_{\bm{p}\geq\bm{0}}f(\bm{p}):=\bm{d}^{\top}\bm{p}+\mathbb{E}\left[(r-\bm{a}^{\top}\bm{p})^{+}\right],$
$\bm{p}_{t}^{*}\in\arg\min_{\bm{p}_{t}\geq\bm{0}}f_{\bm{d}_{t}}(\bm{p}_{t}):=\bm{d}_{t}^{\top}\bm{p}_{t}+\mathbb{E}\left[(r-\bm{a}^{\top}\bm{p}_{t})^{+}\right],$
By Lemma 12 in [16], we have $\|\bm{p}_{t}^{*}-\bm{p}^{*}\|_{2}^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\|\bm{d}_{t}-\bm{d}\|_{2}^{2}$.
A.2 First-order Analysis
We provide the analysis for dual solution obtained from LP-based method, so $\bm{p}_{t}=\bm{p}_{t}^{D}$ in this part.
Lemma A.2 (Boundedness of first-order result).
The online dual price $\bm{p}_{t}$ from Algorithm 2 is bounded as
$\lVert\bm{p}_{t}\rVert\leq\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d}).$
Proof. As our initial choice $\bm{p}_{1}=\bm{0}$, according to Lemma 1 in [15], we get the above result. In addition, by Lemma B.1 in [8], we have
$\lVert\bm{p}_{t}\rVert\leq\frac{\bar{r}}{\underline{d}}+\frac{m(\bar{a}+\bar{d})^{2}\alpha_{t}}{2\underline{d}}+\alpha_{t}\sqrt{m}(\bar{a}+\bar{d}).$
Proof for Lemma 3.3.
Based on the updated rule in Algorithm 2, we derive:
$\lVert\bm{p}_{t+1}-\bm{p}^{*}\rVert^{2}$ $\leq\lVert\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t})-\bm{p}^{*}\rVert^{2}$
$=\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}-2\alpha_{t}\langle\bm{d}-\bm{a}_{t}x_{t},\bm{p}_{t}-\bm{p}^{*}\rangle+\alpha_{t}^{2}\lVert\bm{d}-\bm{a}_{t}x_{t}\rVert^{2}$
$\leq\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}-2\alpha_{t}\langle\bm{d}-\bm{a}_{t}x_{t},\bm{p}_{t}-\bm{p}^{*}\rangle+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}.$
1) With convexity of $f$ and $\mathbb{E}[\bm{d}-\bm{a}_{t}x_{t}]\in\partial f(\bm{p}_{t})$, we have:
$f(\bm{p}^{*})\geq f(\bm{p}_{t})+\langle\bm{d}-\bm{a}_{t}x_{t},\bm{p}_{t}-\bm{p}^{*}\rangle.$
2) With quadratic regularity of $f$ as in Lemma 2.2, we have:
$f(\bm{p}_{t})\geq f(\bm{p}^{*})+\nabla f(\bm{p}^{*})^{\top}(\bm{p}_{t}-\bm{p}^{*})+\frac{\nu\lambda}{2}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2},$
which indicates $f(\bm{p}_{t})-f(\bm{p}^{*})\geq\frac{\nu\lambda}{2}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}.$
Combine the above results and take expectation conditioned on history information $\{(r_{j},\bm{a}_{j}),j\leq t\}$, we obtain:
$\mathbb{E}\lVert\bm{p}_{t+1}-\bm{p}^{*}\rVert^{2}$ $\leq\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}-2\alpha_{t}\mathbb{E}[\langle\bm{d}-\bm{a}_{t}x_{t},\bm{p}_{t}-\bm{p}^{*}\rangle]+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}$
$\leq\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}-2\alpha_{t}(f(\bm{p}_{t})-f(\bm{p}^{*}))+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}$
$\leq\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}-\alpha_{t}\nu\lambda\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}$
$=(1-\alpha_{t}\nu\lambda)\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}.$
This proves (18) with the general case of learning rate $\alpha_{t}$.
Case 1. When $\alpha_{t}=\alpha<\frac{1}{\nu\lambda}$ is a constant, we take recursion of (18). Note that $(1-\nu\lambda\alpha)^{t}<1/\nu\lambda\alpha t$, we get:
$\mathbb{E}\lVert\bm{p}_{t+1}-\bm{p}^{*}\rVert^{2}$ $\leq(1-\alpha\nu\lambda)^{t}\lVert\bm{p}_{1}-\bm{p}^{*}\rVert^{2}+\sum_{j=0}^{t-1}\alpha^{2}m(\bar{a}+\bar{d})^{2}(1-\alpha\nu\lambda)^{j}$
$\leq\frac{\lVert\bm{p}_{1}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha t}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}\alpha.$
Since all feasible $\bm{p}\in\mathcal{V}_{\bm{p}}$ is bounded, let $C_{fo}=\frac{\bar{p}^{2}+m(\bar{a}+\bar{d})^{2}}{\nu\lambda}$, we are able to obtain (19).
Case 2. When $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, for any $j\leq t$, Lemma 3.3 gives us
$\mathbb{E}\lVert\bm{p}_{j+1}-\bm{p}^{*}\rVert^{2}\leq\frac{j-1}{j+1}\lVert\bm{p}_{j}-\bm{p}^{*}\rVert^{2}+\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}(j+1)^{2}}.$
Re-arranging the above inequality, we have
$(j+1)^{2}\mathbb{E}\lVert\bm{p}_{j+1}-\bm{p}^{*}\rVert^{2}\leq(j^{2}-1)\lVert\bm{p}_{j}-\bm{p}^{*}\rVert^{2}+\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}},$
then $(j+1)^{2}\mathbb{E}\lVert\bm{p}_{j+1}-\bm{p}^{*}\rVert^{2}-j^{2}\lVert\bm{p}_{j}-\bm{p}^{*}\rVert^{2}\leq\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}}.$
Then by telescoping from $j=2$ to $t$, we have
$\sum_{j=1}^{t}(j+1)^{2}\mathbb{E}\lVert\bm{p}_{j+1}-\bm{p}^{*}\rVert^{2}-j^{2}\lVert\bm{p}_{j}-\bm{p}^{*}\rVert^{2}=(t+1)^{2}\mathbb{E}\lVert\bm{p}_{t+1}-\bm{p}^{*}\rVert^{2}-\lVert\bm{p}_{1}-\bm{p}^{*}\rVert^{2}\leq\frac{4m(\bar{a}+\bar{d})^{2}t}{\nu^{2}\lambda^{2}}$
which then gives us
$\mathbb{E}\lVert\bm{p}_{t+1}-\bm{p}^{*}\rVert^{2}\leq\frac{\lVert\bm{p}_{1}-\bm{p}^{*}\rVert^{2}}{(t+1)^{2}}+\frac{4m(\bar{a}+\bar{d})^{2}t}{\nu^{2}\lambda^{2}(t+1)^{2}},$
thus $\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\leq\frac{\lVert\bm{p}_{1}-\bm{p}^{*}\rVert^{2}}{t^{2}}+\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}t}.$
As $\bm{p}_{1}=\bm{0}$ and $\bm{p}^{*}$ is bounded, taking $C_{fo}=\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}}$ completes the proof for (20).
Appendix B Algorithm Design and Analysis
In this section, we present the cohesive mathematical framework to integrate the LP-based method and first-order method by developing a new performance metric. We also provide some unique properties for our algorithm design of parallel multi-phase structure.
As stated in Algorithm 3, we have independent two paths for online learning and online decision-making conducted simultaneously. Specifically, for the online learning path, we use the LP-based method to resolve the updated linear programs at a fixed frequency and update this result to the decision-making path; For online decision-making path, we restart this process every time when receiving an updated learning result. We apply first-order method in the initial and final batches, and use the latest dual price for decision-making in the intermediate resolving intervals. This process is illustrated in Figure (1).
B.1 Optimality Gap Analysis
First, we construct an upper bound for the offline optimal objective value. The challenge comes from the intractable dependency of constraints on objective value under the online setting. We tackle this issue by introducing a Lagrangian function to integrate constraints into the objective and balance revenue maximization with constraint satisfaction. The formulation is stated as follows.
Lemma B.1 (Lagrangian Upper Bound).
Under Assumptions (1) and (2), define the deterministic Lagrangian dual function as
$\ell(\bm{p})\coloneqq\mathbb{E}\left[rI(r>\bm{a}^{\top}\bm{p})+(\bm{d}-\bm{a}I(r>\bm{a}^{\top}\bm{p}))^{\top}\bm{p}^{*}\right].$
Then for any feasible $\bm{p}\in\mathcal{V}_{\bm{p}}$, we have:
$(a)\quad$ $\mathbb{E}\bigg{[}\sum_{t=1}^{T}r_{t}x_{t}^{*}\bigg{]}\leq T\ell(\bm{p}^{*}),$
$(b)\quad$ $\ell(\bm{p}^{*})-\ell(\bm{p})\leq\mu\bar{a}^{2}\lVert\bm{p}-\bm{p}^{*}\rVert^{2}.$ (27)
Proof. This lemma is proved as Lemma 3 in [16] by strong duality and optimality of $\bm{p}_{T}^{*}$.
Theorem B.1 (Decomposition of Optimality Gap).
Under Assumptions (1) and (2), let $C_{r}=\max\{\frac{\bar{r}}{\underline{d}},\frac{\nu\lambda\bar{a}^{2}}{2}\}$, we derive an upper bound for the optimality gap $r(\bm{x})$ as:
$r(\bm{x})\leq C_{r}\bigg{[}\mathbb{E}\Big{\lVert}\Big{(}\bm{b}-\sum_{t=1}^{T}\bm{a}_{t}x_{t}\Big{)}^{B^{+}}\Big{\rVert}+\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\bigg{]}.$ (28)
Proof. By the definition of $\ell(\bm{p})$, we derive the online objective as
$\mathbb{E}\bigg{[}\sum_{t=1}^{T}r_{t}x_{t}\bigg{]}$ $=\sum_{t=1}^{T}\mathbb{E}[\mathbb{E}[r_{t}x_{t}|\bm{p}_{t}]]$
$=\sum_{t=1}^{T}\mathbb{E}[\ell(\bm{p}_{t})-(\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}^{*}].$
Then by Lemma B.1, we derive the upper bound for the optimality gap $r(\bm{x})$ as defined in (8) to be:
$r(\bm{x})$ $=\mathbb{E}\bigg{[}\sum_{t=1}^{T}r_{t}x_{t}^{*}-r_{t}x_{t}\bigg{]}$
$\leq T\ell(\bm{p}^{*})-\sum_{t=1}^{T}\mathbb{E}[\ell(\bm{p}_{t})-(\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}^{*}]$
$=\sum_{t=1}^{T}\mathbb{E}[(\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}^{*}]+\sum_{t=1}^{T}\mathbb{E}[\ell(\bm{p}^{*})-\ell(\bm{p}_{t})]$
$\leq\mathbb{E}\bigg{[}\Big{(}\bm{b}-\sum_{t=1}^{T}\bm{a}_{t}x_{t}\Big{)}^{\top}\bm{p}^{*}\bigg{]}+\frac{\mu\bar{a}^{2}}{2}\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}$
$\leq\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\bigg{\lVert}\Big{(}\bm{b}-\sum_{t=1}^{T}\bm{a}_{t}x_{t}\Big{)}^{B^{+}}\bigg{\rVert}+\frac{\mu\bar{a}^{2}}{2}\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}$
where $(\cdot)^{B+}$ denotes the positive part only for binding constraints.
As $\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}}$ is bounded, taking the constant $C_{r}=\max\{\frac{\bar{r}}{\underline{d}},\frac{\mu\bar{a}^{2}}{2}\}$ completes the proof.
B.2 Total Regret Analysis
We consider the constraint violation as $v(\bm{x})=\lVert(\bm{Ax}-\bm{b})^{+}\rVert$ in (9). Then the total regret is the combination of optimality gap and constraint violation. Since Theorem B.1 holds for any unknown distribution $\mathcal{P}\in\Xi$, we have:
$\Delta_{T}$ $=\sup_{\mathcal{P}\in\Xi}\mathbb{E}_{\mathcal{P}}[r(\bm{x})+v(\bm{x})]$
$\leq C_{r}\bigg{[}\mathbb{E}\Big{\lVert}\Big{(}\bm{b}-\sum_{t=1}^{T}\bm{a}_{t}x_{t}\Big{)}^{B^{+}}\Big{\rVert}+\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\bigg{]}+\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=1}^{T}\bm{a}_{t}x_{t}-\bm{b}\Big{)}^{+}\Big{\rVert}$
$\leq C_{r}\bigg{[}\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\mathbb{E}\big{\lVert}(\bm{b}-\bm{Ax})^{B^{+}}\big{\rVert}+\mathbb{E}\big{\lVert}(\bm{Ax}-\bm{b})^{+}\big{\rVert}\bigg{]}.$ (29)
This new framework is adaptive to all OLP methods. We will proceed based on this structure to analyze various methods respectively.
Theorem B.2 (Horizon Division).
Based on the change of methods, we separate the horizon into three intervals $T_{1},T_{2}$, and $T_{3}$, and reorganize the total regret for analysis to be
$\Delta_{T}\leq C_{r}\big{[}\Delta_{T_{1}}+\Delta_{T_{2}}+\Delta_{T_{3}}\big{]}.$ (30)
where $\Delta_{T_{1}},\Delta_{T_{2}}$, and $\Delta_{T_{3}}$ are the total regret during each interval.
Proof. We split the total horizon into three intervals of initial batch $T_{1}=[0,f]$, intermediate process $T_{2}=[f,kf]$, and final batch $T_{3}=[kf,T]$. We use the first-order method for the first and final batches and LP-based method for intermediate process. Notice that for the initial resolving batch, we maintain the classical analysis for first-order method since it has not reached to the mixture with LP-based method. Then based on (B.2), we define
$\Delta_{T_{1}}$ $\coloneqq\sum_{t=1}^{f}\mathbb{E}[r_{t}x_{t}^{*}-r_{t}x_{t}]+\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=1}^{f}\bm{a}_{t}x_{t}-f\bm{d}\Big{)}^{+}\Big{\rVert}$
$\Delta_{T_{2}}$ $\coloneqq\sum_{t=f}^{kf}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\mathbb{E}\Big{\lVert}\Big{(}(k-1)f\bm{d}-\sum_{t=f}^{kf}\bm{a}_{t}x_{t}\Big{)}^{B^{+}}\Big{\rVert}+\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=f}^{kf}\bm{a}_{t}x_{t}-(k-1)f\bm{d}\Big{)}^{+}\Big{\rVert}$ (31)
$\Delta_{T_{3}}$ $\coloneqq\sum_{t=kf}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\mathbb{E}\Big{\lVert}\Big{(}(T-kf)\bm{d}-\sum_{t=kf}^{T}\bm{a}_{t}x_{t}\Big{)}^{B^{+}}\Big{\rVert}+\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=kf}^{T}\bm{a}_{t}x_{t}-(T-kf)\bm{d}\Big{)}^{+}\Big{\rVert}$
Then we have
$\Delta_{T}\leq C_{r}\big{[}\Delta_{T_{1}}+\Delta_{T_{2}}+\Delta_{T_{3}}\big{]}$
since $\lVert(a+b)^{+}\rVert\leq\lVert a^{+}\rVert+\lVert b^{+}\rVert$ for any $a,b$.
B.3 Regret for LP-based Method
In this section, we are going to bound $\Delta_{T_{2}}$ for the regret from LP-based method. To provide a tighter analysis, we examine the real-time average resource capacity $\bm{d}_{t}$ instead of the original process $\bm{b}_{t}$. Specifically, by Lemma 2.1, we constraint $\bm{d}_{t}$ within a feasible range $\mathcal{D}$. If $\bm{d}_{t}\notin\mathcal{D}$, we set the dual price to zero to accept all subsequent orders. We analyze the remaining resource and constraint violation due to this strategy. Note that this is required solely for rigor analysis purpose and is not necessary when running the algorithm.
Lemma B.2 (Dynamics of Resource Usage).
Under Assumption 1 and 2, there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda_{1},\mu$, and $C_{lp}$ such that
$\sum_{t=f}^{kf}\mathbb{E}[(d_{i,t}-d_{i})^{2}]\leq C\log(k).$ (32)
Proof. During the re-solving process, we denote each interval to be $[jf,(j+1)f]$ where $j=1,2,...,k$. As the resource usage follows $\bm{b}_{t+1}=\bm{b}_{t}-\bm{a}_{t+1}I(r_{t+1}>\bm{a}_{t+1}^{T}\bm{p}_{{t+1}})$, normalizing both sides, we derive the update of average resource consumption to be:
$d_{i,{(j+1)f}}=d_{i,jf}+\frac{\sum_{\ell=jf+1}^{(j+1)f}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(k+1)f})}{T-(j+1)f}.$
Subtracting $d_{i}$ on both sides, it becomes
$d_{i,(j+1)f}-d_{i}$ $={d}_{i,jf}-{d}_{i}+\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}$
$={d}_{i,jf}-{d}_{i}+\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}$
$\ \ +\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}.$
Taking expectations of squares, we have
$\mathbb{E}(d_{i,(j+1)f}-d_{i})^{2}=$ $\mathbb{E}(d_{i,jf}-d_{i})^{2}+\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}}{(T-(j+1)f)^{2}}\right]$
$+$ $\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}))^{2}}{(T-(j+1)f)^{2}}\right]$
$+2$ $\mathbb{E}\left[(d_{i,jf}-d_{i})\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\right)\right]$
$+2$ $\mathbb{E}\left[(d_{i,jf}-d_{i})\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}\right)\right]$
$+2$ $\mathbb{E}\Bigg{[}\frac{\sum_{\ell=jf+1}^{(j+1)f}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\cdot\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}\Bigg{]}.$ (33)
By Lemma D.1, we obtain the recursion relation as
$\mathbb{E}(d_{i,(j+1)f}-d_{i})^{2}$ $\leq\mathbb{E}({d}_{i,jf}-d_{i})^{2}+\frac{C_{rec}}{(k-j-1)^{2}f}+\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-d_{i})^{2}]}$ (34)
where $C_{rec}>0$ is a constant defined in Lemma D.1.
Then according to Lemma D.2, take $C=12\max\{C_{rec},16\mu^{2}\bar{a}^{4}C_{lp}\}$, we solve recursion (34) and obtain the upper bound of total deviation from original $d_{i}$ in the re-solving process to be:
$\sum_{j=1}^{k}\mathbb{E}\left[({d}_{i,jf}-{d}_{i})^{2}\right]$ $\leq\frac{C}{f}\log(k).$ (35)
Therefore, we sum the whole re-solving process and obtain:
$\sum_{t=f}^{kf}\mathbb{E}[(d_{i,t}-d_{i})^{2}]$ $=\sum_{j=1}^{k}\sum_{\ell=jf}^{(j+1)f}\mathbb{E}[(d_{i,\ell}-d_{i})^{2}]$
[…]
$\leq C\log(k).$ (36)
This completes the proof.
Theorem B.3 (Regret of Intermediate Process).
Under Assumption 1 and 2, following the result in (B.2), we prove the regret satisfies
$\Delta_{T_{2}}\leq\log\Big{(}\frac{T}{f}\Big{)}=\mathcal{O}(\log(k)).$ (37)
Proof. We analyze the three terms in (B.2) respectively.
  1. 1.
    For dual convergence, by Lemma 3.2, we have
    $\sum_{t=f}^{kf}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}$ $=\mathbb{E}\left[\sum_{j=1}^{k-1}\sum_{t=jf+1}^{(j+1)f}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert_{2}^{2}\right]$
    $=\sum_{j=1}^{k-1}f\ \mathbb{E}\Big{[}\lVert\bm{p}_{(j+1)f}-\bm{p}^{*}\rVert_{2}^{2}\Big{]}$
    $\leq\sum_{j=1}^{k-1}f\ \mathbb{E}\Big{[}\lVert\bm{p}_{(j+1)f}-\bm{p}_{jf}^{*}\rVert_{2}^{2}+\lVert\bm{p}_{jf}^{*}-\bm{p}^{*}\rVert_{2}^{2}\Big{]}$
[…]
  2. 2.
    For the constraint violation, by Chebyshev’s inequality, we bound the probability as
    $\sum_{t=f}^{kf}\mathbb{P}(|d_{i,t}-d_{i}|\leq\delta)$ $\leq\sum_{t=f}^{kf}\frac{\mathbb{E}[(d_{i,t}-d_{i})^{2}]}{\delta^{2}}$
[…]
    Since our strategy is to automatically set $\bm{p}_{t}=0$ once $\bm{d}_{t}\notin\mathcal{D}$, we will accept all the subsequent orders according to our decision condition (6). Denote $R$ as the total number of orders processed specifically due to this rule, then we have:
    $\mathbb{E}[R]\leq\sum_{t=f}^{kf}\mathbb{P}(|d_{i,t}-d_{i}|\leq\delta)\leq\frac{C}{\delta^{2}}\log(k)$ (40)
    which indicates that the resource violation is at most
    $\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=f}^{kf}\bm{a}_{t}x_{t}-(k-1)f\bm{d}\Big{)}^{+}\Big{\rVert}\leq\mathbb{E}[\bar{a}R]\leq\frac{C\bar{a}}{\delta^{2}}\log(k).$ (41)
  3. 3.
    For the remaining resource, we derive its upper bound by considering the opposite extreme case of our strategy. If we reject all those $R$ orders, then no resource is used on them. Thus, by (40), the positive projection of remaining resource at $t=kf$ is at most:
    $\mathbb{E}\Big{\lVert}\Big{(}(k-1)f\bm{d}-\sum_{t=f}^{kf}\bm{a}_{t}x_{t}\Big{)}^{B^{+}}\Big{\rVert}\leq\mathbb{E}[(d_{i}+\delta)R]\leq\frac{C(\bar{d}+\delta)}{\delta^{2}}\log(k).$ (42)
Combining the results of (1), (41), and (42), we derive the final result as
$\Delta_{T_{2}}\leq\Big{(}C_{lp}+\frac{mC}{\nu^{2}\lambda^{2}}+(\bar{a}+\bar{d}+\delta)\frac{C}{\delta^{2}}\Big{)}\log(k).$ (43)
This completes the proof.
B.4 Regret for First-Order Method
In this section, we are going to bound $\Delta_{T_{1}}$ and $\Delta_{T_{3}}$ for the regret from first-order method. The key point is to analyze the connection between dual prices based on the gradient-updated rule. Here $\bm{p}_{t}=\bm{p}_{t}^{fo}$.
Theorem B.4 (Regret of Initial Batch).
Under Assumption 1 and 2, we have
$\Delta_{T_{1}}\leq\left(\frac{m(\bar{a}+\bar{d})^{2}}{2}+\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right)\sqrt{f}.$ (44)
Proof. We analyze $\Delta_{T_{1}}$ in (B.2) by parts. By Lemma B.3, B.4 in [8], we have
$\sum_{t=1}^{f}\mathbb{E}[r_{t}x_{t}^{*}-r_{t}x_{t}]\leq\frac{m(\bar{a}+\bar{d})^{2}}{2}\alpha_{t}f,$
$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=1}^{f}\bm{a}_{t}x_{t}-f\bm{d}\Big{)}^{+}\Big{\rVert}\leq\frac{1}{\alpha_{t}}\bigg{[}\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\bigg{]}.$
To achieve the tight upper bound, we select the optimal step size $\alpha_{t}=\frac{1}{\sqrt{f}}$, which then gives us the desired result.
Lemma B.3 (Constraint Violation).
Under Assumption 1 and 2, take the learning rate $\alpha_{t}=\alpha$, we bound the constraint violation for the last batch as
$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=kf}^{T}\bm{a}_{t}x_{t}-(T-kf)\bm{d}\Big{)}^{+}\Big{\rVert}\leq\Big{(}\frac{C_{lp}}{\sqrt{kf}}+\frac{C_{b}}{T^{2}}\Big{)}\cdot\frac{1}{\alpha}+\frac{C_{b}}{\sqrt{(T-kf)kf}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+\frac{C_{b}}{\sqrt{\alpha}}.$ (45)
Proof. With the update rule of first-order method in Algorithm 2, we have
$\bm{p}_{t+1}=[\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t})]^{+}\geq\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t}),$
which gives us $\bm{a}_{t}x_{t}-\bm{d}\leq\frac{1}{\alpha}(\bm{p}_{t+1}-\bm{p}_{t}).$
Summarizing on both sides and apply the telescoping, we derive
$\sum_{t=kf}^{T}(\bm{a}_{t}x_{t}-\bm{d})\leq\frac{1}{\alpha}\sum_{t=kf}^{T}(\bm{p}_{t+1}-\bm{p}_{t})=\frac{1}{\alpha}(\bm{p}_{T+1}-\bm{p}_{kf}).$
By Lemma D.3, take $C_{b}=\max\{\frac{C_{lp}}{\sqrt{\nu\lambda}},\frac{m(\bar{a}+\bar{d})}{\sqrt{\nu\lambda}},\bar{p}\}$, we have
$\mathbb{E}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert\leq C_{b}\Big{[}\frac{1}{\sqrt{(T-kf)kf}}\cdot\frac{1}{\sqrt{\alpha}}+\sqrt{\alpha}+\frac{1}{T^{2}}\Big{]}.$ (46)
Thus, according to Lemma 3.2 and (46), the expectation of constraint violation satisfies
$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=kf}^{T}(\bm{a}_{t}x_{t}-\bm{d})\Big{)}^{+}\Big{\rVert}$ $\leq\mathbb{E}\Big{\lVert}\sum_{t=kf}^{T}(\bm{a}_{t}x_{t}-\bm{d})\Big{\rVert}$
$\leq\frac{1}{\alpha}\mathbb{E}\lVert\bm{p}_{T+1}-\bm{p}_{kf}\rVert$
$\leq\frac{1}{\alpha}\mathbb{E}\Big{[}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert+\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert\Big{]}$
$\leq\frac{1}{\alpha}\Big{(}\frac{C_{lp}}{\sqrt{kf}}+C_{b}\Big{[}\frac{1}{\sqrt{(T-kf)kf}}\cdot\frac{1}{\sqrt{\alpha}}+\sqrt{\alpha}+\frac{1}{T^{2}}\Big{]}\Big{)}$
$\leq\frac{C_{lp}}{\sqrt{kf}}\cdot\frac{1}{\alpha}+\frac{C_{b}}{\sqrt{(T-kf)kf}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+C_{b}\cdot\frac{1}{\sqrt{\alpha}}+\frac{C_{b}}{T^{2}}\cdot\frac{1}{\alpha}.$ (47)
This completes the proof.
Theorem B.5 (Regret of Final Batch).
Under Assumption 1 and 2, we have
$\Delta_{T_{3}}\leq\bigg{(}\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+\frac{2C_{lp}}{\sqrt{k}}+\frac{C_{lp}\log(f)}{\nu\lambda k}+\frac{4C_{b}}{\sqrt{k}}+2C_{b}\bigg{)}\cdot f^{1/3}.$ (48)
Proof. We decompose $\Delta_{T_{3}}$ into three parts according to (B.2) and analyze each term respectively. Since the first-order method re-starts from $\bm{p}_{kf}$, by Lemma 3.2, we have:
$\sum_{t=kf}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}$ $\leq\sum_{t=kf}^{T}\Big{[}\frac{\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha t}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}\alpha\Big{]}$
$\leq\frac{\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha}\log(T-kf)+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}(T-kf)\alpha$
$\leq\frac{\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha}\log(f)+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha$
$\leq\frac{C_{lp}\log(f)}{\nu\lambda kf}\cdot\frac{1}{\alpha}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha.$ (49)
Since $t\in[kf,T]$, by Proposition 3.3 in [8], the dual price is far from the origin and thus the positive projection still equals to itself. This indicates that:
$\bm{p}_{t+1}=[\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t})]^{+}=\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t}),$
which then gives us $\bm{d}-\bm{a}_{t}x_{t}=\frac{1}{\alpha}(\bm{p}_{t}-\bm{p}_{t+1}).$
Summarizing on both sides and apply the telescoping, we derive
$\sum_{t=kf}^{T}(\bm{d}-\bm{a}_{t}x_{t})\leq\frac{1}{\alpha}\sum_{t=kf}^{T}(\bm{p}_{t}-\bm{p}_{t+1})=\frac{1}{\alpha}(\bm{p}_{kf}-\bm{p}_{T+1}).$
Then for the positive binding terms, by Lemma 3.2 and (46), we have
$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=kf}^{T}(\bm{d}-\bm{a}_{t}x_{t})\Big{)}^{B^{+}}\Big{\rVert}$ $\leq\mathbb{E}\Big{\lVert}\sum_{t=kf}^{T}(\bm{d}-\bm{a}_{t}x_{t})\Big{\rVert}$
$\leq\frac{1}{\alpha}\mathbb{E}\lVert\bm{p}_{kf}-\bm{p}_{T+1}\rVert$
$\leq\frac{1}{\alpha}\mathbb{E}\Big{[}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert+\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert\Big{]}$
$\leq\frac{C_{lp}}{\sqrt{kf}}\cdot\frac{1}{\alpha}+\frac{C_{b}}{\sqrt{(T-kf)kf}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+C_{b}\cdot\frac{1}{\sqrt{\alpha}}+\frac{C_{b}}{T^{2}}\cdot\frac{1}{\alpha}.$ (50)
Combining the results of (B.4), Lemma B.3, and (B.4) together, we obtain the final result as:
$\Delta_{T_{3}}\leq\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha+\Big{(}\frac{2C_{lp}}{\sqrt{kf}}+\frac{2C_{b}}{T^{2}}+\frac{C_{lp}\log(f)}{\nu\lambda kf}\Big{)}\cdot\frac{1}{\alpha}+\frac{2C_{b}}{\sqrt{(T-kf)kf}}\cdot\frac{2}{\alpha\sqrt{\alpha}}+\frac{2C_{b}}{\sqrt{\alpha}}.$ (51)
We select the optimal learning rate $\alpha=f^{-2/3}$ to minimize $\Delta_{T_{3}}$ in (51). We prove this result in cases.
  1. 1.
    If $T-kf\leq f^{1/3}$, the total regret must be smaller than the length of the batch, so $\Delta_{T_{3}}\leq T-kf\leq f^{1/3}$.
  2. 2.
    If $T-kf>f^{1/3}$, then we use this property to bound the term $\frac{2C_{b}}{\sqrt{(T-kf)kf}}\cdot\frac{2}{\alpha\sqrt{\alpha}}$. We derive
    $\Delta_{T_{3}}$ $\leq\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\cdot f^{-2/3}+\Big{(}\frac{2C_{lp}}{\sqrt{kf}}+\frac{2C_{b}}{T^{2}}+\frac{C_{lp}\log(f)}{\nu\lambda kf}\Big{)}\cdot f^{2/3}+\frac{4C_{b}}{\sqrt{f^{1/3}\cdot kf}}\cdot f+2C_{b}\cdot f^{1/3}$
    $=\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f^{1/3}+\frac{2C_{lp}}{\sqrt{k}}f^{1/6}+\frac{2C_{b}}{T}+\frac{C_{lp}\log(f)}{\nu\lambda k}f^{1/3}+\frac{4C_{b}}{\sqrt{k}}\cdot f^{1/3}+2C_{b}\cdot f^{1/3}$
    $\leq\Big{(}\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+\frac{2C_{lp}}{\sqrt{k}}+\frac{C_{lp}\log(f)}{\nu\lambda k}+\frac{4C_{b}}{\sqrt{k}}+2C_{b}\Big{)}\cdot f^{1/3}.$ (52)
This completes the proof.
Lemma B.4 (Warm-Start First-Order Regret).
Assume that we have a warm start for the initial dual price $\bm{p}_{0}$ in the first batch, which satisfies $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$. Then the regret of first batch follows
$\Delta_{T_{1}\text{warm}}\leq\Big{(}\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+4C_{w}\Big{)}\cdot f^{1/3}+\log(f).$ (53)
Proof. According to Lemma D.3, with $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ and $\alpha<\frac{1}{\nu\lambda}$, we have
$\mathbb{E}\big{[}\lVert\bm{p}_{f+1}-\bm{p}^{*}\rVert^{2}\mid\bm{p}_{0}\big{]}\leq\frac{\lVert\bm{p}_{0}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha f}+\frac{\alpha m(\bar{a}+\bar{d})^{2}}{\nu\lambda},$
$\mathbb{E}\lVert\bm{p}_{f+1}-\bm{p}^{*}\rVert\leq\frac{1}{f^{1/3}\sqrt{\nu\lambda f}}\cdot\frac{1}{\sqrt{\alpha}}+\frac{m(\bar{a}+\bar{d})}{\sqrt{\nu\lambda}}\cdot\sqrt{\alpha}+\frac{\bar{p}}{T^{2}}.$ (54)
Then according to Lemma B.3, take $C_{w}=\max\{\frac{1}{\nu\lambda},\frac{m(\bar{a}+\bar{d})}{\sqrt{\nu\lambda}},\bar{p}\}$, we derive the constraint violation follows
$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=0}^{f}(\bm{a}_{t}x_{t}-\bm{d})\Big{)}^{+}\Big{\rVert}$ $\leq\mathbb{E}\Big{\lVert}\sum_{t=0}^{f}(\bm{a}_{t}x_{t}-\bm{d})\Big{\rVert}$
$\leq\frac{1}{\alpha}\mathbb{E}\lVert\bm{p}_{f+1}-\bm{p}_{0}\rVert$
$\leq\frac{1}{\alpha}\mathbb{E}\Big{[}\lVert\bm{p}_{0}-\bm{p}^{*}\rVert+\lVert\bm{p}_{f+1}-\bm{p}^{*}\rVert\Big{]}$
$\leq\frac{1}{\alpha}\Big{(}\frac{1}{f^{1/3}}+C_{w}\Big{[}\frac{1}{f^{1/3}\sqrt{f}}\cdot\frac{1}{\sqrt{\alpha}}+\sqrt{\alpha}+\frac{1}{T^{2}}\Big{]}\Big{)}$
$\leq\frac{1}{f^{1/3}}\cdot\frac{1}{\alpha}+\frac{C_{w}}{f^{5/6}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+C_{w}\cdot\frac{1}{\sqrt{\alpha}}+\frac{C_{w}}{T^{2}}\cdot\frac{1}{\alpha}.$ (55)
Similar to (B.4) in Theorem B.5, we derive the dual distance as
$\sum_{t=0}^{f}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}$ $\leq\sum_{t=0}^{f}\Big{[}\frac{\lVert\bm{p}_{0}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha t}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}\alpha\Big{]}$
$\leq\frac{\lVert\bm{p}_{0}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha}\log(f)+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha$
$\leq\frac{\log(f)}{\nu\lambda f^{2/3}}\cdot\frac{1}{\alpha}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha.$ (56)
Similar to (B.4) in Theorem B.5, we derive the positive projection for binding terms as
$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=0}^{f}(\bm{d}-\bm{a}_{t}x_{t})\Big{)}^{B^{+}}\Big{\rVert}$ $\leq\mathbb{E}\Big{\lVert}\sum_{t=0}^{f}(\bm{d}-\bm{a}_{t}x_{t})\Big{\rVert}$
$\leq\frac{1}{\alpha}\mathbb{E}\lVert\bm{p}_{0}-\bm{p}_{f+1}\rVert$
$\leq\frac{1}{\alpha}\mathbb{E}\Big{[}\lVert\bm{p}_{0}-\bm{p}^{*}\rVert+\lVert\bm{p}_{f+1}-\bm{p}^{*}\rVert\Big{]}$
$\leq\frac{1}{f^{1/3}}\cdot\frac{1}{\alpha}+\frac{C_{w}}{f^{1/3}\sqrt{f}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+C_{w}\cdot\frac{1}{\sqrt{\alpha}}+\frac{C_{w}}{T^{2}}\cdot\frac{1}{\alpha}.$ (57)
Therefore, combining the results in (B.4), (B.4), and (B.4) together, we have
$\Delta_{T_{1}\text{warm}}$ $\leq\frac{\log(f)}{\nu\lambda f^{2/3}}\cdot\frac{1}{\alpha}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha$
$\quad+\frac{2}{f^{1/3}}\cdot\frac{1}{\alpha}+\frac{2C_{w}}{f^{1/3}\sqrt{f}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+2C_{w}\cdot\frac{1}{\sqrt{\alpha}}+\frac{2C_{w}}{T^{2}}\cdot\frac{1}{\alpha}$
$\leq\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha+\Big{(}\frac{\log(f)}{\nu\lambda f^{2/3}}+\frac{2}{f^{1/3}}+\frac{2C_{w}}{T^{2}}\Big{)}\cdot\frac{1}{\alpha}+\frac{2C_{w}}{f^{1/3}\sqrt{f}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+2C_{w}\cdot\frac{1}{\sqrt{\alpha}}.$ (58)
Thus, take the optimal learning rate $\alpha=f^{-2/3}$, we obtain the regret for the first batch with warm start as:
$\Delta_{T_{1}\text{warm}}$ $\leq\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f^{1/3}+\Big{(}\frac{\log(f)}{\nu\lambda f^{2/3}}+\frac{2}{f^{1/3}}+\frac{2C_{w}}{T^{2}}\Big{)}\cdot f^{2/3}+\frac{2C_{w}}{f^{1/3}\sqrt{f}}\cdot f+2C_{w}\cdot f^{1/3}$
$\leq\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f^{1/3}+\frac{\log(f)}{\nu\lambda}+2f^{1/3}+\frac{2C_{w}}{T}+2C_{w}f^{1/6}+2C_{w}\cdot f^{1/3}$
$\leq\Big{(}\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+4C_{w}\Big{)}\cdot f^{1/3}+\log(f).$ (59)
This completes the proof.
Appendix C Main Results
In this section, we demonstrate the proof for all theoretical results that we proposed in the main body of the paper. We instate Assumption 1 and 2 are satisfied.
C.1 Proof of Lemma 3.1
By the boundedness results in Lemma A.1 and Lemma A.2, define
$\bar{p}\coloneqq\max\left(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right),$
then we complete the proof.
C.2 Proof of Theorem 3.1
By Theorem B.1, we obtain
$\Delta_{T}$ $=\mathbb{E}[r(\bm{x})+v(\bm{x})]$
$\leq\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\bigg{\lVert}\Big{(}\bm{b}-\sum_{t=1}^{T}\bm{a}_{t}x_{t}\Big{)}^{B^{+}}\bigg{\rVert}+\frac{\mu\bar{a}^{2}}{2}\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=1}^{T}\bm{a}_{t}x_{t}-\bm{b}\Big{)}^{+}\Big{\rVert}$
$\leq\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\left[\lVert(\bm{b}-\bm{Ax})^{B+}\rVert\right]+\mu\bar{a}^{2}\sum_{t=1}^{T}\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]+\mathbb{E}\left[\lVert(\bm{Ax}-\bm{b})^{+}\rVert\right].$
This completes the proof.
C.3 Proof of Theorem 3.2
Combining the results of Theorem B.3, Theorem B.4, and Theorem B.5, for some constant $C_{reg}>0$, we obtain:
$\Delta_{T}$ $=\Delta_{T_{1}}+\Delta_{T_{2}}+\Delta_{T_{3}}$
$\leq\left(\frac{m(\bar{a}+\bar{d})^{2}}{2}+\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right)\sqrt{f}$
$\quad+\Big{(}C_{lp}+\frac{mC}{\nu^{2}\lambda^{2}}+(\bar{a}+\bar{d}+\delta)\frac{C}{\delta^{2}}\Big{)}\log(k)$
$\quad+\Big{(}\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+\frac{2C_{lp}}{\sqrt{k}}+\frac{C_{lp}\log(f)}{\nu\lambda k}+\frac{4C_{b}}{\sqrt{k}}+2C_{b}\Big{)}\cdot f^{1/3}.$ (60)
Therefore, we achieve the worst-case regret of:
$\Delta_{T}=\mathcal{O}(\log(k)+\sqrt{f}+f^{1/3}).$ (61)
For special cases,
  1. 1.
    If we use LP-based method on the first batch, then we will have $\Delta_{T_{1}}\leq\log(f)$. Total regret follows
    $\Delta_{T}\leq\log(\max\{f,k\})+f^{1/3})\leq\log(\sqrt{T})+f^{1/3}\in\mathcal{O}(\log(T)+f^{1/3}).$ (62)
  2. 2.
    If we have a warm start for the first batch, with the initialization $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$, then the first batch regret achieves $\Delta_{T_{1}}\in\mathcal{O}(f^{1/3}+\log(f))$ by Lemma B.4. Thus, the total regret follows
    $\Delta_{T}\leq\log(\max\{f,k\})+2f^{1/3})\in\mathcal{O}(\log(T)+f^{1/3}).$ (63)
This completes our proof.
Appendix D Auxiliary Results
In this section, we provide some auxiliary results to support the proof in the previous three sections. These lemmas focus on pure mathematical derivations.
D.1 Technical Support for LP-based Analysis
Lemma D.1.
Denote $d_{i,(j+1)f},d_{i,jf}$ as the average consumption of $i$-th type resource at time $(j+1)f$ and $jf$. There exists a constant $C_{rec}$ depending on $\bar{d},\bar{a},m,\nu,\lambda,\mu$, and $C_{lp}$ such that:
$\mathbb{E}(d_{i,(j+1)f}-d_{i})^{2}$ $\leq\mathbb{E}({d}_{i,jf}-d_{i})^{2}+\frac{C_{rec}}{(k-j-1)^{2}f}+\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-d_{i})^{2}]}.$
Proof. We analyze each term in (B.3). The key technique we used is to take conditional expectations and simplify the double summations.
  1. (a)
    Term 1.
    $\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}}{(T-(j+1)f)^{2}}\right]$
    =\ = $\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\sum_{s=jf+1}^{(j+1)f}\mathbb{E}\Big{[}(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))(d_{i,jf}-{a}_{i,s}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}))\mid\bm{d}_{jf}\Big{]}\right]$
    =\ = $\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big{[}(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}\mid\bm{d}_{jf}\Big{]}\right]$
    $+\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell\neq j,\ell=jf+1}^{(j+1)f}\mathbb{E}\Big{[}(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))(d_{i,jf}-{a}_{i,j}I(r_{j}>\bm{a}_{j}^{\top}\bm{p}_{jf}^{*}))\mid\bm{d}_{jf}\Big{]}\right]$
[…]
    =\ = $\frac{f(\bar{a}+\bar{d})^{2}}{(T-(j+1)f)^{2}}\ \ \ (\text{since }\mathbb{E}\Big{[}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})|\bm{d}_{jf}\Big{]}=0\text{ for binding terms})$
[…]
  2. (b)
    Term 2.
    $\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}))^{2}}{(T-(j+1)f)^{2}}\right]$
    =\ = $\frac{\mathbb{E}\Bigg{[}\sum_{\ell,s=jf+1}^{jk+1)f}\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))\cdot({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\Big{|}\bm{d}_{jf}\Big{]}\Bigg{]}}{(T-(j+1)f)^{2}}.$
    When $\ell=s$,
    $\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))^{2}\mid\bm{d}_{jf}\Big{]}\leq\bar{a}^{2}.$
    When $\ell\neq s$, by Assumption 2,
    $\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\mid\bm{d}_{jf}\Big{]}$
    =\ = $\mathbb{E}\Bigg{[}\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\mid\bm{a}_{i},\bm{a}_{s}\Big{]}\mid\bm{d}_{jf}\Bigg{]}$
    =\ = $\mathbb{E}\Bigg{[}{a}_{i,\ell}\mathbb{E}\Big{[}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big{]}\cdot{a}_{i,s}\mathbb{E}\Big{[}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{s}\Big{]}\mid\bm{d}_{jf}\Bigg{]}$
    =\ = $\mathbb{E}\Bigg{[}{a}_{i,\ell}(P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{\ell})-P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{\ell}))\cdot{a}_{i,s}(P(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{s})-P(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{s}))\mid\bm{d}_{jf}\Bigg{]}$
[…]
    Combining these two cases together and by the convergence of LP-based method, we obtain the bound for Term 2 as
    $\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}))^{2}}{(T-(j+1)f)^{2}}\right]$
[…]
  3. (c)
    Term 3.
    $2\mathbb{E}\left[({d}_{i,jf}-d_{i})\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\right)\right]$
    =\ = $\frac{2}{T-(j+1)f}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big{[}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})\mid\bm{d}_{jf}\Big{]}\cdot({d}_{i,jf}-d_{i})\right]$
[…]
  4. (d)
    Term 4.
    $2\mathbb{E}\left[({d}_{i,jf}-d_{i})\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}\right)\right]$
    =\ = $\frac{2}{T-(j+1)f}\ \mathbb{E}\Bigg{[}\sum_{\ell=jf+1}^{(j+1)f}({d}_{i,jf}-{d}_{i}){a}_{i,\ell}\ \mathbb{E}\Big{[}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big{]}\Bigg{]}$
[…]
    =\ = $\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}.$
  5. (e)
    Term 5.
    $2\mathbb{E}\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\cdot\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}I(jf<\tau)\right)$
[…]
    =\ = $\frac{2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\frac{1}{j}\mu^{2}\bar{a}^{2}C_{lp}}}{(k-j-1)^{2}f}.$
Combining all the terms, we obtain the upper bound as:
$\mathbb{E}\left[({d}_{i,(j+1)f}-{d}_{i})^{2}\right]$ $\leq\mathbb{E}({d}_{i,jf}-{d}_{i})^{2}+\frac{(\bar{a}+\bar{d})^{2}}{(k-j-1)^{2}f}+\frac{\bar{a}^{2}+\frac{1}{j}\mu^{2}\bar{a}^{4}C_{lp}}{(k-j-1)^{2}f}$
$\ \ +\frac{2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\frac{1}{j}\mu^{2}\bar{a}^{2}C_{lp}}}{(k-j-1)^{2}f}+\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}.$
Taking $C_{rec}=(\bar{a}+\bar{d})^{2}+\bar{a}^{2}+\mu^{2}\bar{a}^{4}C_{lp}+2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\mu^{2}\bar{a}^{2}C_{lp}}$ completes the proof.
Lemma D.2.
With the recursion relation in (34), there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda,\mu$, and $C_{lp}$ such that the summation of the total deviation of $\bm{d}_{t}$ with the original $\bm{d}$ satisfies:
$\sum_{j=1}^{k}\mathbb{E}\left[({d}_{i,jf}-{d}_{i})^{2}\right]$ $\leq\frac{C}{f}\log(k).$
Proof. We consider a general sequence $\{z_{j}\}_{j=1}^{k}$ with
$z_{j+1}\leq z_{j}+\frac{R}{(k-j-1)^{2}f}+\frac{\sqrt{R}\sqrt{z_{j}}}{(k-j-1)\sqrt{(j+1)f}}$
where $R>0$ is a constant.
Taking sum on both sides of the inequality and re-arranging, we have
$\sum_{j=1}^{k}(k-j+1)(z_{j+1}-z_{j})$ $\leq\sum_{j=1}^{k}\frac{16R}{(k-j+1)f}+\sqrt{\frac{16R}{f}}\sum_{j=1}^{k}\frac{\sqrt{z_{j}}}{\sqrt{j+1}}$
$\leq\frac{16R}{f}\log(k)+\sqrt{\frac{16R}{f}}\sum_{j=1}^{k}\frac{\sqrt{z_{j}}}{\sqrt{j+1}}.$
Noticing that $\sum_{j=1}^{k}(k-j+1)(z_{j+1}-z_{j})=\sum_{j=1}^{k}z_{j}$, we have
$\frac{16R}{f}\Big{(}\sum_{j=1}^{k}\frac{1}{j+1}\Big{)}\cdot\Big{(}\sum_{j=1}^{k}z_{j}\Big{)}$ $\geq\left(\sqrt{\frac{16R}{f}}\sum_{j=1}^{k}\frac{\sqrt{z_{j}}}{\sqrt{j+1}}\right)^{2}$
$\geq\bigg{(}\sum_{j=1}^{k}z_{j}-\frac{4R}{f}\log(k)\bigg{)}^{2}.$
We treat $\sum_{j=1}^{k}z_{j}$ as the variable, then solve and get
$\sum_{j=1}^{k}z_{j}\leq\frac{12R}{f}\log(k).$
With this result, consider our recursion in (34), take the constant $C=12\max\{C_{rec},16\mu^{2}\bar{a}^{4}C_{lp}\}>0$, we obtain:
$\sum_{j=1}^{k}\mathbb{E}\left[({d}_{i,jf}-{d}_{i})^{2}\right]$ $\leq\frac{C}{f}\log(k).$
D.2 Technical Support for First-order Analysis
Lemma D.3.
Following the updated rule of first-order method, we derive last dual price satisfies:
$\mathbb{E}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{\nu\lambda(T-kf)kf}}\cdot\frac{1}{\sqrt{\alpha}}+\frac{m(\bar{a}+\bar{d})}{\sqrt{\nu\lambda}}\cdot\sqrt{\alpha}+\frac{\bar{p}}{T^{2}}.$
Proof. According to Lemma 3.3, take $\alpha_{t}=\alpha<\frac{1}{\nu\lambda}$, we derive the conditional expectation as
$\mathbb{E}\big{[}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert^{2}\mid\bm{p}_{kf}\big{]}$ $\leq(1-\nu\lambda\alpha)\mathbb{E}\big{[}\lVert\bm{p}_{T}-\bm{p}^{*}\rVert^{2}\mid\bm{p}_{kf}\big{]}+\alpha^{2}m(\bar{a}+\bar{d})^{2}$
$\leq(1-\nu\lambda\alpha)^{T-kf}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert^{2}+\sum_{j=0}^{T-kf-1}\alpha^{2}m(\bar{a}+\bar{d})^{2}(1-\nu\lambda\alpha)^{j}$
$\leq(1-\nu\lambda\alpha)^{T-kf}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert^{2}+\frac{\alpha^{2}m(\bar{a}+\bar{d})^{2}}{\nu\lambda\alpha}$
$\leq\frac{1}{\nu\lambda\alpha(T-kf)}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert^{2}+\frac{\alpha m(\bar{a}+\bar{d})^{2}}{\nu\lambda}$ (64)
where we use the technique of $\sum_{j=0}^{T-kf-1}\alpha^{2}m(\bar{a}+\bar{d})^{2}(1-\nu\lambda\alpha)^{j}\leq\frac{1-(1-\nu\lambda\alpha)^{T-kf}}{\nu\lambda\alpha}\leq\frac{1}{\nu\lambda\alpha}$ and $(1-\nu\lambda\alpha)^{T-kf}\leq\frac{1}{1+\nu\lambda\alpha(T-kf)}\leq\frac{1}{\nu\lambda\alpha(T-kf)}$.
By LP-convergence result in Lemma 3.2, we know $\mathbb{E}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$. By Proposition 3.3 in [8], we know the event $E\coloneqq\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$ with probability $\mathbb{P}\geq 1-\frac{1}{T^{4}}$. By Lemma 3.1, we know $\lVert\bm{p}_{t}\rVert\leq\bar{p}$.
Thus, we have
$\mathbb{E}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert^{2}$ $\leq\mathbb{E}\big{[}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert^{2}|E\big{]}\cdot\mathbb{P}(E)+\mathbb{E}\big{[}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert^{2}|{\bar{E}}\big{]}\cdot\mathbb{P}(\bar{E})$
$\leq\frac{C_{lp}}{\nu\lambda(T-kf)kf}\cdot\frac{1}{\alpha}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}\cdot\alpha+\frac{\bar{p}}{T^{4}}.$ (65)
Thus, as $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for any $a,b>0$, by (D.2), we have
$\mathbb{E}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert$ $\leq\frac{C_{lp}}{\sqrt{\nu\lambda(T-kf)kf}}\cdot\frac{1}{\sqrt{\alpha}}+\frac{m(\bar{a}+\bar{d})}{\sqrt{\nu\lambda}}\cdot\sqrt{\alpha}+\frac{\bar{p}}{T^{2}}.$
This completes the proof.

### v3

@@
-appReferences in the Appendix
+Wait-Less Offline Tuning and Re-solving for Online Decision Making
 Abstract
@@
 However, LP-based methods are computationally expensive and often inefficient for large-scale applications.
-In contrast, recent first-order OLP algorithms are more computationally efficient but typically suffer from worse regret guarantees.
-To address these shortcomings, we propose a new algorithm that combines the strengths of LP-based and first-order OLP methods.
-The algorithm re-solves the LP subproblems periodically at a predefined frequency $f$ and uses the latest dual prices to guide online decision-making.
-In addition, a first-order method runs in parallel during each interval between LP re-solves, smoothing resource consumption.
-Our algorithm achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret, delivering a “wait-less” online decision-making process that balances the computational efficiency of first-order methods and the superior regret guarantee of LP-based methods.
+By contrast, recent first-order OLP algorithms are more computationally efficient but typically suffer from weaker regret guarantees.
+To address these shortcomings, we propose a new algorithm that combines the strengths of LP-based and first-order OLP algorithms.
+Our algorithm re-solves the LP subproblems periodically at a predefined frequency $f$ and uses the latest dual prices to guide online decision-making.
+In parallel, a first-order method runs during each interval between LP re-solves and smooths resource consumption.
+Our algorithm achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret and delivers a “wait-less” online decision-making process that balances computational efficiency and regret guarantees.
+Extensive experiments demonstrate at least $10$-fold improvements in regret over first-order methods and $100$-fold improvements in runtime over LP-based methods.
+Online Linear Programming, Sequential Decision-Making, Machine Learning
 1 Introduction
-Sequential decision-making has garnered increasing attention for its utility in guiding optimal strategies in dynamic environments.
-It addresses the problems of identifying the optimal decisions and policies in environments where knowledge of the system continuously accumulates and evolves over time.
-In this paper, we study Online Linear Programming (OLP) [1], a powerful paradigm that encapsulates the core principles of sequential decision-making.
-OLP methods have been extensively applied to diverse areas, including resource allocation [2], online advertising [19], and inventory management [21].
-In the OLP problem, customers arrive sequentially, each presenting a resource request and an associated bidding price.
-The objective is to determine which resource requests to fulfill to maximize total revenue while ensuring that no resource is ever over-allocated.
-The key challenge is that decisions must be made immediately and irrevocably, relying solely on historical data without knowledge of future arrivals.
-The goal is to minimize the cumulative regret with respect to the optimal hindsight linear programming (LP) solution.
-Most OLP algorithms use online learning methods to gradually refine their estimates of the optimal dual prices, using these values to guide near-optimal decision-making.
-Commonly, these algorithms rely either on LP-based or first-order methods.
-LP-based methods repeatedly solve linear programs at each timestep with all available information, thereby updating the dual prices.
-However, the substantial computational demands of LP-based methods often restrict their application in time-sensitive settings.
-In contrast, first-order methods leverage gradient information for quick, incremental updates to the dual prices.
-However, they generally cannot achieve the same $\mathcal{O}(\log{T})$ regret guarantee as the LP-based methods [16].
-These trade-offs motivate an open question at the intersection of online learning and decision making:
-Can we simultaneously achieve low regret and efficiency?
-We answer this question affirmatively in the well-studied stochastic input model–where the resource request and bidding price for each customer are sampled identically and independently from an unknown distribution—by proposing a decision-making framework that combines LP-based and first-order OLP algorithms.
-Our algorithm iteratively optimizes the dual prices, which progressively converge to the optimal dual solution.
-1.1 Main Contributions
-We propose a parallel multi-stage framework for OLP that integrates online learning and decision-making.
-Building on this framework, we develop a mathematical analysis that reconciles existing OLP approaches, ultimately allowing us to present an algorithm that achieves logarithmic worst-case regret.
-We formalize the trade-off between decision quality and computational efficiency, enabling end-users to quantify how allocating additional computational resources will reduce regret.
-In more detail, our contributions can be summarized as follows:
-  • Parallel Multi-stage Framework: We separate online learning and decision-making into two separate yet interlinked processes, connected through a feedback loop.
-We re-solve the LP subproblems periodically at a fixed frequency and pass the result to the decision-making path.
-These results guide decision-making until the next time that the LP is re-solved.
-Additionally, to enhance the algorithm’s efficiency, we apply a first-order method during the initial and final stages, restarting with the most recent learning outcomes.
-By integrating the LP-based and first-order techniques, this approach leverages their respective strengths.
-  • Regret Analysis: We develop a mathematical analysis for our algorithm where we derive a new performance metric to account for the inter-dependency between the LP-based and first-order methods.
-We unify the analysis of these two methods and provide a “spectrum theorem” that bounds regret for any feasible choice of the re-solving frequency in $[1,T]$:
+Sequential decision-making has garnered significant attention for its utility in guiding optimal strategies in dynamic environments.
+The goal is to identify effective decisions and policies in environments where knowledge of the system continuously accumulates and evolves.
+Online Linear Programming (OLP) (Agrawal et al., 2014) offers a powerful framework that encapsulates the core principles of sequential decision-making and has been extensively applied to different domains, including resource allocation (Balseiro et al., 2022a), online advertising (Mehta et al., 2007), and inventory management (Talluri et al., 2004).
+We study an OLP problem where customers arrive sequentially, each requesting a combination of resources and offering a bidding price.
+The objective is to determine which resource requests to fulfill to maximize revenue while respecting resource constraints.
+The challenge is that decisions must be made immediately and irrevocably, relying solely on historical data without knowledge of future arrivals.
+The goal is to minimize regret with respect to the optimal hindsight linear programming (LP) solution.
+To guide real-time decision-making, state-of-the-art OLP algorithms estimate optimal dual prices and use them to make decisions.
+These algorithms fall into two main categories: LP-based and first-order methods.
+Specifically, LP-based methods update dual prices by repeatedly solving linear programs at each time step with all available information so far.
+However, the substantial computational demands limit their application in time-sensitive settings.
+First-order methods offer quick, incremental updates to dual prices using gradient information, but generally fall short of achieving the strong regret guarantee of LP-based methods.
+These trade-offs motivate an open question at the intersection of online learning and decision-making:
+Can we simultaneously achieve
+low regret and computational efficiency?
+Our contributions.
+We answer this question in the affirmative. We summarize our contributions as follows:
+  • Parallel Multi-Stage Framework: We separate online learning and decision-making into distinct processes that interact via a feedback loop.
+We re-solve the LP subproblems periodically at a fixed frequency and feed the updated dual price to guide decision-making until the next time that the LP is re-solved.
+To further enhance efficiency, we apply a first-order method during the initial and final stages, restarting with the most recent learning outcomes.
+By integrating the LP-based and first-order techniques, our approach leverages their respective strengths and achieves a balance between regret performance and computational costs.
+  • Regret Analysis: We unify the analysis of LP-based and first-order methods by deriving a new performance metric to account for their inter-dependency.
+Our analysis yields a “spectrum theorem” that bounds regret for any feasible choice of the re-solving frequency:
     Theorem 1.1 (Informal version of Theorem 3.2).
-    If we re-solve the updated linear programs every $f$ timesteps over the total time horizon of length $T$ and utilize gradient information in the initial and final $f$ steps, our algorithm is guaranteed to achieve a worst-case regret of
-    $\mathcal{O}\big{(}\log\big{(}\tfrac{T}{f}\big{)}+\sqrt{f}\big{)}.$ (1)
-    When $f=1$, the algorithm is a pure LP-based method and reduces to [16], and when $f=T$, the algorithm is a pure first-order method with regret $\mathcal{O}(\sqrt{T})$.
-In our analysis, one can choose the “optimal” $f$ to balance the LP-solving times and regret.
-Compared to [23], our method enables a “wait-less” decision-making system for customers during the whole process.
+    If we re-solve the LPs every $f$ time steps within horizon $T$ and apply a first-order method in the initial and final $f$ steps, we achieve a worst-case regret of $\mathcal{O}\big(\log(T/f)+\sqrt{f}\big).$
+    In particular, $f=1$ yields a pure LP-based method with $\mathcal{O}(\log T)$ regret, while $f=T$ reduces to a pure first-order method with $\mathcal{O}(\sqrt{T})$ regret.
+By choosing an “optimal” $f$, one can achieve the best possible regret based on available computational resources and enable a “wait-less” decision-making system across all time steps.
+  • Experiments: Through experiments across diverse distributions, we demonstrate that our algorithm achieves over a $10$-fold improvement in regret compared to first-order methods and over a $100$-fold improvement in runtime with comparable regret to LP-based methods.
+Key Challenges.
+OLP methods face a fundamental challenge in balancing efficient real-time decision-making and accurate dual-price learning.
+Specifically, LP-based methods ensure high-quality decisions with $\mathcal{O}(\log T)$ regret, but their computational costs increase quadratically with problem size.
+A natural alternative—batching customers and solving the LP once every $f$ arrivals—reduces this cost, but leaves customers waiting until the batch concludes.
+In contrast, first-order methods update dual prices via gradient information, allowing for faster computation.
+However, these methods require small step sizes to maintain decision quality, which slows their adaptation to new data.
+Even using different step sizes for learning and decision-making only achieves $\mathcal{O}(T^{1/3})$ regret when the distribution of customers has continuous support.
+Our parallel framework addresses these issues: an LP-based method periodically refines dual prices, while a first-order method immediately processes arriving customers using the most recent dual updates, thereby eliminating delays.
+A second challenge emerges when analyzing the regret of this hybrid approach.
+LP-based methods rely on a stopping-time analysis (halting when resources are depleted), whereas first-order methods track constraint violations as part of the regret.
+Since these two formulations are fundamentally different, existing regret bounds cannot be naively combined.
+We overcome this challenge by introducing a unified performance metric that decomposes regret into three components: dual convergence, constraint violation, and leftover resources.
+This approach yields the first integrated analysis of LP-based and first-order OLP, culminating in our $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret bound.
 1.1 Related Literature
-Online resource allocation and OLP have been widely studied under two predominant models: the stochastic input model [9, 6] and the stochastic permutation model [1, 11].
-In our paper, we study the stochastic input model, where each customer’s resource request and bidding price—represented as columns of the constraint matrix—are drawn i.i.d.
-from an unknown distribution $\mathcal{P}\in\Xi$.
-Table 1: Performances in current OLP literature
+Online resource allocation and OLP problems have been widely studied under two predominant models: the stochastic input (Goel & Mehta, 2008; Devanur et al., 2019) and stochastic permutation models (Agrawal et al., 2014; Gupta & Molinaro, 2016).
+We study the former, where each customer’s resource request and bidding price are drawn i.i.d.
+from an unknown distribution.
+We use expected regret and constraint violation as the performance metric.
+We summarize some recent works in Table˜1.
+LP-based Methods.
+These methods solve the OLP dual repeatedly to update dual prices and make decisions.
+Early approaches enforced a fixed average resource constraint (Agrawal et al., 2014), whereas recent work dynamically tracks remaining resources and enables LP-based methods to achieve $\mathcal{O}(\log T)$ regret under continuous support and non-degeneracy assumptions (Li & Ye, 2022).
+Variants include multi-secretary problems (Bray, 2022), regularized resource constraints (Ma et al., 2024), and finite-support distributions yielding constant regret (Chen et al., 2024).
+First-order Methods.
+These methods generate decisions using gradient updates without solving LPs, enabling efficient computation.
+They achieve $\mathcal{O}(\sqrt{T})$ regret with mirror descent (Li et al., 2020; Balseiro et al., 2022a) and $\mathcal{O}(T^{3/8})$ under finite support and non-degeneracy assumptions (Sun et al., 2020).
+Variants include proximal updates (Gao et al., 2023), momentum-based mirror descent (Balseiro et al., 2022b), resource adjustments (Ma et al., 2024), and restart strategies yielding $\mathcal{O}(T^{1/3})$ regret (Gao et al., 2024).
+Delay in Decision-making.
+Table 1: Performances of Dual Algorithms in Recent OLP Literature
 Paper Setting Algorithm Regret Decision-Making
-[16] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T\log\log T)$ Delay
-[4] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
-[17] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
-[12] Bounded, finite support LP-based $\mathcal{O}(\log^{2}T)$ Delay
-[5] Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
-[14] Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
-[23] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T/f)$ Delay
+(Li & Ye, 2022) Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T\log\log T)$ Delay
+(Bray, 2022) Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
+(Chen et al., 2024) Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
+(Li et al., 2024) Bounded, finite support LP-based $\mathcal{O}(1)$ Delay
+(Xu et al., 2024) Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log(T/f))$ Delay
 This paper Bounded, continuous support, non-degeneracy LP-based & First-order $\color[rgb]{1,0,0}\mathcal{O}(\log(T/f)+\sqrt{f})$} No Delay
-[15] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
-[2] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
-[7] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
-[3] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
-[20] Bounded, finite support, non-degeneracy First-order $\mathcal{O}(T^{3/8})$ No Delay
-[8] Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(T^{1/3})$ No Delay
-[17] Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(\log^{2}T)$ No Delay
-LP-based Methods
-This class of algorithms focuses on the OLP dual problem, repeatedly solving it to update the decision variables.
-While these methods achieve low regret, they are computationally expensive.
-The dual LPs are formulated based on the information revealed by each arriving customer.
-Early versions of these algorithms maintained a fixed average resource consumption rate throughout the time horizon [1], whereas more recent variants adaptively adjust the consumption rate based on the remaining resources, updating the LP accordingly [16].
-Under standard non-degeneracy and continuous support assumptions, these LP-based methods can achieve a regret bound of $\mathcal{O}(\log T)$.
-Several related works explore variants of this approach: [4] studies a multi-secretary problem and also attains $\mathcal{O}(\log T)$ regret; [17] incorporates a regularization term on the resource and also achieves $\mathcal{O}(\log T)$ regret; [12] assumes that the resource distribution has a finite support and obtains $\mathcal{O}(\log^{2}T)$ regret; and [5] makes a stronger assumption that both the resource and price distributions have finite support, leading to a constant regret.
-First-order Methods
-These algorithms leverage gradient information to update the decision variables directly without explicitly solving LPs.
-Although less accurate than LP-based methods, they enable more efficient computation and are therefore well-suited for large-scale or time-sensitive applications.
-First-order methods can attain a regret of $\mathcal{O}(\sqrt{T})$ with mirror descent and subgradient information under distributions with continuous support [15, 2], and $\mathcal{O}(T^{3/8})$ under non-degeneracy assumptions for finite support distributions [20].
-Furthermore, [7] attains a $\mathcal{O}(\sqrt{T})$ regret under proximal point updates, while [3] obtains the same regret via a momentum variant of mirror descent.
-[8] utilizes a restart strategy to decouple the learning and decision-making processes and improves the regret to $\mathcal{O}(T^{1/3})$.
-[17] integrates the resource consumption rate adjustment from LP-based methods into first-order methods and obtains $\mathcal{O}(\log^{2}T)$ regret.
-Delay in Decision-making
-This line of research examines the impact of decision delays in online resource allocation.
-[10] analyzes an online setting with a mix of impatient customers, whose requests must be decided upon immediately and irrevocably, and partially patient customers, for whom decisions can be postponed for several time periods.
-[22] demonstrates that deferring decisions—either by introducing explicit delays or batching customers—can substantially reduce regret.
-In concurrent research, [23] partially solves this delay issue by solving LPs in batches.
-While their approach is similar in spirit to ours, they make a stronger assumption that each resource request has a positive lower bound in expectation and their algorithm still requires customers to wait for decisions in the initial and final batches.
-Compared with the aforementioned works, our framework imposes minimal assumptions on the problem parameters, attains a regret bound of $\mathcal{O}(\log(T/f)+\sqrt{f})$, and provides a “wait-less” decision-making process throughout the entire time horizon.
+(Li et al., 2020) Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
+(Balseiro et al., 2022a) Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
+(Gao et al., 2023) Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
+(Sun et al., 2020) Bounded, finite support, non-degeneracy First-order $\mathcal{O}(T^{3/8})$ No Delay
+(Gao et al., 2024) Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(T^{1/3})$ No Delay
+Delays arise from the time-consuming process of solving the large-scale, up-to-date LP subproblems for each customer.
+Golrezaei & Yao (2021) study a mix of impatient and partially patient customers, while Xie et al.
+(2023) show that batching requests can reduce regret.
+Concurrent work (Xu et al., 2024) reduces delay by solving LPs in batches but assumes a lower bound on resource requests and still requires waiting in initial and final batches.
+To achieve delay-free decisions, we shift the re-solving process offline and only fine-tune the solution online.
+Compared with previous works, our framework imposes standard assumptions, achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret, and ensures “wait-less” decision-making throughout the entire horizon.
+Paper organization.
 The rest of the paper is organized as follows.
-Section 2 introduces the problem formulation, preliminary dual convergence results, and the main assumptions.
-Section 3 proposes our novel framework to integrate the LP-based and first-order methods and provides our “spectrum theorem,” which bounds the regret of our algorithm by $\mathcal{O}(\log(T/f)+\sqrt{f})$ for any choice of $f$.
-Section 4 presents experiments that illustrate the strong performance of our algorithm and validate our theoretical conclusions.
+Section˜2 introduces the problem formulation and assumptions.
+Section 3 proposes our algorithms and main theoretical guarantee: a $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret bound.
+Section 4 presents experiments to validate our theory.
 2 Problem Setup
-[…]
-2.1 Online LP Formulation
-Consider an online resource allocation problem over the time horizon $T$.
-Initially, we have an inventory vector $\bm{b}\in\mathbb{R}^{m}$, representing $m$ types of resources.
-At each time step $t$, a customer arrives with a request characterized by $(r_{t},\bm{a}_{t})\sim\mathcal{P}$, where $r_{t}\in\mathbb{R}$ is the offered payment (bid), $\bm{a}_{t}\in\mathbb{R}^{m}$ is the customer’s demand for each resource, and $\mathcal{P}$ is a fixed, unknown distribution.
-We must decide whether to accept or reject this request, represented by the decision variable $x_{t}\in\{0,1\}$.
+We use $\|\cdot\|$ to denote the Euclidean norm and $\langle\cdot\rangle$ to denote Euclidean inner product.
+Bold letters $\bm{A}$ and $\bm{a}$ denote matrices and vectors, respectively.
+Subscript $(\cdot)_{it}$ denotes the index for resource type $i$ at time $t$.
+The notation $(\cdot)^{+}=\max\{\cdot,0\}$ denotes the element-wise positive part function, and $\mathbb{I}(\cdot)$ denotes the 0-1 indicator function.
+2.1 OLP Formulation
+We study an online resource allocation problem over the time horizon $T$ under a stochastic input model.
+Initially, we have an inventory vector $\bm{b}\in\mathbb{R}^{m}$, representing $m$ resource types.
+At each time step $t$, a customer arrives with a request sampled i.i.d.
+as $(r_{t},\bm{a}_{t})\sim\mathcal{P}$, where $\bm{r}=(r_{1},\ldots,r_{T})^{\top}\in\mathbb{R}^{T}$ is the offered payment (bid), $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{T})\in\mathbb{R}^{m\times T}$ is the matrix of customers’ resource demands, and $\mathcal{P}$ is a fixed, unknown distribution.
+We must decide whether to accept or reject each request, represented by the decision variables $\bm{x}=(x_{1},\ldots,x_{T})\in\{0,1\}^{T}$.
 The goal is to maximize the cumulative reward.
-This problem can be formulated as the following OLP, which we refer to as the primal linear program (PLP):
-$\max_{\bm{x}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{t=1}^{T}r_{t}x_{t}$ (PLP)
-subject to $\leavevmode\nobreak\ \quad\sum_{t=1}^{T}a_{it}x_{t}\leq b_{i},i=1,...,m$
-[…]
+This problem can be formulated as the following OLP, referred to as the primal linear program (PLP):
+$\max_{\bm{0}\leq\bm{x}\leq\bm{1}}\langle\bm{r},\bm{x}\rangle\quad\text{s.t.}\quad\bm{Ax}\leq\bm{b}.$ (PLP)
 The dual problem of (PLP) is given by
-$\min_{\bm{p},\bm{y}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=1}^{m}b_{i}p_{i}+\sum_{t=1}^{T}y_{t}$ (DLP)
-subject to $\leavevmode\nobreak\ \quad\sum_{i=1}^{m}a_{it}p_{i}+y_{t}\geq r_{t},t=1,...,T$
-$\quad\ \ p_{i},y_{t}\geq 0,\forall i,t,$
-where $\bm{p}$ is the vector of dual prices. Let $\bm{d}=\bm{b}/T\in\mathbb{R}^{m}$ denote the initial average resource capacity. As demonstrated by [15], (DLP) can be written as
-$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f_{T}(\bm{p})\coloneqq\frac{1}{T}\sum_{t=1}^{T}\left[\sum_{i=1}^{m}d_{i}p_{i}+\left(r_{t}-\sum_{i=1}^{m}a_{it}p_{i}\right)^{+}\right]$ (2)
-subject to $\leavevmode\nobreak\ \quad p_{i}\geq 0,i=1,...,m$
-This can be regarded as a $T$-sample approximation of the following stochastic program:
-$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f(\bm{p})\coloneqq\mathbb{E}[f_{T}(\bm{p})]=\bm{d}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (3)
-subject to $\leavevmode\nobreak\ \quad\bm{p}\geq\bm{0}$
-where the expectation is taken with respect to the coefficient pair $(r,\bm{a})$ drawn from the unknown distribution $\mathcal{P}$.
-We define the optimal solutions to the $T$-sample approximation problem (2) and the stochastic program (3) respectively as
-$\bm{p}_{T}^{*}=\arg\min_{\bm{p}\geq\textbf{0}}\leavevmode\nobreak\ f_{T}(\bm{p})\quad\text{and}\quad\bm{p}^{*}=\arg\min_{\bm{p}\geq\textbf{0}}\leavevmode\nobreak\ f(\bm{p}).$
-The decision variables $x_{t}^{*}$ can be established from the complementary slackness condition as
-$x_{t}^{*}=\left\{\begin{array}[]{ll}\ 0,&r_{t}<\bm{a}_{t}^{\top}\bm{p}_{T}^{*},\vspace{3 pt}\\ \ 1,&r_{t}>\bm{a}_{t}^{\top}\bm{p}_{T}^{*}.\end{array}\right.$ (6)
+$\min_{(\bm{p},\bm{y})\geq\bm{0}}\langle\bm{b},\bm{p}\rangle+\langle\bm{1},\bm{y}\rangle\quad\text{s.t.}\quad\bm{A}^{\top}\bm{p}+\bm{y}\geq\bm{r}$ (DLP)
+where $\bm{p}$ is the vector of dual prices. Let $\bm{d}=\bm{b}/T\in\mathbb{R}^{m}$ denote the initial average resource. As demonstrated by (Li et al., 2020), (DLP) can be written as
+$\min_{\bm{p}\geq\bm{0}}f_{T}(\bm{p})\coloneqq\tfrac{1}{T}\textstyle\sum_{t=1}^{T}\left[\langle\bm{d},\bm{p}\rangle+(r_{t}-\langle\bm{a}_{t},\bm{p}\rangle)^{+}\right]$ (1)
+This formulation can be written as a $T$-sample approximation of the following stochastic program:
+$\min_{\bm{p}\geq\bm{0}}f(\bm{p})\coloneqq\mathbb{E}[f_{T}(\bm{p})]=\bm{d}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (2)
+where the expectation is taken with respect to $(r,\bm{a})\sim\mathcal{P}$.
+We define optimal solutions to the $T$-sample approximation problem (1) and stochastic program (2) respectively as
+$\bm{p}_{T}^{*}=\arg\min_{\bm{p}\geq\bm{0}}\penalty 10000\ f_{T}(\bm{p})\quad\text{and}\quad\bm{p}^{*}=\arg\min_{\bm{p}\geq\bm{0}}\penalty 10000\ f(\bm{p}).$
+The decision variable $x_{t}^{*}$ can be established from the complementary slackness condition as
+$x_{t}^{*}=\left\{\begin{array}[]{ll}\ 0,&r_{t}<\bm{a}_{t}^{\top}\bm{p}_{T}^{*},\vskip 2.0pt\\ \ 1,&r_{t}>\bm{a}_{t}^{\top}\bm{p}_{T}^{*}\end{array}\right$. (5)
 and $x_{t}^{*}\in[0,1]$ if $r_{t}=\bm{a}_{t}^{\top}\bm{p}_{T}^{*}$.
-This connection between the primal-dual optimal solutions motivates us to propose dual-based online LP algorithms and then provide a convergence analysis of $\bm{p}_{T}^{*}$ to $\bm{p}^{*}$.
-2.2 Online Dual Algorithms
-We introduce two basic dual-based algorithms for the OLP problem. We calculate and update the dual prices $\{\bm{p}_{t}\}_{t=1}^{T}$ and compute decisions $\{x_{t}\}_{t=1}^{T}$ based on rule (6).
-2.2.1 LP-based Algorithm
-In this method, we calculate the dual price by re-solving the online LP problem at each time step.
-The model is dynamically updated to incorporate real-time capacity constraints derived from the Action-history-dependent learning (AHDL) algorithm [16].
-Define $d_{it}=\frac{b_{it}}{T-t}$ to be the average remaining resource capacity for resource $i$ at time $t$.
+This connection between primal and dual solutions inspires OLP algorithms that make decisions based on dual prices:
+$x_{t}=\mathbb{I}(r_{t}\geq\bm{a}_{t}^{\top}\bm{p}_{t}).$ (6)
+2.2 Algorithms for OLP
+We summarize two main dual-based OLP algorithms.
+LP-based Method.
+The LP-based method (Li & Ye, 2022) calculates dual prices by re-solving the online LP problem at each time step.
+Specifically, define $d_{it}=b_{it}/(T-t)$ as the average remaining resource for type $i$ at time $t$.
 The resulting optimization problem can be viewed as a $t$-sample approximation to the stochastic program specified by $\bm{d}_{t}=(d_{it},...,d_{mt})^{\top}$ as
-$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f_{\bm{d}_{t}}(\bm{p})\coloneqq\bm{d}_{t}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (7)
-subject to $\leavevmode\nobreak\ \quad\bm{p}\geq\bm{0}.$
-We denote the optimal solution to the stochastic program (7) as $\bm{p}_{t}^{*}$.
-Thus, instead of focusing on a static problem with fixed initial $\bm{d}$, we formulate a dynamic stochastic program that evolves according to the changing constraints.
-We outline this method in Algorithm 1.
-Algorithm 1 LP-based Online AHDL algorithm
-Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0}.$
-for $t$ = $1$ to $T$  do
-       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6) if constraints are not violated.
-Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Solve the updated dual problem and obtain dual price $\bm{p}_{t+1}$:
-$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
-end for
-This method incorporates the past decisions into the optimization of $\bm{p}_{t}$.
-If resources were over-utilized in earlier periods, the remaining supply decreases, prompting the algorithm to raise the dual price and become more selective with future orders.
-Conversely, if ample resources remain, the future dual price will be lowered, allowing more consumer requests to be accepted.
-This adaptive mechanism accounts for past actions by adjusting the available resource capacity.
-2.2.2 Subgradient-based Algorithm
-In this method, we calculate the dual price by iteratively moving in the direction of the steepest cost gradient.
-We maintain a static resource constraint, as the historical information is integrated through the gradient updates.
-Due to its lower computational overhead, the subgradient-based is well-suited for high-dimensional and real-time applications.
-Algorithm 2 summarizes this method, which is also known as the “first-order” method in [15].
-Algorithm 2 First-Order Online algorithm
-Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0}.$
-for $t$ = $1$ to $T$  do
-       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6). Update learning rate $\alpha_{t}=\frac{1}{t+1}$. Compute subgradient and obtain dual price $\bm{p}_{t+1}$:
-$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ (\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}+\frac{1}{2\alpha_{t}}\lVert\bm{p}-\bm{p}_{t}\rVert^{2}$
-end for
-We can derive a closed-form solution for the updated dual price from the above rule as
-$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$
-where the subgradient term evaluated at $\bm{p}_{t}$ is calculated as
+$\min_{\bm{p}\geq\bm{0}}$ $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ f_{\bm{d}_{t}}(\bm{p})\coloneqq\bm{d}_{t}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (7)
+By updating $\bm{d}_{t}$, this method incorporates past decisions when computing $\bm{p}_{t}$.
+If resources are over-utilized in earlier periods, the supply decreases, prompting us to raise the dual price and be more selective with future orders.
+We outline this method in Algorithm˜3 of Section˜A.2.
+First-order Method.
+The first-order method (Li et al., 2020) calculates dual prices via online subgradient updates.
+We maintain a static average resource $\bm{d}=\bm{b}/T$, compute $x_{t}$ as per (6), and update the dual price as
+$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))^{+}$
+where the subgradient term evaluated at $\bm{p}_{t}$ is
 $\bm{d}-\bm{a}_{t}x_{t}\in\partial_{\bm{p}=\bm{p}_{t}}\big(\bm{d}^{\top}\bm{p}+(r_{t}-\bm{a}_{t}^{\top}\bm{p})^{+}\big).$
-Thus, this process can be interpreted as a projected stochastic subgradient descent method for solving the optimization problem (2).
-The method reduces computational costs by requiring only a single pass through the data and eliminates the need for solving LPs.
+This process can be interpreted as a projected stochastic subgradient method for solving (1).
+It reduces computational cost by requiring only a single pass through the data and eliminates solving LPs explicitly.
+A restart strategy improves the first-order method to have $\mathcal{O}(T^{1/3})$ regret (Gao et al., 2024).
+Algorithm˜4 and Algorithm˜5 in Section˜A.3 summarizes this method.
 2.3 Performance Metrics
-We evaluate the algorithms using a bi-objective performance measure that involves the optimality gap and constraint violation.
-The optimality gap measures the difference between the objective value of the algorithm’s output and that of the true optimal solution, while the constraint violation measures the degree to which the algorithm’s output fails to meet the given constraints.
-We denote the order-request matrix by $\bm{A}=[\bm{a}_{1},...,\bm{a}_{T}]^{\top}$, the offline optimal solution to problem (PLP) by $\bm{x}^{*}=(x_{1}^{*},...,x_{T}^{*})^{\top}$, and the online optimal solution by $\bm{x}=(x_{1},...,x_{T})^{\top}$.
-Then, we define the regret $r(\bm{x})$ and resource violation $v(\bm{x})$ as:
-$r(\bm{x})$ $\coloneqq\max_{\mathbf{A}\mathbf{x}\leq\bm{b},\mathbf{0}\leq\mathbf{x}\leq\mathbf{1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{t=1}^{T}r_{t}x_{t}^{*}-r_{t}x_{t},$ (8)
+We aim to design algorithms that optimize a bi-objective performance measure involving regret and constraint violation.
+The regret measures the difference between the objective value of the algorithm’s output and that of the true optimal solution, while the constraint violation measures the degree to which the algorithm’s output fails to meet the given constraints.
+We denote the offline optimal solution to (PLP) by $\bm{x}^{*}=(x_{1}^{*},\ldots,x_{T}^{*})$, and the online algorithm output by $\bm{x}=(x_{1},\ldots,x_{T})$.
+Then, we define the regret $r(\bm{x})$ and resource violation $v(\bm{x})$ as
+$r(\bm{x})$ $\coloneqq\langle\bm{r},\bm{x}^{*}\rangle-\langle\bm{r},\bm{x}\rangle,$ (8)
 $v(\bm{x})$ $\coloneqq\lVert(\bm{Ax}-\bm{b})^{+}\rVert.$ (9)
-Therefore, we define the following bi-objectives for evaluating the total regret and worst-case total regret of the algorithms we propose:
+Therefore, we define the following bi-objectives for evaluating an algorithm’s worst-case performance:
 $\Delta_{T}=\sup_{\mathcal{P}\in\Xi}\mathbb{E}_{\mathcal{P}}[r(\bm{x})+v(\bm{x})]$ (10)
-where $\mathcal{P}$ is the unknown distribution of $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ and $\Xi$ denotes the family of distribution satisfying some regularity assumptions specified later.
-This metric is commonly used for first-order OLP algorithms in [7, 15] and is also aligned with the literature on the online convex optimization with constraints in [18, 24].
-For LP-based methods, [16] use stopping time to bound the regret.
-Equation (10) integrates the analysis of subgradient methods and LP-based methods into a single performance metric, enabling more consistent analysis across online learning scenarios.
-This integrated method promotes more balanced resource consumption over the decision horizon.
-This flexibility makes our framework well-suited for scenarios without prior knowledge of the constraint support, thereby expanding its applicability to a broad range of online learning and resource allocation problems.
+where $\Xi$ denotes a family of distributions satisfying regularity assumptions specified later.
+This metric is commonly used for first-order OLP algorithms (Gao et al., 2023; Li et al., 2020) and is also aligned with the literature on online convex optimization with constraints (Mahdavi et al., 2012; Yu et al., 2017).
+Integrating $r(\bm{x})$ and $v(\bm{x})$ into a single performance measure promotes balanced resource consumption over the decision horizon.
+In Section˜3, we derive a decomposition of (10) that unifies the regret analysis for all OLP methods.
 2.4 Assumptions and Auxiliary Results
-We propose the following assumptions regarding the stochastic inputs used throughout the paper.
-These minimal assumptions are standard in the online learning literature [16, 13, 23].
+We adopt the following assumptions regarding the stochastic inputs.
+These assumptions are standard in the online learning literature (Li & Ye, 2022; Jiang et al., 2022; Xu et al., 2024).
 In particular, we require the input data to be bounded, follow a linear growth, and be non-degenerate.
-Assumption 1 (Boundedness).
+Assumption 2.1 (Boundedness).
 We assume
@@
   2. (b)
-    There exist constants $\bar{r},\bar{a}>0$ such that $|r_{t}|\leq\bar{r}$ and $\lVert\bm{a}_{t}\rVert_{\infty}\leq\bar{a}$ almost surely for $t=1,...,T$.
+    There exist constants $\bar{r},\bar{a}>0$ such that $|r_{t}|\leq\bar{r}$ and $\lVert\bm{a}_{t}\rVert_{\infty}\leq\bar{a}$ almost surely for t=1,,Tt=1,...,T.
   3. (c)
-    The average resource capacity $\bm{d}=\bm{b}/T$ satisfies $d_{i}\in[\underline{d},\bar{d}]$ for some $\bar{d}>\underline{d}>0$ for any $i=1,...,m$.
+    The average resource capacity $\bm{d}=\bm{b}/T$ satisfies $d_{i}\in[\underline{d},\bar{d}]$ for some $\bar{d}>\underline{d}>0$ for any i=1,,mi=1,...,m.
 In this assumption, (a) states that $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ are independent of each other, but we allow dependencies between their components.
-Part (b) introduces the bounds $\bar{r},\bar{a}$ solely for analytical purposes; they are not used within the algorithm itself.
-This is a minimal requirement on $(r_{t},\bm{a}_{t})$ compared to previous work [23].
-Part (c) requires the average resource capacity to grow linearly with $T$, ensuring that a constant fraction of the $x_{t}$ values can be set to 1 in the optimal solutions.
+Part (b) introduces the bounds $\bar{r},\bar{a}$ solely for analytical purposes.
+This is a minimal requirement on $(r_{t},\bm{a}_{t})$ compared to previous work (Agrawal et al., 2014; Li & Ye, 2022; Xu et al., 2024).
+Part (c) requires the average resource to grow linearly with $T$, ensuring that a constant fraction of $x_{t}$ values can be set to 1.
 Consequently, the number of fulfillable orders is proportional to $T$, facilitating a stable service level over time.
-Assumption 2 (Uniform Non-degeneracy).
+Assumption 2.2 (Uniform Non-degeneracy).
 We assume
   1. (a)
-    The second moment matrix $\mathbb{E}[\bm{a}\bm{a}^{\top}]$ is positive definite with minimum eigenvalue $\lambda$.
+    The second-moment matrix $\mathbb{E}[\bm{a}\bm{a}^{\top}]$ is positive definite with minimum eigenvalue $\lambda$.
   2. (b)
-    There exists constants $\mu,\nu$ such that for any $(r,\bm{a})\sim\mathcal{P}$,
-    $\nu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|\leq|\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}\mid\bm{a})-\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}^{*}\mid\bm{a})|\leq\mu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|$
-    holds for all $\bm{p}\in\mathcal{V}_{\bm{p}}\coloneqq\{\bm{p}\in\mathbb{R}^{m}:\bm{p}\geq\bm{0},\lVert\bm{p}\rVert\leq\frac{\bar{r}}{\underline{d}}\}$ and $\bm{d}\in\mathcal{V}_{\bm{d}}\coloneqq[\underline{d},\bar{d}]^{m}$.
+    There are constants $\mu,\nu$ such that for any $(r,\bm{a})\sim\mathcal{P}$,
+    $\nu|\bm{a}^{\top}(\bm{p}-\bm{p}^{*})|$
+    $\leq\$$\left|\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}\mid\bm{a})-\mathbb{P}(r\geq\bm{a}^{\top}\bm{p}^{*}\mid\bm{a})\right|$
+[…]
+    holds for all
+    $\bm{p}\in\mathcal{V}_{\bm{p}}\coloneqq\left\{\bm{p}\in\mathbb{R}^{m}:\bm{p}\geq\bm{0},\lVert\bm{p}\rVert\leq\frac{\bar{r}}{\underline{d}}\right\}$
+    and $\bm{d}\in\mathcal{V}_{\bm{d}}\coloneqq[\underline{d},\bar{d}]^{m}$.
   3. (c)
-    The optimal solution $\bm{p}^{*}$ satisfies $p_{i}^{*}=0$ if and only if $d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]>0$ for all $\bm{d}\in\mathcal{V}_{\bm{d}}$ and $i=1,...,m$.
+    The optimal $\bm{p}^{*}$ satisfies $p_{i}^{*}=0$ if and only if $d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]>0$ for all $\bm{d}\in\mathcal{V}_{\bm{d}}$ and $i\in[m]$.
 In this assumption, part (b) ensures that the cumulative distribution of the reward given the resource consumption request $r|\bm{a}$ is continuous and exhibits a stable growth rate.
-Part (c) requires strict complementarity for the optimal solutions of the stochastic program in (3), which is a non-degeneracy condition for both the primal and dual LP.
+Part (c) requires strict complementarity for the optimal solutions of the stochastic program in (2), which is a non-degeneracy condition for both the primal and dual LPs.
 According to stochastic program (2), we define the binding and non-binding index sets as:
-$I_{B}$ =$=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})=0]\},$
-$I_{N}$ =$=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})>0]\}.$ (11)
-Here, $I_{B}$ denotes the set of binding resources—those that are limited and require careful allocation—while $I_{N}$ denotes the set of non-binding resources, which are unlimited and can fulfill any order’s request.
-By Assumption 2(c), these sets are complements, as $I_{B}\cap I_{N}=\emptyset$ and $I_{B}\cup I_{N}=\{1,...,m\}$.
-Let $\Xi$ denote the family of distributions satisfying Assumptions 1 and 2. We propose the following lemma.
-Lemma 2.1 (Dimension Stability).
-Under Assumptions 1 and 2, there exists a constant $\delta>0$ such that
+$I_{B}$ $=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]=0\},$
+$I_{N}$ $=\{i:d_{i}-\mathbb{E}[a_{i}\mathbb{I}(r>\bm{a}^{\top}\bm{p}^{*})]>0\}.$ (11)
+By ˜2.2(c), these sets are complements, as $I_{B}\cap I_{N}=\emptyset$ and $I_{B}\cup I_{N}=\{1,...,m\}$.
+3 Parallel Multi-Phase OLP Algorithm
+We now present our algorithm for parallel multi-phase online learning and decision-making.
+Our approach is motivated by the challenges of existing methods.
+Specifically, the LP-based method has a strong worst-case regret bound of $\mathcal{O}(\log T)$, but its high computational costs lead to decision-making delays.
+Meanwhile, the first-order method updates decisions efficiently but suffers a high regret of $\mathcal{O}(\sqrt{T})$.
+By combining the strengths of these two methods, our new framework balances decision-making quality and computational efficiency.
+3.1 Algorithm Design
+Figure 1: Algorithm 1 illustration of parallel paths and the interactions between online learning and decision-making.
+Decisions are generated based on 1) the LP-based method (blue) with frequency $f$, 2) the first-order method (red) for the initial and final phases (with a warm start), and 3) employing the latest dual price (yellow) during intermediate phases.
+We establish our framework with the following design:
+  1. 1.
+    We maintain two parallel paths of online learning and online decision-making.
+The online learning results are periodically sent to the decision-making path as a re-start point to guide subsequent updates.
+  2. 2.
+    In the online learning path, we employ a streamlined LP-based method, which only re-solves the updated OLP problem according to a predefined frequency, reducing the computational overhead.
+  3. 3.
+    In the online decision-making path, we apply the first-order method during the initial and final batches. The learning rate is optimally tuned for these two intervals.
+Figure 1 illustrates our framework, which generates two parallel sequences of dual prices: $\{\bm{p}_{t}^{D}\}_{t=1}^{T}$ from the first-order method and $\{\bm{p}_{t}^{L}\}_{t=1}^{T}$ from the LP-based method.
+The algorithm proceeds in batches of length $f$.
+In the first batch, it uses the first-order method to iteratively update the dual prices, thereby guiding decision-making.
+At the end of this batch, it applies the LP-based method to obtain a refined dual price and passes it to the decision-making path.
+The LP re-solving occurs only once per batch, making the approach computationally efficient.
+Formally, we re-solve the OLP at every time $t$ satisfying $t\leq kf$ and $t\bmod f=0,$ where $k=\lfloor T/f\rfloor$ is the number of batches.
+During intermediate batches, decisions are made based on the most recent dual prices computed from the LP at the end of the previous batch.
+The algorithm restarts the subgradient updates only in the final batch to guide the remaining decisions.
+Algorithm 1 summarizes this approach.
+Algorithm 1 Parallel Multi-Phase OLP Algorithm
+Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of batches $k=\lfloor T/f\rfloor$
+for $t=$ 11 to $T$ do
+   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ as rule (6) Update constraint for i=1,,mi=1,...,m:   remaining resource $b_{it}=b_{i,t-1}-a_{it}x_{t}$   average remaining resource $d_{it}=\frac{b_{it}}{T-t}$ if $t\leq f$ or $t\geq kf$ then
+      Update learning rate $\alpha_{t}=\begin{cases}\mathcal{O}(1/f^{1/2})&t\leq f\\ \mathcal{O}(1/f^{2/3})&t\geq kf\end{cases}$ Compute subgradient and update dual price $\bm{p}_{t+1}$:
+$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))^{+}$
+  end if
+ if $t\bmod f=0$ and constraints are not violated then
+      Solve OLP and update dual price $\bm{p}_{t+1}$:
+$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\tfrac{1}{t}\textstyle\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
+  end if
+ else
+      Update dual price $\bm{p}_{t+1}$ to be the most recent solution $\bm{p}_{t+1}=\bm{p}_{t}$.
+   end if
+end for
+Algorithm 1 balances online learning and efficient decision-making.
+It is responsive to dynamic environments, adapts to the latest information, and reduces computational cost by periodic re-solving.
+We thus establish a “wait-less” online decision-making framework where each customer’s order is processed immediately without delays from earlier requests or large-scale LP computations.
+3.2 Algorithm Analysis
+We decompose the performance metric of Algorithm 1 into three key components. All proofs (including essential properties of the dual price $\bm{p}_{t}$) are in Appendix B and C.
+Theorem 3.1 (Performance Metric).
+Under Assumptions 2.1 and 2.2, the performance $\Delta_{T}$ of Algorithm 1 satisfies
+$\Delta_{T}\$$\leq\ \mu\bar{a}^{2}\textstyle\sum_{t=1}^{T}\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]+\mathbb{E}\left[\lVert(\bm{Ax}-\bm{b})^{+}\rVert\right]$
+$\quad+\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\left[\lVert(\bm{b}-\bm{Ax})^{B+}\rVert\right].$ (12)
+where $(\cdot)^{B+}$ indicates the projection of binding terms onto the positive orthant.
+Based on the definition of $\Delta_{T}$ (10), we derive the performance metric as the sum of three key components: dual convergence, remaining resources, and binding constraint violation.
+This bound demonstrates that our algorithm encourages 1) smooth and balanced resource utilization, and 2) full resource consumption by the end of the time horizon.
+Building on Theorem 3.1, we provide a “spectrum theorem” for the algorithm’s performance.
+Theorem 3.2.
+Under Assumptions 2.1 and 2.2, the worst-case performance $\Delta_{T}$ of Algorithm 1 is bounded by
+$\Delta_{T}\in\mathcal{O}\big(\log\left(\tfrac{T}{f}\right)+\sqrt{f}\big)$ (13)
+where $T/f$ represents the number of re-solving batches and $f$ is the length of each batch.
+Remark 3.3 (Spectrum Theorem).
+We elucidate the trade-offs in total regret induced by using LP-based methods with $\mathcal{O}(\log T)$ regret and first-order methods with $\mathcal{O}(\sqrt{T})$ regret.
+Our algorithm introduces a re-solving frequency $f\in[1,T]$ and achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret, which recovers previous results with extreme cases of $f=1$ and $f=T$.
+Specifically, the first-order method gives us the regret of $\sqrt{f}$ for the first batch and $f^{1/3}$ for the last batch, and the LP-based method contributes the regret of $\log(T/f)$.
+Remark 3.4 (Warm Start).
+We can obtain a tighter regret bound of $\mathcal{O}(\log(T/f)+f^{1/3})$ given a warm start of the initial dual price satisfying $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ or if we use the LP-based method at each time step for the first batch.
+Remark 3.5 (Learning Rate Selection).
+We select the best learning rate to minimize the regret upper bound consisting of (8) and (9).
+For the first batch, the regret grows linearly with $\alpha_{t}$ while the constraint violation is inversely proportional to $\alpha_{t}$.
+To achieve the tightest bound, we balance this trade-off by selecting the learning rate that minimizes the overall expression, which yields the optimal choice $\alpha_{t}=1/\sqrt{f}$ as derived in Theorem B.7.
+A similar analysis for the final batch leads to the choice $\alpha_{t}=1/f^{2/3}$ as shown in Theorem B.9.
+Technical Intuitions.
+Theorem 3.1 decomposes the performance metrics to make the regret analysis compatible between the LP-based and first-order methods.
+  • LP-based Method: As Theorem 3.1 suggests, achieving small regret requires $\bm{Ax}-\bm{b}$ to be close to zero.
+To enforce this, we impose a stronger condition to track the average remaining resource—if $\bm{d}_{t}$ exceeds the allowed limit, we manually set the dual prices to zero to accept all subsequent orders.
+This approach enables us to eliminate the “stopping time” argument and make the analysis compatible with our new framework.
+  • First-order Method: The regret improvement comes from the final batch since it starts with the LP-derived learning result $\bm{p}_{kf}$, which lies within a $\mathcal{O}(1/\sqrt{kf})$-sized neighborhood around $\bm{p}^{*}$.
+Using the subgradient method, we bound the three components in ˜3.1 in terms of $\lVert\bm{p}_{t}-\bm{p}^{*}\rVert$ and express them in terms of $f,k,T,$ and $\alpha_{t}$.
+Optimizing the learning rate to $\alpha_{t}=f^{-2/3}$ yields an improved regret of $\mathcal{O}(f^{1/3})$ in the final batch.
+3.3 Algorithm Extension
+Figure 2: Algorithm 2 illustration of parallel paths with multi-time restart.
+Decisions are generated based on 1) the LP-based method (blue) with frequency $f$ and 2) the first-order method (red) during re-solving intervals with a warm start.
+We propose an enhanced version of Algorithm 1 that employs the first-order method for decision-making between two consecutive LP resolves.
+Rather than making decisions solely with the most recently solved dual price from the learning path, Algorithm 2 treats it as a new starting point for each re-solving interval and adopts a smaller step size for subgradient updates to avoid deviating too far from the LP-guided solutions.
+Figure 2 illustrates this framework.
+Specifically, for each interval $[jf,(j+1)f]$, we attain the new start $\bm{p}_{jf}$ from the learning path at time $jf$ and continue to fine-tune the dual price $\{\bm{p}_{t}\}_{t=jf}^{(j+1)f}$ using the first-order method with a learning rate $\alpha_{t}=\mathcal{O}(1/t)$.
+Algorithm 2 summarizes this approach.
+As we illustrate in our experiments (Section 4), the incorporation of an intermediate first-order method improves the algorithm’s stability and ensures smooth resource consumption.
+Therefore, the multi-restart mechanism results in better performance during the final batch and improves the algorithm’s total performance.
+Algorithm 2 Enhanced Multi-Start OLP Algorithm
+Input: total resource $\bm{b}$, horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$
+for $t=$ 11 to $T$ do
+   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ as rule (6) Update constraint for i=1,,mi=1,...,m:   remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$   average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ if $t\bmod f\neq 0$ then
+      Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$ Compute subgradient and update dual price $\bm{p}_{t+1}$:
+$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))^{+}$
+  end if
+ if $t\bmod f=0$ and constraints are not violated then
+      Solve OLP and update dual price $\bm{p}_{t+1}$:
+$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\tfrac{1}{t}\textstyle\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
+  end if
+end for
+3.4 Algorithm Application
+Our motivation in designing Algorithm 1 is to effectively balance computational efficiency and decision optimality.
+Building on the Spectrum Theorem 3.2, this section aims to translate our theoretical results into practical applications.
+We formulate a new optimization problem to find the optimal re-solving frequency that minimizes the regret in ˜3.2, taking into account computational resource capacities.
+Denote $c_{1}(\cdot),c_{2}(\cdot)$ as computational cost functions for LP-based and first-order methods, and $R$ as the total computational resource capacity.
+The optimal value of $f$ is determined by solving the following optimization problem:
+$\min_{f\in\{1,\dots,T\}}$ $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \log\Big(\tfrac{T}{f}\Big)+\sqrt{f}$
+s.t. $\penalty 10000\ \quad c_{1}(k)+2c_{2}(f)\leq R.$ (14)
+Specifically, if we use the interior-point method or the simplex method as the LP solver, the computational cost is $m^{2}(m+t)$ for any time $t$.
+The first-order method updates gradients in constant time.
+The following proposition provides a concrete example in practice.
+Proposition 3.6 (Optimal Re-solving Frequency).
+Given a fixed computation resource capacity $R$, if we use the interior-point or simplex method as the LP solver in Algorithm 1, we can instantiate Constraint (14) as
+$\textstyle\sum_{j=1}^{k}(m^{2}(m+jf))+2mf\leq R.$
+This proposition enables users to determine the optimal re-solving frequency that balances regret and computational cost based on available computational resources.
+4 Numerical Experiments
+Figure 3: Algorithms comparison.
+$T$ Regret Algorithm Compute Time (s)
+$0^{4} 11$5.32 $\mathcal{O}(T^{1/2})$ First-Order $0.008$
+60.39 $\mathcal{O}(T^{1/3})$ First-Order $0.013$
+3.50 LP-based $123.497$
+9.12 Algorithm 1 1.8
+5.09 Algorithm 2 1.8
+$10^{5}$ 203.20 $\mathcal{O}(T^{1/2})$ First-Order $0.118$
+73.07 $\mathcal{O}(T^{1/3})$ First-Order $0.108$
+3.88 LP-based >3600>3600
+9.41 Algorithm 1 56.9
+6.36 Algorithm 2 56.8
+$10^{6}$ 351.91 $\mathcal{O}(T^{1/2})$ First-Order $1.211$
+115.39 $\mathcal{O}(T^{1/3})$ First-Order $1.577$
+5.50 LP-based >100000>100000
+11.65 Algorithm 1 2155.9
+7.09 Algorithm 2 2242.1
+Figure 5: Regret for various algorithms.
+We conduct extensive experiments to evaluate our algorithm’s performance and validate our theoretical results.
+This section is divided into two parts.
+In the first part (Section 4.1), we evaluate Algorithms 1 and 2 across different choices of re-solving frequency.
+In the second part (Section 4.2), we compare our algorithm with LP-based and first-order methods in terms of regret and running time.
+All implementations can be found at GitHub Link.
+We consider the following distributions:
+$\text{Input I:}\ \ a_{it}\sim\text{Unif}[0,2],r_{t}\sim\text{Unif}[0,10]$
+$\text{Input II:}\ \ a_{it}\sim\mathcal{N}(0.5,1),r_{t}\sim\mathcal{N}(0.5m,m)$
+Learning rates are selected as specified in Section 3:
+$\text{Algorithm 1: }\alpha_{t}=\begin{cases}\mathcal{O}(1/f^{1/2})&t\leq f\\ \mathcal{O}(1/f^{2/3})&t\geq kf\end{cases}$
+$\text{Algorithm 2: }\alpha_{t}=\mathcal{O}(1/t)$
+4.1 Regret under Varying Re-solving Frequencies
+We choose $m=1$ and generate the sequence $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ randomly from the uniform (Input I) and normal (Input II) distributions, and include more complex distributions in Appendix E.1.
+The time horizon $T$ spans evenly over $10^{2},10^{6}]$, the initial average resource is sampled as $d_{i}\sim\text{Uniform}[1/3,2/3]$, and $f\in\{T^{1/3},T^{1/2},T^{2/3}\}$ representing high, medium, and low re-solving frequencies.
+We report the average result over 100 trials for each experiment.
+We use the classic first-order method with $\mathcal{O}(T^{1/2})$ regret (Li et al., 2020) as a baseline.
+Results are summarized in Table 1 and Figure 4 on a logarithmic scale.
+We analyze the absolute value of regret and its growth rate over time.
+We observe that regret decreases as the re-solving frequency increases.
+This trend holds consistently across both algorithms and input types.
+In addition, while regret accumulates over longer time horizons, the increasing rate remains stable for algorithms employing higher re-solving frequencies.
+These findings are consistent with the guarantees of ˜3.2, as more frequent updates enable better adaptation to dynamic environments.
+4.2 Comparative Analysis with Baseline Methods
+Table 1: Regret of algorithms with various re-solving frequencies.
+$T$ First-Order Low freq Mid freq High freq
+Input I $10^{3}$ 12.13 7.76 5.77 4.86
+$10^{4}$ 38.50 10.96 7.55 5.67
+$10^{5}$ 122.44 23.90 9.92 8.36
+$10^{6}$ 404.59 56.70 21.90 8.99
+Input II $0^{3} 11$.44 6.28 4.86 3.95
+$10^{4}$ 36.50 10.21 7.34 3.81
+$0^{5} 11$5.57 14.61 11.78 4.66
+$10^{6}$ 365.99 35.20 15.68 6.26
+(a) Algorithm 1
+$T$ First-Order Low freq Mid freq High freq
+Input I $10^{3}$ 12.13 6.78 5.20 4.50
+$10^{4}$ 38.50 10.37 8.03 5.99
+$10^{5}$ 122.44 22.33 11.57 6.36
+$10^{6}$ 404.59 48.21 22.44 7.09
+Input II $0^{3} 11$.44 3.20 2.56 1.75
+$10^{4}$ 36.50 5.48 4.30 2.52
+$0^{5} 11$5.57 12.35 4.48 3.86
+$10^{6}$ 365.99 30.48 13.20 4.77
+(b) Algorithm 2
+Figure 4: Evaluations of Algorithm˜1 and 2 across various horizons, re-solving frequencies, and stochastic inputs, validating the positive relationship between regret and frequency stated in ˜3.2.
+We next compare our algorithm’s regret and computation time with a classic LP-based (Li & Ye, 2022) and two first-order methods with $\mathcal{O}(T^{1/2})$ (Li et al., 2020) and $\mathcal{O}(T^{1/3})$ regrets (Gao et al., 2024).
+We generate $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ from a uniform distribution (Input I).
+We set the resource types to $m=5$, time $T$ to range evenly over $10^{2},10^{6}]$, average resource $d_{i}\sim\text{Uniform}[1/3,2/3]$, and re-solving frequency to $f=T^{1/3}$ from Section 4.1.
+Each result is averaged over 100 trial runs.
+We summarize the findings in Table 5 and Figure 5, and we include more comparisons with an infrequent re-solving method (Li et al., 2024) in Appendix E.2.
+We observe the effectiveness of Algorithm˜1 and 2:
+  1. 1.
+    Our algorithms exhibit strong performance in terms of decision optimality.
+They achieve over a 20-fold and 10-fold improvement in regret compared to the $\mathcal{O}(T^{1/2})$ first-order method and the $\mathcal{O}(T^{1/3})$ first-order method respectively.
+These numerical results also corroborate the theoretical bounds in Theorem 3.2.
+  2. 2.
+    Our algorithms are computationally efficient. Their runtimes exhibit over 100-fold improvement compared to the LP-based method with only a minimal increase (less than 2-fold) in regret.
+Therefore, we achieve a balance between effective decision-making and efficient computation.
+Our algorithms demonstrate better regret than the first-order method and obtain substantial computational speed-ups compared to the LP-based method.
+They exhibit strong scalability and adaptability across various re-solving frequencies and stochastic input models, consistently delivering superior performance as the problem size grows.
+5 Conclusion and Discussion
+This paper presents a new approach to online linear programming for dynamic resource allocation, addressing the inherent trade-offs between decision-making optimality and computational efficiency.
+Recognizing the limitations of existing methods, we propose a parallel framework that decouples online learning and decision-making into independent yet complementary processes.
+By integrating LP-based and first-order methods, our framework effectively balances total regret and computational cost.
+We establish rigorous theoretical guarantees, proving that our algorithm achieves a worst-case regret bound of $\mathcal{O}(\log(T/f)+\sqrt{f})$ under continuous support.
+This result highlights our method’s ability to interpolate between LP-based and first-order methods based on computational capability.
+Furthermore, extensive experiments validate the effectiveness of our approach, demonstrating improvements in both regret minimization and runtime efficiency over competitive baselines.
+Beyond these contributions, our work opens avenues for further research in adaptive online decision-making.
+Future directions include refining the re-solving frequency based on real-time computational constraints and extending our framework to broader classes of online optimization problems.
+Overall, our results underscore the potential of hybrid algorithms in high-dimensional, large-scale environments, offering practical insights for applications in operations research, machine learning, and beyond.
+Discussion of Unknown Horizon
+In this paper, we consider decision-making under a finite horizon, since the average resource ˜2.1(c) widely used in OLP literature relies on a fixed horizon to be well-defined.
+When the horizon is unknown, this assumption becomes ill-posed, and prior work (Balseiro et al., 2023) shows that it may not be possible to achieve sublinear regret.
+Addressing these challenges would likely require first adapting LP-based and first-order OLP methods to uncertain horizons before tackling a hybrid approach.
+In practice, it is often possible to make a prior prediction of the horizon using data-driven approaches or based on the resources available.
+For example, in online advertising, the number of customers can be predicted from historical statistics.
+Besides, in retail or event-based sales, the selling horizon may be externally decided by upper-level decision-makers.
+These practical applications suggest a reasonable modeling choice for finite time horizons.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Acknowledgments
+This work was supported in part by NSF grant CCF-2338226.
+Appendix
+Structure of the Appendix
+We organize the appendix as follows.
+In Section A, we present some primary results for the OLP problem and the foundation analysis for LP-based and first-order methods; Section B characterizes the unique properties from our parallel multi-phase structure, and presents the cohesive mathematical framework which decomposes the performance metrics to unify all the OLP methods; Section C demonstrates the outline of Algorithm˜2, combines all the previous results and proves the main results in our paper; Section D provides auxiliary results and technical support for the previous three sections; Section E includes supplementary experiments of more general input distributions and comparisons with recent algorithms.
+Appendix A Primary Results and Properties
+In this section, we present some primary results for the problem and OLP dual algorithms which will help present our main results.
+We also include basic properties and convergence of the LP-based and first-order methods.
+All the analyses are stated under Assumption 2.1 and 2.2.
+A.1 Preliminary results
+We propose the following lemma as directed results from Assumptions. Let $\Xi$ denote the family of distributions satisfying ˜2.1 and 2.2. We propose the following lemma.
+Lemma A.1 (Dimension Stability).
+Under Assumptions 2.1 and 2.2, there exists a constant $\delta>0$ such that
 […]
-This lemma is a consequence of Lemmas 12 and 13 in [16].
+This lemma is a consequence of Lemmas 12 and 13 in (Li & Ye, 2022).
 The existence of $\delta$ comes from the continuity of $f_{\bm{d}_{t}}(\bm{p})$.
@@
 We will analyze our feasible dual solutions derived from the online algorithm with $\bm{d}_{t}\in\mathcal{D}$.
-Lemma 2.2 (Quadratic Regularity, Proposition 2 in [16]).
-Under Assumptions 1 and 2, for any $\bm{p}\in\mathcal{V}_{\bm{p}}$,
+Lemma A.2 (Quadratic Regularity, Proposition 2 in (Li & Ye, 2022)).
+Under Assumptions 2.1 and 2.2, for any $\bm{p}\in\mathcal{V}_{\bm{p}}$,
 $f(\bm{p})$ $\leq f(\bm{p}^{*})+\nabla f(\bm{p}^{*})^{\top}(\bm{p}-\bm{p}^{*})+\frac{\mu\bar{a}^{2}}{2}\lVert\bm{p}-\bm{p}^{*}\rVert^{2},$
@@
 In later sections, we will leverage this result to derive our regret bound.
-3 Parallel Multi-Phase Online Learning and Decision-Making
-In this section, we present our novel framework for parallel multi-phase online learning and decision-making.
-Our approach is motivated by the difficulties of existing online resource allocation methods, which involve either high regret or heavy computational costs.
-Specifically, the LP-based AHDL Algorithm 1 has a strong worst-case regret bound of $\mathcal{O}(\log T)$, but its high computational costs delay decision-making.
-Meanwhile, the first-order online Algorithm 2 updates decisions efficiently but suffers a high regret of $\mathcal{O}(\sqrt{T})$.
-By combining the strengths of these two methods, our new framework balances decision-making quality and computational efficiency, thereby overcoming the limitations of existing approaches.
-3.1 Algorithm Design
-We establish our framework with the following design:
-  1. 1.
-    We maintain two parallel processes—one dedicated to online learning and the other to online decision-making.
-The online learning results are periodically sent to the decision-making path to guide subsequent updates.
-  2. 2.
-    In the online learning path, we employ an enhanced LP-based AHDL algorithm, which only re-solves the updated OLP problem according to a predefined frequency, significantly reducing the computational overhead.
-  3. 3.
-    In the online decision-making path, we apply the first-order method exclusively during the initial and final re-solving batches, where the learning rate is optimally tuned for these two intervals.
-Figure 1: Parallel Multi-Phase Online Learning and Decision-Making Algorithm
-Figure 1 illustrates our framework, which generates two parallel sequences of dual solutions: $\{\bm{p}_{t}^{D}\}_{t=1}^{T}$ from a first-order method and $\{\bm{p}_{t}^{L}\}_{t=1}^{T}$ from an LP-based approach.
-The algorithm proceeds in batches of length $f$.
-In the first batch, it uses a first-order method to iteratively update dual prices, thereby guiding the decision-making process.
-At the end of this batch, it applies the LP-based AHDL algorithm to obtain a refined set of dual prices, passing these updated values to the decision-making path.
-This LP-based re-solving occurs only once per batch, making the approach computationally efficient.
-Formally, we re-solve the OLP at every time $t$ satisfying:
-$t\leq kf\quad\text{and}\quad t\bmod f=0,$ (12)
-where $k=\left\lfloor\frac{T}{f}\right\rfloor$ represents the number of batches.
-After the first batch, the algorithm does not employ the first-order method until the final batch.
-Instead, during intermediate batches, the decision-making path uses the most recent dual prices computed from the LP at the end of the previous batch.
-Only in the final batch does the algorithm restart the first-order updates to guide the final set of decisions.
-Algorithm 3 summarizes this approach.
-Algorithm 3 Parallel Multi-Phase Online Learning and Decision-Making
-Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of re-solving batches $k=\lfloor T/f\rfloor$.
+A.2 LP-based analysis
+We state the LP-based method in Algorithm˜3 and its properties of boundedness and dual convergence results.
+Algorithm 3 LP-based method
+Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0}.$
 for $t$ = $1$ to $T$ do
-       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6).
-Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$.
-if $t\leq f$ or $t\geq kf$ then
-             Compute the subgradient and update the dual price $\bm{p}_{t+1}$:
-$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$ (13)
-       end if
-      if $t\bmod f=0$ and constraints are not violated then
-             Solve the online dual problem and update dual price $\bm{p}_{t+1}$ as
-$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$ (14)
-       end if
-      else
-             Update dual price $\bm{p}_{t+1}$ to be the most recent solution $\bm{p}_{t+1}=\bm{p}_{t}$.
-       end if
+   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6) if constraints are not violated.
+Update constraint for i=1,,mi=1,...,m:   remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$   average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Solve the updated dual problem and obtain dual price $\bm{p}_{t+1}$:
+$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
 end for
-We achieve a strong synergy between an effective online learning method and an efficient decision-making procedure.
-This structure enables the algorithm to remain responsive to dynamic environments and adapt to the latest information.
-Our strategy of periodic re-solving greatly reduces the computational costs.
-Therefore, we establish a “wait-less” online decision-making framework in which each customer’s order is processed immediately without waiting for the completion of earlier requests.
-3.2 Algorithm Analysis
-We next present several lemmas that characterize the convergence of the online dual solutions $\{\bm{p}_{t}\}_{t=1}^{T}$ generated by Algorithm 3 toward the optimal solution $\bm{p}^{*}$ of the stochastic program (3).
-In particular, let $\bm{p}_{t}^{D}$ be the dual prices updated via the first-order rule (13), and let $\bm{p}_{t}^{L}$ be those obtained from the LP-based method (14) at time $t$.
-These lemmas will then be used to establish upper bounds on the worst-case regret of our algorithm.
-Lemma 3.1 (Dual Price Boundedness).
-Under Assumption 1, the online and offline optimal dual prices are bounded respectively by:
-$\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}},\ \ \ \lVert\bm{p}_{t}\rVert\leq\bar{p}\coloneqq\max\left(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right).$ (15)
-This lemma establishes that the optimal dual prices remain bounded.
-Our algorithm maintains these bounds because if $\bm{p}_{t}$ grows large, the algorithm responds by accepting more orders, which in turn reduces $\bm{p}_{t+1}$.
-This self-correcting mechanism keeps the dual prices within these limits.
-Moreover, drawing on results of [15], we have that
-$\lVert\bm{p}_{t}^{D}\rVert\leq\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})$
-and from [16], we have that
-$\lVert\bm{p}_{t}^{L}\rVert\leq\frac{\bar{r}}{\underline{d}-\delta}.$
-Consequently, we can prove the following convergence lemmas.
-Lemma 3.2 (Dual Convergence of LP-based algorithm).
-Under Assumptions 1 and 2, there exists a constant $C_{lp}>0$ depending on $\bar{r},\bar{a},\underline{d},m,\nu$, and $\lambda$ such that
-$\mathbb{E}\left[\lVert\bm{p}_{t}^{L}-\bm{p}_{t}^{*}\rVert^{2}\right]\leq\frac{C_{lp}}{t}.$ (16)
-In addition, the difference between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ satisfies
-$\lVert\bm{p}_{t}^{*}-\bm{p}^{*}\rVert^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\lVert\bm{d}_{t}-\bm{d}\rVert^{2}.$ (17)
-This lemma establishes that the LP-based online AHDL algorithm [13] produces dual solutions $\bm{p}_{t}$ that converge to $\bm{p}_{t}^{*}$.
-This convergence highlights the stability of the online dual variables during the intermediate stages of decision-making and ensures a warm start for the first-order method in the last re-solving batch.
-Moreover, we bound the distance between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ by relating their respective average resource capacities, $\bm{d}_{t}$ and $\bm{d}$, in the associated stochastic programs [16].
-These results provide the foundation for analyzing $\lVert\bm{p}_{t}-\bm{p}^{*}\rVert$ in our final result.
-Lemma 3.3 (Dual convergence of First-order algorithm).
-Under Assumptions 1 and 2, if $\alpha_{t}<\nu\lambda$, the subgradient updates satisfy the following recursion rule:
-$\mathbb{E}\left[\lVert\bm{p}_{t+1}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq(1-\alpha_{t}\nu\lambda)\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}.$ (18)
-Case 1. if $\alpha_{t}\equiv\alpha<\frac{1}{\nu\lambda}$, then there exists a constant $C_{fo}=\frac{\bar{p}^{2}+m(\bar{a}+\bar{d})^{2}}{\nu\lambda}$ such that
-$\mathbb{E}\left[\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq C_{fo}\left(\frac{1}{\alpha t}+\alpha\right).$ (19)
-Case 2. if $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, then there exists a constant $C_{fo}=\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}}$ such that
-$\mathbb{E}\left[\lVert\bm{p}_{t}^{D}-\bm{p}^{*}\rVert^{2}\right]\leq\frac{C_{fo}}{t}.$ (20)
-This lemma shows that the first-order method [8] guarantees the convergence of $\bm{p}_{t}$ to $\bm{p}^{*}$.
-It also highlights a key trade-off in choosing the learning rate $\alpha_{t}$, an aspect that will be central to our later optimality analysis.
-With this groundwork in place, we now move on to bounding the total regret of our algorithm.
-Theorem 3.1 (Total Regret).
-Under Assumptions 1 and 2, the total regret $\Delta_{T}$ of Algorithm 3 satisfies:
-$\Delta_{T}\ \leq\ \mu\bar{a}^{2}\sum_{t=1}^{T}\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]+\lVert\bm{p}^{*}\rVert\cdot\mathbb{E}\left[\lVert(\bm{b}-\bm{Ax})^{B+}\rVert\right]+\mathbb{E}\left[\lVert(\bm{Ax}-\bm{b})^{+}\rVert\right].$ (21)
-where $(\cdot)^{B+}$ indicates that we only consider the binding terms and project them onto the positive orthant.
-We express the total regret as the sum of two key components: dual convergence and resource consumption.
-This bound demonstrates that the algorithm encourages (1) smooth and balanced resource utilization, especially for the binding terms, and (2) full resource consumption by the end of the time horizon.
-Building on Theorem 3.1, we next introduce a “spectrum theorem” for the algorithm’s total regret.
-Theorem 3.2 (Spectrum Theorem).
-Under Assumptions 1 and 2, the worst-case regret $\Delta_{T}$ of Algorithm 3 is bounded by
-$\Delta_{T}\in\mathcal{O}\left(\sqrt{f}+\log\left(\frac{T}{f}\right)+f^{1/3}\right)=\mathcal{O}\left(\log\left(\frac{T}{f}\right)+\sqrt{f}\right),$ (22)
-where $T/f$ represents the number of re-solving batches and $f$ is the length of each batch.
-Remark 3.1.
-The Spectrum Theorem elucidates the trade-offs in total regret induced by our use of LP-based and first-order methods.
-As stated in Section 2, utilizing pure LP-based method at every step yields a regret of $\mathcal{O}(\log T)$, while applying pure first-order method continuously attains a regret of $\mathcal{O}(\sqrt{T})$.
-Our algorithm sets a re-solving frequency $f\in[1,T]$ and achieves the regret of $\mathcal{O}(\log(T/f)+\sqrt{f})$, which aligns with previous results with the extreme cases of $f=1$ and $f=T$.
-By selecting an appropriate re-solving frequency $f$, our algorithm leverages the strengths of both approaches.
-Specifically, first-order method gives us the regret of $\sqrt{f}$ for the first batch and $f^{1/3}$ for the last batch, and LP-based method produces the regret of $\log(T/f)$.
-Remark 3.2.
-We can obtain a tighter bound of
-$\Delta_{T}\in\mathcal{O}(\log(T)+f^{1/3})$ (23)
-if we have a warm start of the initial dual price for the first batch satisfying $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ or we use the LP-based method to learn at each time.
-Both these approaches can give us the improvement from $\sqrt{f}$ to $f^{1/3}$ for the first-batch regret.
-Technical Intuitions.
-We prove this theorem using several novel techniques.
-As stated in Theorem 3.1, we develop a comprehensive framework that accommodates both the first-order and LP-based methods, allowing us to decompose the total regret into distinct components associated with each approach.
-Specifically:
-  • For the first-order method, we relate the distance $\bm{Ax}-\bm{b}$ to the distance between dual prices.
-  • For the LP-based method, as the algorithm progresses toward the end of the time horizon, if the current average resource consumption rate $\bm{d}_{t}$ exceeds the allowed limit, we manually set the dual prices to zero to accept all future orders.
-This approach enables us to adapt the concept of a “stopping time” within our new framework and derive a rigorous bound of regret.
-Through this new framework, we integrate the LP-based and subgradient-based methods to achieve “wait-less” decision-making, where each customer’s order is processed immediately without delays, while also attaining the best-known logarithmic regret bound.
-3.3 Algorithm Extension
-Building on Algorithm 3, we propose an enhanced multi-stage learning algorithm that employs the first-order method for decision-making during each re-solving interval.
-This multi-restart mechanism further improves the accuracy and efficiency of our framework.
-We illustrate this algorithm in the following figure.
-Figure 2: Enhanced Parallel Multi-Phase Online Learning and Decision-Making Algorithm
-Figure 2 illustrates this variant of Algorithm 3 with multi-stage exploration and exploitation.
-Rather than make decisions solely based on the most recently solved dual prices, the algorithm treats the updated dual prices at each re-solving interval as a new starting point for the first-order learning process.
-For example, in Algorithm 3, the dual price $\bm{p}_{f}^{L}$ is used to guide decision-making during the time interval $t\in[f,2f]$.
-In contrast, the enhanced algorithm restarts the first-order method by setting $\bm{p}_{f}^{L}=\bm{p}_{f}^{D}$ and then continues to fine-tune the dual prices $\{\bm{p}_{t}^{D}\}_{t=f}^{2f}$ to make up-to-date decisions.
-Algorithm 4 summarizes this approach.
-Algorithm 4 Enhanced Parallel Multi-Phase Online Learning and Decision-Making
-Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of re-solving batches $k=\lfloor T/f\rfloor$.
-for $t=$ $1$ to $T$  do
-       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6).
-Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$.
-if $t\bmod f\neq 0$ then
-             Compute the subgradient and update dual price $\bm{p}_{t+1}$:
-$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$
-       end if
-      if $t\bmod f=0$ and constraints are not violated then
-             Solve the online dual problem and update dual price $\bm{p}_{t+1}$:
-$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
-       end if
-end for
-This algorithm establishes an online framework of multi-stage learning and decision-making.
-The incorporation of an intermediate first-order method improves the algorithm’s stability and ensures smooth resource consumption.
-These improvements result in better performance during the final batch, thereby reducing the algorithm’s total regret.
-3.4 Algorithm Application
-Our motivation in designing Algorithm 3 is to develop a hybrid decision-making system that effectively balances computational efficiency and decision accuracy.
-By integrating the first-order method, we aim to reduce computational costs, while the LP-based method enhances the precision of our decisions.
-Building on the Spectrum Theorem 3.2, this section aims to translate our theoretical results into practical applications.
-Specifically, we formulate a new optimization problem to determine the optimal re-solving frequency, taking into account computational resource constraints.
-We use $c_{1}(\cdot),c_{2}(\cdot)$ to denote the computational cost functions for the LP-based method and first-order method, respectively, and use $R$ to denote the total computational resource capacity.
-Then we define the following optimization problem to determine the optimal re-solving frequency $f$:
-$\min_{f\in\mathbb{Z}^{+}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \log\Big{(}\frac{T}{f}\Big{)}+\sqrt{f}$ (24)
-subject to $\leavevmode\nobreak\ \quad c_{1}(k)+2c_{2}(f)\leq R$
-$\leavevmode\nobreak\ \quad 1\leq f\leq T$
-We can further analyze the computational requirements of our algorithm by considering the specific LP solvers employed.
-Recall that $m$ denotes the dimension of the dual price vector.
-At any given time step $t$, utilizing either the interior-point method or the simplex method as the LP-based solver incurs a computational cost of $m^{2}(m+t)$.
-In contrast, the first-order method incurs a computational cost of $2mf$.
-We encapsulate these trade-offs in the following proposition.
-Proposition 3.1 (Optimal Re-solving Frequency).
-Given a fixed computational resource capacity $R$, if we use the interior-point method or simplex method as the LP solver in Algorithm 3, we can formulate the optimization problem (24) as
-$\min_{f\in\mathbb{Z}^{+}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \log\Big{(}\frac{T}{f}\Big{)}+\sqrt{f}$ (25)
-subject to $\leavevmode\nobreak\ \quad\sum_{b=1}^{k}\left(m^{2}(m+bf)\right)+2mf\leq R,$
-$\leavevmode\nobreak\ \quad 1\leq f\leq T$ (26)
-By solving optimization problem (25), we can select an optimal re-solving frequency $f$ that achieves the lowest regret within the computational resource limit $R$.
-Therefore, we extend our original optimization problem for resource allocation to include computational resource allocation.
-This proposition enables users to determine the optimal re-solving frequency that balances regret with computational cost based on their available computational resources.
-4 Numerical Experiments
-We conduct extensive experiments under various settings to illustrate the performance of our algorithm and validate the theoretical results of our framework.
-This section is divided into two parts.
-In the first part (Section 4.1), we evaluate Algorithms 3 and 4 across different choices of the re-solving frequency.
-In the second part (Section 4.2), we compare our algorithm with the LP-based method in terms of running time and with the first-order method in terms of regret.
-Both sets of experiments are conducted under different distributions of $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ that satisfy the assumptions in Section 2.
-4.1 Performance Evaluation under Varying Re-solving Frequencies
-We generate the sequence $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ randomly from distributions that satisfy Assumptions 1 and 2.
-Specifically, we use the uniform and normal distribution, with details in Table 2.
-We select the types of resource $m\in\{1,5\}$, time horizon $T$ evenly spanned over $[10^{2},10^{6}]$, and initial average resource capacity $d_{i}$ sampled i.i.d.
-from $\text{Uniform}[1/3,2/3]$.
-We run our algorithm for $100$ trials and use the average as the final result for each experiment.
-Table 2: Stochastic Input in Experiments
-Distribution Model $\bm{a}_{t}$ $r_{t}$
-Input I $a_{it}\sim\text{Uniform}[0,2]$ $r_{t}\bot\bm{a}_{t},r_{t}\sim\text{Uniform}[0,10]$
-Input II $a_{it}\sim\text{Normal}(0.5,1)$ $r_{t}=\sum_{i=1}^{m}a_{it},r_{t}\sim\text{Normal}(0.5m,m)$
-We evaluate our algorithm with $m=1$ under the stochastic input models I and II from Table 2.
-The experiments are conducted across various choices of $f\in\{T^{1/3},T^{1/2},T^{2/3}\}$ representing high, medium, and low re-solving frequencies.
-We analyze the absolute value of the regret and its growth rate over time.
-We illustrate our results in Table 3 and Figure 3 on a logarithmic scale.
-Table 3: Regret of Algorithms with various re-solving frequency
-Offline Low freq Mid freq High freq
-$T=10^{3}$ 12.13 7.76 5.77 4.86
-$T=10^{4}$ 38.50 10.96 7.55 5.67
-$T=10^{5}$ 122.44 23.90 9.92 8.36
-$T=10^{6}$ 404.59 56.70 21.90 8.99
-(a) Algorithm 3, Input I
-Offline Low freq Mid freq High freq
-$T=10^{3}$ 11.44 6.28 4.86 3.95
-$T=10^{4}$ 36.50 10.21 7.34 3.81
-$T=10^{5}$ 115.57 14.61 11.78 4.66
-$T=10^{6}$ 365.99 35.20 15.68 6.26
-(b) Algorithm 3, Input II
-Offline Low freq Mid freq High freq
-$T=10^{3}$ 12.13 6.78 5.20 4.50
-$T=10^{4}$ 38.50 10.37 8.03 5.99
-$T=10^{5}$ 122.444 22.33 11.57 6.36
-$T=10^{6}$ 404.59 48.21 22.44 7.09
-(c) Algorithm 4, Input I
-Offline Low freq Mid freq High freq
-$T=10^{3}$ 11.44 3.20 2.56 1.75
-$T=10^{4}$ 36.50 5.48 4.30 2.52
-$T=10^{5}$ 115.57 12.35 4.48 3.86
-$T=10^{6}$ 365.99 30.48 13.20 4.77
-(d) Algorithm 4, Input II
-(a) Algorithm 3 with input I
-(b) Algorithm 3 with input II
-(c) Algorithm 4 with input I
-(d) Algorithm 4 with input II
-Figure 3: Main Algorithms with Various Frequency
-These experiments assess the performance of our two main algorithms under varying re-solving frequencies and across multiple time horizons.
-As presented in Table 3 and Figure 3, we observe that regret decreases as the re-solving frequency increases from $f=T^{2/3}$ to $f=T^{1/3}$.
-This trend holds consistently across both algorithms and input types, with higher re-solving frequencies achieving the lowest regret in all cases.
-In addition, while regret accumulates over longer time horizons, the rate of increase remains stable for algorithms employing higher re-solving frequencies.
-These findings illustrate the robustness of our algorithms, as more frequent updates enable better adaptation to dynamic environment.
-4.2 Comparative Analysis with Baseline Methods
-After examining the algorithm under various re-solving frequencies, we compare its performance with the classic first-order method and LP-based method in terms of regret and computation time.
-We generate $\{(r_{t},\bm{a}_{t}\}_{t=1}^{T}$ from the uniform distribution under the Input I model (Table 2).
-We set the number of resource types to $m=5$, the time horizon $T$ to range evenly over $[10^{2},10^{6}]$, and the average resource capacity $d_{i}$ to be sampled i.i.d from $\text{Uniform}[1/3,2/3]$.
-The re-solving frequency is fixed to be $f=T^{1/3}$.
-Each result is averaged over 100 trial runs, and the findings are summarized in Table 5 and Figure 5.
-$T$ Regret Algorithm Compute Time(s)
-$10^{3}$ $66.88$ Offline $0.001$
-$2.80$ LP-based (Algorithm 1) $4.185$
-$25.11$ First-Order (Algorithm 2) $0.002$
-7.20 Algorithm 3 0.3
-4.50 Algorithm 4 0.3
-$10^{4}$ 115.32 Offline $0.008$
-3.50 LP-based (Algorithm 1) $123.497$
-60.39 First-Order (Algorithm 2) $0.013$
-9.12 Algorithm 3 1.8
-5.09 Algorithm 4 1.8
-$10^{5}$ 203.20 Offline $0.118$
-3.88 LP-based (Algorithm 1) $>3600$
-73.07 First-Order (Algorithm 2) $0.108$
-9.41 Algorithm 3 56.9
-6.36 Algorithm 4 56.8
-$10^{6}$ 351.91 Offline 1.211
-5.50 LP-based (Algorithm 1) $>100000$
-115.39 First-Order (Algorithm 2) 1.577
-11.65 Algorithm 3 2155.9
-7.09 Algorithm 4 2242.1
-Table 4: Algorithms Comparison.
-Figure 5: Regret for Various Algorithms
-Table 5 and Figure 5 provide a comprehensive comparison of all the algorithms discussed in this paper. Focusing on our primary algorithms, Algorithm 3 and 4, we observe:
-  1. 1.
-    Our algorithms exhibit strong performance in terms of decision optimality.
-They achieve over a 20-fold improvement in regret compared to the offline solution and more than a $10$-fold improvement compared to the first-order method.
-These numerical results also corroborate the theoretical bounds in Theorem 3.2.
-  2. 2.
-    Our algorithms are computationally efficient.
-Their computation times exhibit a 100-fold improvement compared to the LP-based method while maintaining regret levels that are remarkably close to those of the LP-based method.
-This significant improvement in efficiency makes our algorithms highly suitable for real-time and large-scale applications.
-Therefore, we achieve a balance between effective decision-making and efficient computation.
-Our algorithms demonstrate better regret than the first-order method and obtain substantial computational speed-ups compared to the LP-based method.
-Our algorithms exhibit robust scalability and adaptability under various re-solving frequencies and stochastic input models, consistently maintaining superior performance as the problem size grows.
-This adaptability ensures their effectiveness in diverse scenarios.
-5 Conclusion
-This paper addresses an online linear programming problem for decision-making in resource allocation.
-Identifying existing methods’ limitations, we develop a novel two-path framework that decouples online learning and decision-making into independent processes.
-Our cohesive framework integrates the LP-based method and first-order method, thereby balancing decision-making optimality and computational efficiency.
-We prove that our algorithm attains $\mathcal{O}(\log(T/f)+\sqrt{f})$ worst-case regret under continuous distributions, and our experimental results demonstrate strong regret and runtime improvements over competitive baselines.
-Appendix
-[…]
-Structure of the Appendix
-We organize the appendix as follows.
-In Section A, we introduce the foundation analysis for LP-based method and first-order method for the discussion of our framework; Section B presents the cohesive mathematical framework which unify all the OLP methods, and the unique properties derived from our parallel multi-phase algorithm; Section C combines all the previous results and proves the main results in our paper; Section D provides some auxiliary results and technical support in the previous three parts.
-Appendix A Primary Results and Properties
-In this section, we provide primary results of LP-based method and first-order method.
-We also include some basic analysis of dual algorithms that will help present our main results.
-We start from the convergence of regular LP-based method.
-All the analysis are under Assumption (1) and (2).
-A.1 LP-based Analysis
-We provide the analysis for dual solution obtained from LP-based method, so $\bm{p}_{t}=\bm{p}_{t}^{L}$ in this part.
+LP-based method incorporates the past decisions into the optimization of $\bm{p}_{t}$.
+If resources were over-utilized in earlier periods, the remaining supply decreases, prompting the algorithm to raise the dual price and become more selective with future orders.
+Conversely, if ample resources remain, the future dual price will be lowered, allowing more consumer requests to be accepted.
+This adaptive mechanism accounts for past actions by adjusting the available resource capacity.
 Lemma A.3 (Boundedness of LP result).
@@
 $\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}},\ \ \ \lVert\bm{p}_{t}\rVert\leq\frac{\bar{r}}{\underline{d}-\delta}.$
-Proof. By the optimality of $\bm{p}^{*}$ and boundedness of $\bm{d}$, we have
+Lemma A.4 (Dual Convergence of LP-based algorithm).
+Under Assumptions 2.1 and 2.2, $\bm{p}_{t}$ represents the online solution from LP-based method, there exists a constant $C_{lp}>0$ depending on $\bar{r},\bar{a},\underline{d},m,\nu$, and $\lambda$ such that
+$\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}_{t}^{*}\rVert^{2}\right]\leq\frac{C_{lp}}{t}.$ (15)
+In addition, the difference between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ satisfies
+$\lVert\bm{p}_{t}^{*}-\bm{p}^{*}\rVert^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\lVert\bm{d}_{t}-\bm{d}\rVert^{2}.$ (16)
+This lemma establishes that the LP-based online AHDL algorithm (Jiang et al., 2022) produces dual solutions $\bm{p}_{t}$ that converge to $\bm{p}_{t}^{*}$.
+This convergence highlights the stability of the online dual variables during the intermediate stages of decision-making and ensures a warm start for the first-order method in the last re-solving batch.
+Moreover, we bound the distance between $\bm{p}_{t}^{*}$ and $\bm{p}^{*}$ by relating their respective average resource capacities, $\bm{d}_{t}$ and $\bm{d}$, in the associated stochastic programs (Li & Ye, 2022).
+These results provide the foundation for analyzing $\lVert\bm{p}_{t}-\bm{p}^{*}\rVert$ in our final result.
+Proof of ˜A.3.
+By the optimality of $\bm{p}^{*}$ and boundedness of $\bm{d}$, we have
 $\underline{d}\lVert\bm{p}^{*}\rVert_{1}\leq\bm{d}^{\top}\bm{p}^{*}\leq\mathbb{E}[r]\leq\bar{r}.$
@@
 Given the non-negativeness of $\bm{p}^{*}$ and $\bm{p}^{*}\in\mathcal{V}_{\bm{d}}$, we obtain $\lVert\bm{p}^{*}\rVert\leq\lVert\bm{p}^{*}\rVert_{1}$, and thus $\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}}$.
-Similarly, by the optimality of $\bm{p}_{t}^{*}$ and its associated $\bm{d}_{t}\in\mathcal{D}$ in Lemma 2.1, we know
+Similarly, by the optimality of $\bm{p}_{t}^{*}$ and its associated $\bm{d}_{t}\in\mathcal{D}$ in Lemma A.1, we know
 $(\underline{d}-\delta)\lVert\bm{p}_{t}\rVert_{1}\leq{\bm{d}}_{t}^{\top}\bm{p}_{t}\leq\mathbb{E}[r]\leq\bar{r},$
 so we get the bound as $\lVert\bm{p}_{t}\rVert\leq\frac{\bar{r}}{\underline{d}-\delta}$.
-Proof of Lemma 3.2.
-According to the latest results in [13], the online dual price $\bm{p}_{t}$ achieves a sublinear convergence $\mathcal{O}(\frac{1}{\sqrt{t}})$ to $\bm{p}_{t}^{*}$. Since
+Proof of ˜A.4.
+According to the latest results in (Jiang et al., 2022), the online dual price $\bm{p}_{t}$ achieves a sublinear convergence $\mathcal{O}(\frac{1}{\sqrt{t}})$ to $\bm{p}_{t}^{*}$. Since
 $\bm{p}^{*}\in\arg\min_{\bm{p}\geq\bm{0}}f(\bm{p}):=\bm{d}^{\top}\bm{p}+\mathbb{E}\left[(r-\bm{a}^{\top}\bm{p})^{+}\right],$
 $\bm{p}_{t}^{*}\in\arg\min_{\bm{p}_{t}\geq\bm{0}}f_{\bm{d}_{t}}(\bm{p}_{t}):=\bm{d}_{t}^{\top}\bm{p}_{t}+\mathbb{E}\left[(r-\bm{a}^{\top}\bm{p}_{t})^{+}\right],$
-By Lemma 12 in [16], we have $\|\bm{p}_{t}^{*}-\bm{p}^{*}\|_{2}^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\|\bm{d}_{t}-\bm{d}\|_{2}^{2}$.
-A.2 First-order Analysis
-We provide the analysis for dual solution obtained from LP-based method, so $\bm{p}_{t}=\bm{p}_{t}^{D}$ in this part.
+By Lemma 12 in (Li & Ye, 2022), we have $\|\bm{p}_{t}^{*}-\bm{p}^{*}\|_{2}^{2}\leq\frac{1}{\nu^{2}\lambda^{2}}\|\bm{d}_{t}-\bm{d}\|_{2}^{2}$.
+A.3 First-order Analysis
+We present the classic first-order method (Li et al., 2020) in Algorithm˜4 and the first-order method with restart strategy (Gao et al., 2024) in Algorithm˜5.
+We also show their properties of boundedness and dual convergence results.
+Algorithm 4 First-Order Online algorithm
+Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0}.$
+for $t$ = $1$ to $T$ do
+   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (5). Update learning rate $\alpha_{t}=\mathcal{O}(1/\sqrt{T})$. Compute subgradient and obtain dual price $\bm{p}_{t+1}$:
+$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ (\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}+\frac{1}{2\alpha_{t}}\lVert\bm{p}-\bm{p}_{t}\rVert^{2}$
+end for
+Algorithm 5 First-Order Restart algorithm
+Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, and initial dual price $\bm{p}_{1}=\bm{0},\bm{p}^{L}_{1}=\bm{0}$
+for $t$ = $1$ to $T$ do
+   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6) Update learning rate $\alpha_{t}=\begin{cases}\mathcal{O}(1/T^{1/3})&t\leq T^{2/3}\\ \mathcal{O}(1/T^{2/3})&t>T^{2/3}\end{cases}$ Compute subgradient and obtain dual price $\bm{p}_{t+1}$:
+$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ (\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}+\frac{1}{2\alpha_{t}}\lVert\bm{p}-\bm{p}_{t}\rVert^{2}$
+Run subgradient method with stepsize $\alpha_{t}=\mathcal{O}(1/t)$ and update $\{\bm{p}_{t}^{L}\}$. At $t=T^{2/3}$, restart $\bm{p}_{T^{2/3}}=\bm{p}^{L}_{T^{2/3}}$.
+end for
 Lemma A.5 (Boundedness of first-order result).
@@
 $\lVert\bm{p}_{t}\rVert\leq\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d}).$
-Proof. As our initial choice $\bm{p}_{1}=\bm{0}$, according to Lemma 1 in [15], we get the above result. In addition, by Lemma B.1 in [8], we have
+Lemma A.6 (Dual convergence of First-order algorithm).
+Under Assumptions 2.1 and 2.2, $\bm{p}_{t}$ represents the online solution from first-order method, if $\alpha_{t}<\nu\lambda$, the subgradient updates satisfy the following recursion rule:
+$\mathbb{E}\left[\lVert\bm{p}_{t+1}-\bm{p}^{*}\rVert^{2}\right]$ $\leq(1-\alpha_{t}\nu\lambda)\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\alpha_{t}^{2}m(\bar{a}+\bar{d})^{2}.$ (17)
+Case 1. if $\alpha_{t}\equiv\alpha<\frac{1}{\nu\lambda}$, then there exists a constant $C_{fo}=\frac{\bar{p}^{2}+m(\bar{a}+\bar{d})^{2}}{\nu\lambda}$ such that
+$\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]\leq C_{fo}\left(\frac{1}{\alpha t}+\alpha\right).$ (18)
+Case 2. if $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, then there exists a constant $C_{fo}=\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}}$ such that
+$\mathbb{E}\left[\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\right]\leq\frac{C_{fo}}{t}.$ (19)
+This lemma shows that the first-order method (Gao et al., 2024) guarantees the convergence of $\bm{p}_{t}$ to $\bm{p}^{*}$.
+It also highlights a key trade-off in choosing the learning rate $\alpha_{t}$, an aspect that will be central to our later optimality analysis.
+It also highlights a key trade-off in choosing the learning rate $\alpha_{t}$, an aspect that will be central to our later optimality analysis.
+With this groundwork in place, we now move on to bounding the total regret of our algorithm.
+Proof of ˜A.5.
+As our initial choice $\bm{p}_{1}=\bm{0}$, according to Lemma 1 in (Li et al., 2020), we get the above result. In addition, by Lemma B.1 in (Gao et al., 2024), we have
 $\lVert\bm{p}_{t}\rVert\leq\frac{\bar{r}}{\underline{d}}+\frac{m(\bar{a}+\bar{d})^{2}\alpha_{t}}{2\underline{d}}+\alpha_{t}\sqrt{m}(\bar{a}+\bar{d}).$
-Proof for Lemma 3.3.
+Proof of ˜A.6.
 Based on the updated rule in Algorithm 4, we derive:
@@
 $f(\bm{p}^{*})\geq f(\bm{p}_{t})+\langle\bm{d}-\bm{a}_{t}x_{t},\bm{p}_{t}-\bm{p}^{*}\rangle.$
-2) With quadratic regularity of $f$ as in Lemma 2.2, we have:
+2) With quadratic regularity of $f$ as in Lemma A.2, we have:
 $f(\bm{p}_{t})\geq f(\bm{p}^{*})+\nabla f(\bm{p}^{*})^{\top}(\bm{p}_{t}-\bm{p}^{*})+\frac{\nu\lambda}{2}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2},$
@@
 Since all feasible $\bm{p}\in\mathcal{V}_{\bm{p}}$ is bounded, let $C_{fo}=\frac{\bar{p}^{2}+m(\bar{a}+\bar{d})^{2}}{\nu\lambda}$, we are able to obtain (18).
-Case 2. When $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, for any $j\leq t$, Lemma 3.3 gives us
+Case 2. When $\alpha_{t}=\frac{2}{\nu\lambda(t+1)}$, for any $j\leq t$, Lemma A.6 gives us
 $\mathbb{E}\lVert\bm{p}_{j+1}-\bm{p}^{*}\rVert^{2}\leq\frac{j-1}{j+1}\lVert\bm{p}_{j}-\bm{p}^{*}\rVert^{2}+\frac{4m(\bar{a}+\bar{d})^{2}}{\nu^{2}\lambda^{2}(j+1)^{2}}.$
@@
 Appendix B Algorithm Design and Analysis
-In this section, we present the cohesive mathematical framework to integrate the LP-based method and first-order method by developing a new performance metric.
+In this section, we present a cohesive mathematical framework to integrate the LP-based method and first-order method by developing a new performance metric.
 We also provide some unique properties for our algorithm design of parallel multi-phase structure.
-As stated in Algorithm 3, we have independent two paths for online learning and online decision-making conducted simultaneously.
-Specifically, for the online learning path, we use the LP-based method to resolve the updated linear programs at a fixed frequency and update this result to the decision-making path; For online decision-making path, we restart this process every time when receiving an updated learning result.
-We apply first-order method in the initial and final batches, and use the latest dual price for decision-making in the intermediate resolving intervals.
-This process is illustrated in Figure (1).
-B.1 Optimality Gap Analysis
+As stated in Algorithm 1, we construct independent paths for online learning and online decision-making conducted simultaneously.
+Specifically, for the online learning path, we use the LP-based method to resolve updated linear programs at a fixed frequency and send this result to the decision-making path; For the online decision-making path, we apply the first-order method in the initial and final batches, and use the latest dual price for decision-making in the intermediate resolving intervals.
+This process is illustrated in Figure˜1.
+B.1 Regret Analysis
 First, we construct an upper bound for the offline optimal objective value.
@@
 Lemma B.1 (Lagrangian Upper Bound).
-Under Assumptions (1) and (2), define the deterministic Lagrangian dual function as
+Under Assumptions (2.1) and (2.2), define the deterministic Lagrangian dual function as
 $\ell(\bm{p})\coloneqq\mathbb{E}\left[rI(r>\bm{a}^{\top}\bm{p})+(\bm{d}-\bm{a}I(r>\bm{a}^{\top}\bm{p}))^{\top}\bm{p}^{*}\right].$
@@
 $(b)\quad$ $\ell(\bm{p}^{*})-\ell(\bm{p})\leq\mu\bar{a}^{2}\lVert\bm{p}-\bm{p}^{*}\rVert^{2}.$ (20)
-Proof. This lemma is proved as Lemma 3 in [16] by strong duality and optimality of $\bm{p}_{T}^{*}$.
-Theorem B.1 (Decomposition of Optimality Gap).
-Under Assumptions (1) and (2), let $C_{r}=\max\{\frac{\bar{r}}{\underline{d}},\frac{\nu\lambda\bar{a}^{2}}{2}\}$, we derive an upper bound for the optimality gap $r(\bm{x})$ as:
+Lemma B.2 (Dual Price Boundedness).
+Under Assumption 2.1, the online and offline optimal dual prices are bounded respectively by $\lVert\bm{p}^{*}\rVert\leq\tfrac{\bar{r}}{\underline{d}}$ and
+$\lVert\bm{p}_{t}\rVert\leq\bar{p}\coloneqq\max\bigg(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\bigg).$
+This lemma establishes that the optimal dual prices remain bounded.
+Our algorithm maintains these bounds because if $\bm{p}_{t}$ grows large, the algorithm responds by accepting more orders, which in turn reduces $\bm{p}_{t+1}$.
+This self-correcting mechanism keeps the dual prices within these limits.
+With this groundwork in place, we now move on to derive the upper bound for the total performance of Algorithm˜1.
+Theorem B.3 (Decomposition of Regret).
+Under Assumptions (2.1) and (2.2), let $C_{r}=\max\{\frac{\bar{r}}{\underline{d}},\frac{\nu\lambda\bar{a}^{2}}{2}\}$, we derive an upper bound for the regret $r(\bm{x})$ as:
 $r(\bm{x})\leq C_{r}\bigg[\mathbb{E}\Big\lVert\Big(\bm{b}-\sum_{t=1}^{T}\bm{a}_{t}x_{t}\Big)^{B^{+}}\Big\rVert+\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}\bigg].$ (21)
-Proof. By the definition of $\ell(\bm{p})$, we derive the online objective as
+Proof of ˜B.1.
+This lemma is proved as Lemma 3 in (Li & Ye, 2022) by strong duality and optimality of $\bm{p}_{T}^{*}$.
+Proof of Lemma B.2.
+By the boundedness results in Lemma A.3 and Lemma A.5, define
+$\bar{p}\coloneqq\max\left(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right),$
+then we complete the proof.
+Proof of ˜B.3.
+By the definition of $\ell(\bm{p})$, we derive the online objective as
 $\mathbb{E}\bigg[\sum_{t=1}^{T}r_{t}x_{t}\bigg]$ $=\sum_{t=1}^{T}\mathbb{E}[\mathbb{E}[r_{t}x_{t}|\bm{p}_{t}]]$
 $=\sum_{t=1}^{T}\mathbb{E}[\ell(\bm{p}_{t})-(\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}^{*}].$
-Then by Lemma B.1, we derive the upper bound for the optimality gap $r(\bm{x})$ as defined in (8) to be:
+Then by Lemma B.1, we derive the upper bound for the regret $r(\bm{x})$ as defined in (8) to be:
 $r(\bm{x})$ $=\mathbb{E}\bigg[\sum_{t=1}^{T}r_{t}x_{t}^{*}-r_{t}x_{t}\bigg]$
@@
 As $\lVert\bm{p}^{*}\rVert\leq\frac{\bar{r}}{\underline{d}}$ is bounded, taking the constant $C_{r}=\max\{\frac{\bar{r}}{\underline{d}},\frac{\mu\bar{a}^{2}}{2}\}$ completes the proof.
-B.2 Total Regret Analysis
+B.2 Total Performance Analysis
 We consider the constraint violation as $v(\bm{x})=\lVert(\bm{Ax}-\bm{b})^{+}\rVert$ in (9).
-Then the total regret is the combination of optimality gap and constraint violation.
+Then the total performance is the combination of regret and constraint violation.
 Since Theorem B.3 holds for any unknown distribution $\mathcal{P}\in\Xi$, we have:
@@
 $\leq C_{r}\bigg[\sum_{t=1}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}+\mathbb{E}\big\lVert(\bm{b}-\bm{Ax})^{B^{+}}\big\rVert+\mathbb{E}\big\lVert(\bm{Ax}-\bm{b})^{+}\big\rVert\bigg].$ (22)
-This new framework is adaptive to all OLP methods. We will proceed based on this structure to analyze various methods respectively.
+This new framework is adaptive to all OLP methods. We proceed with this structure to analyze various OLP methods.
 Theorem B.4 (Horizon Division).
-Based on the change of methods, we separate the horizon into three intervals $T_{1},T_{2}$, and $T_{3}$, and reorganize the total regret for analysis to be
+Based on the change of methods, we separate the horizon into three intervals $T_{1},T_{2}$, and $T_{3}$, and reorganize the total performance for analysis to be
 $\Delta_{T}\leq C_{r}\big[\Delta_{T_{1}}+\Delta_{T_{2}}+\Delta_{T_{3}}\big].$ (23)
-where $\Delta_{T_{1}},\Delta_{T_{2}}$, and $\Delta_{T_{3}}$ are the total regret during each interval.
-Proof.
+where $\Delta_{T_{1}},\Delta_{T_{2}}$, and $\Delta_{T_{3}}$ are the performances for each interval.
+Proof of ˜B.4.
 We split the total horizon into three intervals of initial batch $T_{1}=[0,f]$, intermediate process $T_{2}=[f,kf]$, and final batch $T_{3}=[kf,T]$.
-We use the first-order method for the first and final batches and LP-based method for intermediate process.
-Notice that for the initial resolving batch, we maintain the classical analysis for first-order method since it has not reached to the mixture with LP-based method.
+We use the first-order method for the first and final batches and the LP-based method for intermediate processes.
+Notice that for the initial resolving batch, we maintain the classical analysis for the first-order method since it has not reached the mixture with the LP-based method.
 Then based on (B.2), we define
@@
 To provide a tighter analysis, we examine the real-time average resource capacity $\bm{d}_{t}$ instead of the original process $\bm{b}_{t}$.
-Specifically, by Lemma 2.1, we constraint $\bm{d}_{t}$ within a feasible range $\mathcal{D}$.
+Specifically, by Lemma A.1, we constraint $\bm{d}_{t}$ within a feasible range $\mathcal{D}$.
 If $\bm{d}_{t}\notin\mathcal{D}$, we set the dual price to zero to accept all subsequent orders.
-We analyze the remaining resource and constraint violation due to this strategy.
-Note that this is required solely for rigor analysis purpose and is not necessary when running the algorithm.
+We analyze the remaining resources and constraint violation due to this strategy.
+Note that this is required solely for rigor analysis purposes and is not necessary when running the algorithm.
 Lemma B.5 (Dynamics of Resource Usage).
-Under Assumption 1 and 2, there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda_{1},\mu$, and $C_{lp}$ such that
+Under Assumption 2.1 and 2.2, there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda_{1},\mu$, and $C_{lp}$ such that
 $\sum_{t=f}^{kf}\mathbb{E}[(d_{i,t}-d_{i})^{2}]\leq C\log(k).$ (25)
-Proof.
-During the re-solving process, we denote each interval to be $[jf,(j+1)f]$ where $j=1,2,...,k$.
+Theorem B.6 (Regret of Intermediate Process).
+Under Assumption 2.1 and 2.2, following the result in (B.2), we prove the regret satisfies
+$\Delta_{T_{2}}\leq\log\Big(\frac{T}{f}\Big)=\mathcal{O}(\log(k)).$ (26)
+Proof of ˜B.5.
+During the re-solving process, we denote each interval to be $[jf,(j+1)f]$ where j=1,2,,kj=1,2,...,k.
 As the resource usage follows $\bm{b}_{t+1}=\bm{b}_{t}-\bm{a}_{t+1}I(r_{t+1}>\bm{a}_{t+1}^{T}\bm{p}_{{t+1}})$, normalizing both sides, we derive the update of average resource consumption to be:
@@
 This completes the proof.
-Theorem B.3 (Regret of Intermediate Process).
-Under Assumption 1 and 2, following the result in (B.2), we prove the regret satisfies
-$\Delta_{T_{2}}\leq\log\Big{(}\frac{T}{f}\Big{)}=\mathcal{O}(\log(k)).$ (37)
-Proof. We analyze the three terms in (B.2) respectively.
+Proof of ˜B.6.
+We analyze the three terms in (B.2) respectively.
   1. 1.
-    For dual convergence, by Lemma 3.2, we have
+    For dual convergence, by Lemma A.4, we have
     $\sum_{t=f}^{kf}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}$ $=\mathbb{E}\left[\sum_{j=1}^{k-1}\sum_{t=jf+1}^{(j+1)f}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert_{2}^{2}\right]$
@@
 B.4 Regret for First-Order Method
-In this section, we are going to bound $\Delta_{T_{1}}$ and $\Delta_{T_{3}}$ for the regret from first-order method.
+In this section, we are going to bound $\Delta_{T_{1}}$ and $\Delta_{T_{3}}$ for the regret from the first-order method.
 The key point is to analyze the connection between dual prices based on the gradient-updated rule.
@@
 Theorem B.7 (Regret of Initial Batch).
-Under Assumption 1 and 2, we have
+Under Assumption 2.1 and 2.2, we have
 $\Delta_{T_{1}}\leq\left(\frac{m(\bar{a}+\bar{d})^{2}}{2}+\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right)\sqrt{f}.$ (37)
-Proof. We analyze $\Delta_{T_{1}}$ in (B.2) by parts. By Lemma B.3, B.4 in [8], we have
-$\sum_{t=1}^{f}\mathbb{E}[r_{t}x_{t}^{*}-r_{t}x_{t}]\leq\frac{m(\bar{a}+\bar{d})^{2}}{2}\alpha_{t}f,$
-$\mathbb{E}\Big{\lVert}\Big{(}\sum_{t=1}^{f}\bm{a}_{t}x_{t}-f\bm{d}\Big{)}^{+}\Big{\rVert}\leq\frac{1}{\alpha_{t}}\bigg{[}\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\bigg{]}.$
-To achieve the tight upper bound, we select the optimal step size $\alpha_{t}=\frac{1}{\sqrt{f}}$, which then gives us the desired result.
 Lemma B.8 (Constraint Violation).
-Under Assumption 1 and 2, take the learning rate $\alpha_{t}=\alpha$, we bound the constraint violation for the last batch as
+Under Assumption 2.1 and 2.2, take the learning rate $\alpha_{t}=\alpha$, we bound the constraint violation for the last batch as
 $\mathbb{E}\Big\lVert\Big(\sum_{t=kf}^{T}\bm{a}_{t}x_{t}-(T-kf)\bm{d}\Big)^{+}\Big\rVert\leq\Big(\frac{C_{lp}}{\sqrt{kf}}+\frac{C_{b}}{T^{2}}\Big)\cdot\frac{1}{\alpha}+\frac{C_{b}}{\sqrt{(T-kf)kf}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+\frac{C_{b}}{\sqrt{\alpha}}.$ (38)
-Proof. With the update rule of first-order method in Algorithm 2, we have
+Theorem B.9 (Regret of Final Batch).
+Under Assumption 2.1 and 2.2, we have
+$\Delta_{T_{3}}\leq\bigg(\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+\frac{2C_{lp}}{\sqrt{k}}+\frac{C_{lp}\log(f)}{\nu\lambda k}+\frac{4C_{b}}{\sqrt{k}}+2C_{b}\bigg)\cdot f^{1/3}.$ (39)
+Lemma B.10 (Warm-Start First-Order Regret).
+Assume that we have a warm start for the initial dual price $\bm{p}_{0}$ in the first batch, which satisfies $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$.
+Then the regret of the first batch follows
+$\Delta_{T_{1}\text{warm}}\leq\Big(\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+4C_{w}\Big)\cdot f^{1/3}+\log(f).$ (40)
+Proof of ˜B.7.
+We analyze $\Delta_{T_{1}}$ in (B.2) by parts. By Lemma B.3, B.4 in (Gao et al., 2024), we have
+$\sum_{t=1}^{f}\mathbb{E}[r_{t}x_{t}^{*}-r_{t}x_{t}]\leq\frac{m(\bar{a}+\bar{d})^{2}}{2}\alpha f,$
+$\mathbb{E}\Big\lVert\Big(\sum_{t=1}^{f}\bm{a}_{t}x_{t}-f\bm{d}\Big)^{+}\Big\rVert\leq\frac{1}{\alpha}\bigg[\frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\bigg].$
+To achieve the tight upper bound, we select the optimal step size $\alpha=\frac{1}{\sqrt{f}}$, which then gives us the desired result.
+Proof of ˜B.8.
+With the update rule of first-order method in Algorithm 4, we have
 $\bm{p}_{t+1}=[\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t})]^{+}\geq\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t}),$
 which gives us $\bm{a}_{t}x_{t}-\bm{d}\leq\frac{1}{\alpha}(\bm{p}_{t+1}-\bm{p}_{t}).$
-Summarizing on both sides and apply the telescoping, we derive
+Summarizing on both sides and applying the telescoping, we derive
 $\sum_{t=kf}^{T}(\bm{a}_{t}x_{t}-\bm{d})\leq\frac{1}{\alpha}\sum_{t=kf}^{T}(\bm{p}_{t+1}-\bm{p}_{t})=\frac{1}{\alpha}(\bm{p}_{T+1}-\bm{p}_{kf}).$
@@
 $\mathbb{E}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert\leq C_{b}\Big[\frac{1}{\sqrt{(T-kf)kf}}\cdot\frac{1}{\sqrt{\alpha}}+\sqrt{\alpha}+\frac{1}{T^{2}}\Big].$ (41)
-Thus, according to Lemma 3.2 and (46), the expectation of constraint violation satisfies
+Thus, according to Lemma A.4 and (41), the expectation of constraint violation satisfies
 $\mathbb{E}\Big\lVert\Big(\sum_{t=kf}^{T}(\bm{a}_{t}x_{t}-\bm{d})\Big)^{+}\Big\rVert$ $\leq\mathbb{E}\Big\lVert\sum_{t=kf}^{T}(\bm{a}_{t}x_{t}-\bm{d})\Big\rVert$
@@
 This completes the proof.
-Theorem B.5 (Regret of Final Batch).
-Under Assumption 1 and 2, we have
-$\Delta_{T_{3}}\leq\bigg{(}\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+\frac{2C_{lp}}{\sqrt{k}}+\frac{C_{lp}\log(f)}{\nu\lambda k}+\frac{4C_{b}}{\sqrt{k}}+2C_{b}\bigg{)}\cdot f^{1/3}.$ (48)
-Proof. We decompose $\Delta_{T_{3}}$ into three parts according to (B.2) and analyze each term respectively. Since the first-order method re-starts from $\bm{p}_{kf}$, by Lemma 3.2, we have:
+Proof of ˜B.9.
+We decompose $\Delta_{T_{3}}$ into three parts according to (B.2) and analyze each term respectively. Since the first-order method re-starts from $\bm{p}_{kf}$, by Lemma A.4, we have:
 $\sum_{t=kf}^{T}\mathbb{E}\lVert\bm{p}_{t}-\bm{p}^{*}\rVert^{2}$ $\leq\sum_{t=kf}^{T}\Big[\frac{\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha t}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}\alpha\Big]$
@@
 $\leq\frac{C_{lp}\log(f)}{\nu\lambda kf}\cdot\frac{1}{\alpha}+\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha.$ (43)
-Since $t\in[kf,T]$, by Proposition 3.3 in [8], the dual price is far from the origin and thus the positive projection still equals to itself. This indicates that:
+Since $t\in[kf,T]$, by Proposition 3.3 in (Gao et al., 2024), the dual price is far from the origin and thus the positive projection still equals itself. This indicates that:
 $\bm{p}_{t+1}=[\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t})]^{+}=\bm{p}_{t}-\alpha(\bm{d}-\bm{a}_{t}x_{t}),$
 which then gives us $\bm{d}-\bm{a}_{t}x_{t}=\frac{1}{\alpha}(\bm{p}_{t}-\bm{p}_{t+1}).$
-Summarizing on both sides and apply the telescoping, we derive
+Summarizing on both sides and applying the telescoping, we derive
 $\sum_{t=kf}^{T}(\bm{d}-\bm{a}_{t}x_{t})\leq\frac{1}{\alpha}\sum_{t=kf}^{T}(\bm{p}_{t}-\bm{p}_{t+1})=\frac{1}{\alpha}(\bm{p}_{kf}-\bm{p}_{T+1}).$
-Then for the positive binding terms, by Lemma 3.2 and (46), we have
+Then for the positive binding terms, by Lemma A.4 and (41), we have
 $\mathbb{E}\Big\lVert\Big(\sum_{t=kf}^{T}(\bm{d}-\bm{a}_{t}x_{t})\Big)^{B^{+}}\Big\rVert$ $\leq\mathbb{E}\Big\lVert\sum_{t=kf}^{T}(\bm{d}-\bm{a}_{t}x_{t})\Big\rVert$
@@
   1. 1.
-    If $T-kf\leq f^{1/3}$, the total regret must be smaller than the length of the batch, so $\Delta_{T_{3}}\leq T-kf\leq f^{1/3}$.
+    If $T-kf\leq f^{1/3}$, the regret must be smaller than the length of the batch, so $\Delta_{T_{3}}\leq T-kf\leq f^{1/3}$.
   2. 2.
@@
 This completes the proof.
-Lemma B.4 (Warm-Start First-Order Regret).
-Assume that we have a warm start for the initial dual price $\bm{p}_{0}$ in the first batch, which satisfies $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$. Then the regret of first batch follows
-$\Delta_{T_{1}\text{warm}}\leq\Big{(}\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}+4C_{w}\Big{)}\cdot f^{1/3}+\log(f).$ (53)
-Proof. According to Lemma D.3, with $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ and $\alpha<\frac{1}{\nu\lambda}$, we have
+Proof of ˜B.10.
+According to Lemma D.3, with $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$ and $\alpha<\frac{1}{\nu\lambda}$, we have
 $\mathbb{E}\big[\lVert\bm{p}_{f+1}-\bm{p}^{*}\rVert^{2}\mid\bm{p}_{0}\big]\leq\frac{\lVert\bm{p}_{0}-\bm{p}^{*}\rVert^{2}}{\nu\lambda\alpha f}+\frac{\alpha m(\bar{a}+\bar{d})^{2}}{\nu\lambda},$
@@
 $\leq\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f\alpha+\Big(\frac{\log(f)}{\nu\lambda f^{2/3}}+\frac{2}{f^{1/3}}+\frac{2C_{w}}{T^{2}}\Big)\cdot\frac{1}{\alpha}+\frac{2C_{w}}{f^{1/3}\sqrt{f}}\cdot\frac{1}{\alpha\sqrt{\alpha}}+2C_{w}\cdot\frac{1}{\sqrt{\alpha}}.$ (51)
-Thus, take the optimal learning rate $\alpha=f^{-2/3}$, we obtain the regret for the first batch with warm start as:
+Thus, taking the optimal learning rate $\alpha=f^{-2/3}$, we obtain the regret for the first batch with a warm start as:
 $\Delta_{T_{1}\text{warm}}$ $\leq\frac{m(\bar{a}+\bar{d})^{2}}{\nu\lambda}f^{1/3}+\Big(\frac{\log(f)}{\nu\lambda f^{2/3}}+\frac{2}{f^{1/3}}+\frac{2C_{w}}{T^{2}}\Big)\cdot f^{2/3}+\frac{2C_{w}}{f^{1/3}\sqrt{f}}\cdot f+2C_{w}\cdot f^{1/3}$
@@
 Appendix C Main Results
-In this section, we demonstrate the proof for all theoretical results that we proposed in the main body of the paper. We instate Assumption 1 and 2 are satisfied.
-C.1 Proof of Lemma 3.1
-By the boundedness results in Lemma A.1 and Lemma A.2, define
-$\bar{p}\coloneqq\max\left(\frac{\bar{r}}{\underline{d}-\delta},\ \frac{2\bar{r}+m(\bar{a}+\bar{d})^{2}}{\underline{d}}+m(\bar{a}+\bar{d})\right),$
-then we complete the proof.
-C.2 Proof of Theorem 3.1
+In this section, we demonstrate the proof for all theoretical results that we proposed in the main body of the paper. We instate Assumption 2.1 and 2.2 are satisfied.
+C.1 Proof of Theorem 3.1.
 By Theorem B.3, we obtain
@@
 This completes the proof.
-C.3 Proof of Theorem 3.2
+C.2 Proof of Theorem 3.2.
 Combining the results of Theorem B.6, Theorem B.7, and Theorem B.9, for some constant $C_{reg}>0$, we obtain:
@@
   1. 1.
-    If we use LP-based method on the first batch, then we will have $\Delta_{T_{1}}\leq\log(f)$. Total regret follows
+    If we use LP-based method on the first batch, then we will have $\Delta_{T_{1}}\leq\log(f)$. Total performance follows
     $\Delta_{T}\leq\log(\max\{f,k\})+f^{1/3})\leq\log(\sqrt{T})+f^{1/3}\in\mathcal{O}(\log(T)+f^{1/3}).$ (55)
@@
     If we have a warm start for the first batch, with the initialization $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$, then the first batch regret achieves $\Delta_{T_{1}}\in\mathcal{O}(f^{1/3}+\log(f))$ by Lemma B.10.
-Thus, the total regret follows
+Thus, total performance follows
     $\Delta_{T}\leq\log(\max\{f,k\})+2f^{1/3})\in\mathcal{O}(\log(T)+f^{1/3}).$ (56)
@@
 Appendix D Auxiliary Results
-In this section, we provide some auxiliary results to support the proof in the previous three sections. These lemmas focus on pure mathematical derivations.
+In this section, we provide auxiliary results to support the proof in the previous three sections. These lemmas focus on pure mathematical derivations.
 D.1 Technical Support for LP-based Analysis
@@
 $\mathbb{E}(d_{i,(j+1)f}-d_{i})^{2}$ $\leq\mathbb{E}({d}_{i,jf}-d_{i})^{2}+\frac{C_{rec}}{(k-j-1)^{2}f}+\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-d_{i})^{2}]}.$
-Proof. We analyze each term in (B.3). The key technique we used is to take conditional expectations and simplify the double summations.
+Lemma D.2.
+With the recursion relation in (28), there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda,\mu$, and $C_{lp}$ such that the summation of the total deviation of $\bm{d}_{t}$ with the original $\bm{d}$ satisfies:
+$\sum_{j=1}^{k}\mathbb{E}\left[({d}_{i,jf}-{d}_{i})^{2}\right]$ $\leq\frac{C}{f}\log(k).$
+Proof of ˜D.1.
+We analyze each term in (B.3). The key technique we use is to take conditional expectations and simplify the double summations.
   1. (a)
@@
     $\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}}{(T-(j+1)f)^{2}}\right]$
-    =\ = $\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\sum_{s=jf+1}^{(j+1)f}\mathbb{E}\Big{[}(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))(d_{i,jf}-{a}_{i,s}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}))\mid\bm{d}_{jf}\Big{]}\right]$
-    =\ = $\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big{[}(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}\mid\bm{d}_{jf}\Big{]}\right]$
+    $=\$$\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\sum_{s=jf+1}^{(j+1)f}\mathbb{E}\Big[(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))(d_{i,jf}-{a}_{i,s}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}))\mid\bm{d}_{jf}\Big]\right]$
+    $=\$$\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big[(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}\mid\bm{d}_{jf}\Big]\right]$
     $+\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell\neq j,\ell=jf+1}^{(j+1)f}\mathbb{E}\Big[(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))(d_{i,jf}-{a}_{i,j}I(r_{j}>\bm{a}_{j}^{\top}\bm{p}_{jf}^{*}))\mid\bm{d}_{jf}\Big]\right]$
 […]
-    =\ = $\frac{f(\bar{a}+\bar{d})^{2}}{(T-(j+1)f)^{2}}\ \ \ (\text{since }\mathbb{E}\Big{[}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})|\bm{d}_{jf}\Big{]}=0\text{ for binding terms})$
 […]
+    $\leq\$$\frac{(\bar{a}+\bar{d})^{2}}{(k-j-1)^{2}f}.\ \ \ (\text{since }\lfloor T\rfloor=k\cdot f)$
   2. (b)
@@
     $\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}))^{2}}{(T-(j+1)f)^{2}}\right]$
-    =\ = $\frac{\mathbb{E}\Bigg{[}\sum_{\ell,s=jf+1}^{jk+1)f}\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))\cdot({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\Big{|}\bm{d}_{jf}\Big{]}\Bigg{]}}{(T-(j+1)f)^{2}}.$
+[…]
     When $\ell=s$,
     $\mathbb{E}\Big[({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))^{2}\mid\bm{d}_{jf}\Big]\leq\bar{a}^{2}.$
-    When $\ell\neq s$, by Assumption 2,
+    When $\ell\neq s$, by Assumption 2.2,
     $\mathbb{E}\Big[({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\mid\bm{d}_{jf}\Big]$
-    =\ = $\mathbb{E}\Bigg{[}\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\mid\bm{a}_{i},\bm{a}_{s}\Big{]}\mid\bm{d}_{jf}\Bigg{]}$
-    =\ = $\mathbb{E}\Bigg{[}{a}_{i,\ell}\mathbb{E}\Big{[}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big{]}\cdot{a}_{i,s}\mathbb{E}\Big{[}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{s}\Big{]}\mid\bm{d}_{jf}\Bigg{]}$
-    =\ = $\mathbb{E}\Bigg{[}{a}_{i,\ell}(P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{\ell})-P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{\ell}))\cdot{a}_{i,s}(P(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{s})-P(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{s}))\mid\bm{d}_{jf}\Bigg{]}$
+[…]
+[…]
+[…]
 […]
@@
     $\mathbb{E}\left[\frac{(\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}))^{2}}{(T-(j+1)f)^{2}}\right]$
-[…]
-[…]
+    $\leq\$$\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\bar{a}^{2}+\sum_{\ell\neq s,\ell=jf+1}^{(j+1)f}\mu^{2}\bar{a}^{4}\ \mathbb{E}\big[(\bm{p}_{(j+1)f}-\bm{p}_{jf}^{*})^{2}\big|\bm{d}_{jf}\big]\Big|\bm{d}_{jf}\right]$
+    $\leq\$$\frac{1}{(T-(j+1)f)^{2}}\ \Big(f\bar{a}^{2}+f^{2}\mu^{2}\bar{a}^{4}\mathbb{E}[(\bm{p}_{(j+1)f}-\bm{p}_{jf}^{*})^{2}\big|\bm{d}_{jf}]\Big)$
 […]
@@
     $2\mathbb{E}\left[({d}_{i,jf}-d_{i})\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\right)\right]$
-    =\ = $\frac{2}{T-(j+1)f}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big{[}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})\mid\bm{d}_{jf}\Big{]}\cdot({d}_{i,jf}-d_{i})\right]$
-[…]
+    $=\$$\frac{2}{T-(j+1)f}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big[{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})\mid\bm{d}_{jf}\Big]\cdot({d}_{i,jf}-d_{i})\right]$
+    $=\$$0.\ \ \ \text{ (by the definition of binding terms)}$
   4. (d)
@@
     $2\mathbb{E}\left[({d}_{i,jf}-d_{i})\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}\right)\right]$
-    =\ = $\frac{2}{T-(j+1)f}\ \mathbb{E}\Bigg{[}\sum_{\ell=jf+1}^{(j+1)f}({d}_{i,jf}-{d}_{i}){a}_{i,\ell}\ \mathbb{E}\Big{[}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big{]}\Bigg{]}$
+    $=\$$\frac{2}{T-(j+1)f}\ \mathbb{E}\Bigg[\sum_{\ell=jf+1}^{(j+1)f}({d}_{i,jf}-{d}_{i}){a}_{i,\ell}\ \mathbb{E}\Big[I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big]\Bigg]$
+    $\leq\$$\frac{2}{T-(j+1)f}\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Bigg[({d}_{i,jf}-{d}_{i}){a}_{i,\ell}\ \Big[P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{\ell})-P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{\ell})\Big]\Bigg]$
 […]
-[…]
-[…]
-[…]
-    =\ = $\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}.$
+    $\leq\$$\frac{2\mu\bar{a}^{2}}{T-(j+1)f}\sum_{\ell=jf+1}^{(j+1)f}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}\cdot\sqrt{\mathbb{E}[(\bm{p}_{(j+1)f}-\bm{p}_{jf}^{*})^{2}]}\ \ \ \text{(by Cauchy's inequality)}$
+    $\leq\$$\frac{2\mu\bar{a}^{2}}{T-(j+1)f}\cdot\frac{\sqrt{C_{lp}}f}{\sqrt{jf}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}$
+    $=\$$\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}.$
   5. (e)
@@
     $2\mathbb{E}\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\cdot\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}I(jf<\tau)\right)$
+    $\leq\$$2\sqrt{\mathbb{E}\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\right)^{2}}\sqrt{\mathbb{E}\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}\right)^{2}}$
 […]
-[…]
-    =\ = $\frac{2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\frac{1}{j}\mu^{2}\bar{a}^{2}C_{lp}}}{(k-j-1)^{2}f}.$
+    $=\$$\frac{2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\frac{1}{j}\mu^{2}\bar{a}^{2}C_{lp}}}{(k-j-1)^{2}f}.$
 Combining all the terms, we obtain the upper bound as:
@@
 Taking $C_{rec}=(\bar{a}+\bar{d})^{2}+\bar{a}^{2}+\mu^{2}\bar{a}^{4}C_{lp}+2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\mu^{2}\bar{a}^{2}C_{lp}}$ completes the proof.
-Lemma D.2.
-With the recursion relation in (34), there exists a constant $C>0$ depending on $\bar{d},\bar{a},m,\nu,\lambda,\mu$, and $C_{lp}$ such that the summation of the total deviation of $\bm{d}_{t}$ with the original $\bm{d}$ satisfies:
-$\sum_{j=1}^{k}\mathbb{E}\left[({d}_{i,jf}-{d}_{i})^{2}\right]$ $\leq\frac{C}{f}\log(k).$
-Proof. We consider a general sequence $\{z_{j}\}_{j=1}^{k}$ with
+Proof of ˜D.2.
+We consider a general sequence $\{z_{j}\}_{j=1}^{k}$ with
 $z_{j+1}\leq z_{j}+\frac{R}{(k-j-1)^{2}f}+\frac{\sqrt{R}\sqrt{z_{j}}}{(k-j-1)\sqrt{(j+1)f}}$
@@
 Lemma D.3.
-Following the updated rule of first-order method, we derive last dual price satisfies:
+Following the updated rule of first-order method, we derive the last dual price satisfies:
 $\mathbb{E}\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{\nu\lambda(T-kf)kf}}\cdot\frac{1}{\sqrt{\alpha}}+\frac{m(\bar{a}+\bar{d})}{\sqrt{\nu\lambda}}\cdot\sqrt{\alpha}+\frac{\bar{p}}{T^{2}}.$
-Proof. According to Lemma 3.3, take $\alpha_{t}=\alpha<\frac{1}{\nu\lambda}$, we derive the conditional expectation as
+Proof of ˜D.3.
+According to Lemma A.6, take $\alpha_{t}=\alpha<\frac{1}{\nu\lambda}$, we derive the conditional expectation as
 $\mathbb{E}\big[\lVert\bm{p}_{T+1}-\bm{p}^{*}\rVert^{2}\mid\bm{p}_{kf}\big]$ $\leq(1-\nu\lambda\alpha)\mathbb{E}\big[\lVert\bm{p}_{T}-\bm{p}^{*}\rVert^{2}\mid\bm{p}_{kf}\big]+\alpha^{2}m(\bar{a}+\bar{d})^{2}$
@@
 where we use the technique of $\sum_{j=0}^{T-kf-1}\alpha^{2}m(\bar{a}+\bar{d})^{2}(1-\nu\lambda\alpha)^{j}\leq\frac{1-(1-\nu\lambda\alpha)^{T-kf}}{\nu\lambda\alpha}\leq\frac{1}{\nu\lambda\alpha}$ and $(1-\nu\lambda\alpha)^{T-kf}\leq\frac{1}{1+\nu\lambda\alpha(T-kf)}\leq\frac{1}{\nu\lambda\alpha(T-kf)}$.
-By LP-convergence result in Lemma 3.2, we know $\mathbb{E}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$.
-By Proposition 3.3 in [8], we know the event $E\coloneqq\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$ with probability $\mathbb{P}\geq 1-\frac{1}{T^{4}}$.
-By Lemma 3.1, we know $\lVert\bm{p}_{t}\rVert\leq\bar{p}$.
+By LP-convergence result in Lemma A.4, we know $\mathbb{E}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$.
+By Proposition 3.3 in (Gao et al., 2024), we know the event $E\coloneqq\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$ with probability $\mathbb{P}\geq 1-\frac{1}{T^{4}}$.
+By Lemma B.2, we know $\lVert\bm{p}_{t}\rVert\leq\bar{p}$.
 Thus, we have
@@
 This completes the proof.
+Appendix E Supplementary Experiments
+In this section, we provide more experiments to further evaluate the performance of our algorithms.
+We consider a more general and complex input distribution and include additional comparisons with recent methods.
+E.1 New Distribution
+As an extension of Section 4.1, our goal is to evaluate the main algorithms with different choices of re-solving frequency $f\in\{T^{1/3},T^{1/2},T^{2/3}\}$.
+We consider a more complex distribution for reward and resource consumption requests and guarantee that Assumptions 2.1 and 2.2 are still satisfied.
+Consider
+Input III: $\ a_{it}\sim\min(1,\max(0,1+z)),\ r_{t}\sim\text{Unif}[0,1]$
+$\text{ where }z\sim t(1):\text{Student's t-distribution with 1 degree of freedom}.$
+We generate $\{r_{t},\bm{a}_{t}\}_{t=1}^{T}$ from Input III and keep other parameters $T\in[10^{2},10^{6}]$ and $d_{i}\sim\text{Uniform}[1/3,2/3]$ the same in Section 4.1.
+We report the average result over 100 trials for each experiment and use the classic first-order method with $\mathcal{O}(T^{1/2})$ regret (Algorithm˜4) as a baseline.
+Figure 5: Algorithms under New Distribution.
+$T$ First-Order Low freq Mid freq High freq
+Algorithm 1 $10^{2}$ 3.78 1.68 1.26 1.20
+$10^{3}$ 10.69 2.60 2.10 1.54
+$10^{4}$ 36.79 4.79 3.67 2.54
+$0^{5} 11$7.13 10.45 6.17 3.56
+$10^{6}$ 377.84 34.54 14.30 5.48
+Algorithm 2 $10^{2}$ 3.78 1.24 1.17 1.13
+$10^{3}$ 10.69 2.30 1.96 1.52
+$10^{4}$ 36.79 4.12 2.99 2.15
+$0^{5} 11$7.13 9.78 4.88 3.05
+$10^{6}$ 377.84 25.96 8.90 4.86
+Figure 7: Regret for various re-solving frequencies.
+Under the new distribution, our algorithms still exhibit a strong regret performance.
+As shown in Table 7 and Figure 7, we observe that regret decreases as the re-solving frequency increases.
+This trend holds consistently across both algorithms and is consistent with the guarantees of Theorem 3.2.
+E.2 More Comparison
+Building on the comparison of baseline methods in Section 4.2, we evaluate our algorithms against recent works.
+As a reminder, our algorithms employ frequent LP-solving to learn online dual price under continuous support.
+Li et al.
+(2024) considers a similar problem using infrequent LP-solving to update the dual variable but under finite support.
+Figure 8: Algorithms comparison.
+$T$ Regret Algorithm Compute Time (s)
+$0^{3} 11$.92 First-Order 0.002
+11.12 Infrequent LP-based 0.724
+3.77 Algorithm 1 0.790
+3.69 Algorithm 2 0.782
+$10^{4}$ 36.37 First-Order 0.01
+14.20 Infrequent LP-based 0.827
+6.69 Algorithm 1 3.015
+6.35 Algorithm 2 3.112
+$0^{5} 11$0.22 First-Order 0.109
+20.90 Infrequent LP-based 1.028
+16.41 Algorithm 1 52.95
+10.84 Algorithm 2 52.39
+$10^{6}$ 312.26 First-Order 1.106
+28.80 Infrequent LP-based 4.929
+20.58 Algorithm 1 1261.0
+14.30 Algorithm 2 1305.4
+Figure 9: Regret for various algorithms.
+To compare them, we adapt our algorithms to finite support and take the re-solving frequency $f=T^{1/3}$ from Section 4.1; We take the best-performed parameters for the infrequent method where the number of customer types $n=50$ and $\alpha=\beta=0.95$ which control the solving times near start and end.
+We take the horizon over $T\in[10^{2},10^{6}]$ and report the average result over 100 trials for each experiment.
+We still use the first-order method in Algorithm˜4 as a baseline.
+Table 9 and Figure 9 demonstrate the algorithm regret and computation time across different horizons.
+While the infrequent LP-based method has faster computation, our algorithms show a competitive performance in decision optimality and achieve lower regret under finite support.